### v2

# GL1-02-06-24-2-supplement

Abstract
In this paper, we develop a Difference-in-Differences model for discrete, ordered outcomes, building upon elements from a continuous Changes-in-Changes model. We focus on outcomes derived from self-reported survey data eliciting socially undesirable, illegal, or stigmatized behaviors like tax evasion, substance abuse, or domestic violence, where too many “false zeros”, or more broadly, underreporting are likely. We provide characterizations for distributional parallel trends, a concept central to our approach, within a general threshold-crossing model framework. In cases where outcomes are assumed to be reported correctly, we propose a framework for identifying and estimating treatment effects across the entire distribution. This framework is then extended to modeling underreported outcomes, allowing the reporting decision to depend on treatment status. A simulation study documents the finite sample performance of the estimators. Applying our methodology, we investigate the impact of recreational marijuana legalization for adults in several U.S. states on the short-term consumption behavior of 8th-grade high-school students. The results indicate small, but significant increases in consumption probabilities at each level. These effects are further amplified upon accounting for misreporting.
JEL Codes: C21, C25, C31, C51
Keywords: Partially Observable Outcomes, Changes-in-Changes, Identification, Measurement Error
1 Introduction
The Difference-in-differences (DiD) method is widely recognized in the econometric literature as a quasi-experimental technique for identifying causal mean parameters. While this method has been central in understanding the average effects of reforms or policy changes, there has been a growing interest in capturing the full spectrum of counterfactual outcomes as policy interventions may have multifaceted impacts across the entire distribution. Consequently, quasi-experimental methods exploiting pre- and post-treatment data have emerged over the past two decades to identify and estimate such distributional treatment effects (see e.g., Roth and Sant’Anna, 2023, for an overview). A prominent example of this more comprehensive approach is the Changes-in-Changes (CiC) model developed by Athey and Imbens (2006). This model is particularly valued for its invariance to transformations of the outcome and its intuitive interpretation, and has gained popularity in recent applied (e.g., Gregory and Zierahn, 2022; Assunçao et al., 2023) and theoretical (e.g., Huber et al., 2022; Ghanem et al., 2023) literature. The key assumption is that the distribution of unobservables governing counterfactual outcomes remains stable over time within the treated and comparison groups.
This paper presents a generalization of the the commonly used Logit or Probit DiD method to discrete, ordered outcomes, building upon elements from a continuous CiC model. Central to our approach is the concept of distributional parallel trends, which we leverage for the latent outcome variable within a general threshold-crossing model class, using it as our key identification assumption. Our modeling approach is distinct from and nonnested within the discrete CiC model of Athey and Imbens (2006), even though it does share the features of the continuous CiC approach. As a second generalization, we consider discrete outcomes that are potentially underreported, nesting the “misreporting-to-zero” or “false-zero” scenario as a special case. This phenomenon of underreporting is prevalent in survey data on discrete ordered outcomes of count, ordinal, or binary nature, especially when socially undesirable, illegal, or stigmatized behaviors and actions are elicited. Common examples include the individual consumption of “social bads” such as marijuana (cannabis) and alcohol, and sensitive issues like abortion or domestic violence incidents. These behaviors or events often lead to a significant proportion of “false zeros” in survey data, due to individuals’ reluctance to report such incidents truthfully. [^1] For example, Brown et al. (2018) found that the actual incidence of marijuana consumption in an Australian survey was nearly double the reported figure (23% vs. 12.2%), while Caudill and Mixon (2005) identified about 20 percentage points underreporting in cheating behavior among undergraduate students in a U.S. survey. See also Nguimkeu et al. (2019) for other examples. Such partial observability, without additional constraints on the Data Generating Process, poses significant challenges for both (point) identification and estimation.
The paper makes several contributions to the literature. Firstly, for the case where outcomes are assumed to be reported correctly, we propose a framework for identification and estimation of their causal effects across the entire distribution. Specifically, we consider a general ordered choice model with strictly stationary error terms, and demonstrate that the distributional parallel trends assumption is a natural choice for discrete outcomes. We provide two characterizations of distributional parallel trends on the latent outcomes within this threshold-crossing framework: we establish that distributional parallel trends hold if and only if a “parallel trends”-type assumption holds for the functional median (mean) and the scaling parameters. The latter holds if and only if these two functional parameters admit a distinct form of separability in group and time. These characterizations are of independent interest and provide a direct pathway to two key point identification results of the paper. In particular, we show that distributional parallel trends point identifies the counterfactual distribution of the latent outcome in a parametric ordered choice setup when no other covariates are present. We also demonstrate that, when additional covariates are present, the counterfactual distribution are point identified even in the absence of distributional assumptions on the error term, i.e., when the model is semiparametric. Our results nest the commonly used Logit or Probit DiD for discrete responses (see Puhani, 2012; Wooldridge, 2023), which feature in the applied literature either as main specification for binary (Meyer and Rosenbaum, 2001; Finkelstein, 2007; Mayer et al., 2014) and ordered choice outcomes (e.g., Sequeira, 2016), or as “robustness” specification (e.g., Marie and Zölitz, 2017; Hvide and Jones, 2018; Hoekstra and Sloan, 2022). These models typically draw on the conventional parallel trends that exclusively affect the mean parameters. Moreover, when additional covariates are present, our identification result can directly be translated into an estimation procedure that accommodates the discrete nature of the outcome. This is in contrast to Athey and Imbens (2006), where the proposed estimators are not directly tailored to these types of outcomes and do not allow for structural estimation.
As a second contribution, this paper extends the DiD framework to account for underreporting, a scenario leading to partially observable data, where the true outcomes are correctly recorded if and only if they are both actual and reported (see Poirier, 1980). We generalize the conventional ”false-zero” reporting by allowing survey respondents to report values that are below the actual level, not limited to zero only. We impose restrictions neither on the dependence structure between the unobservables governing the outcome and reporting the decisions, nor on their evolution across groups and over time. In this respect, our approach is distinct from the literature on binary choice models with two-sided misreporting (or misclassification) that typically assume the independence between the unobservables (see Hausman et al., 1998; Lewbel, 2000a). Addressing the identification and estimation challenges posed by such partially observed outcome models, we propose a set of sufficient conditions for the identification of causal parameters. Our setup, though akin to the parametric bivariate choice model with fully observed outcomes (e.g., Mourifié and Méango, 2014; Han and Vytlacil, 2017; Han and Lee, 2019), is notably more complex in its analysis due to the partial observability. We demonstrate that having an exclusion restriction in each equation (outcome and reporting), where the excluded variables are continuous, along with the real analytic condition on the joint distribution of unobservables, is sufficient for point identification of the distributional treatment effects on the treated. This real analytic assumption, a strong form of smoothness imposed on the copula and the marginal distribution functions of the unobservables, has been used for identification in various contexts including quantile sample selection (Arellano and Bonhomme, 2017), triangular models with endogenous binary treatment (Pereda-Fernández, 2023), and random coefficient discrete choice models (Fox et al., 2012; Fox and Gandhi, 2016). While not being “constructive” and lacking a direct pathway for estimation, it permits the use of a wide range of well-established marginal distribution and copula functions, thereby offering practical flexibility. On the other hand, two exclusion restrictions are required to identify the parameters in both outcome and reporting equations. This is common in multi-index outcome models (e.g., Donkers and Schafgans, 2008; Lewbel and Lin, 2022), but may be hard to find in some applications. However, we note that this requirement for two exclusion restrictions can be relaxed if the focus is solely on the parameters of one of these equations. In addition, we discuss possibilities to device specification tests that allow to check the functional specification in practice. Crucially, our approach accounts for the potential endogeneity of reporting behavior with respect to treatment status, a frequent concern in observational studies (e.g., Arulampalam et al., 2017). This aspect is particularly pertinent in our empirical application, where we examine how reporting behavior may shift in response to treatment.
The theoretical identification results are complemented by an extensive Monte Carlo study that investigates the finite sample performance across various exclusion restriction scenarios and distributional choices related to the real analytic copula function. Corroborating that parameters of the partially observable outcome model remain unidentified without exclusion restrictions as postulated by Poirier (1980), our study reveals a relatively sizable and slow-decreasing bias in the Maximum Likelihood (ML) estimator for certain parameters, when a single continuous excluded variable is employed in only one of the equations. This pattern persists across both Frank and Clayton copula functions. Crucially, the inclusion of two excluded variables, one per equation, substantially improves finite sample performance. This leads to a rapid decline in the bias of the ML estimator for all model parameters, offering practical guidance for applying these methods in real-world scenarios.
Finally, as an empirical contribution, we investigate the impact of recreational marijuana legalization for adults in several U.S. states on the short-term consumption behavior of 8th-grade high-school students. The focus on the last group is particularly relevant as they are widely believed to be most vulnerable to the associated health risks of marijuana consumption. Unlike existing studies which examined the effects of legalization of medicinal or recreational marijuana usage on the mean (cf. Anderson et al., 2015; Wen et al., 2015; Cerdá et al., 2017; Hollingworth et al., 2022), or on the timing of the first (early age) consumption (Williams and Bretteville-Jensen, 2014), we consider the effects of the legalization of recreational marijuana use on the entire consumption distribution of the aforementioned high-school population in the treated states of California, Maine, Nevada, and Massachusetts, thus allowing for the possibility of heterogeneous effects of the reforms. Moreover, in contrast to the existing literature, we address the concerns that consumption behavior may not always be reported truthfully, and that reporting behavior may have been affected by legalization through shifts in stigma, public perception, or general attitudes towards marijuana consumption. Our analysis reveals that, while initial findings without covariates indicate no significant changes, incorporating covariates uncovers a significant decrease in non-consumption probability by around 2 percentage points, and significant increases in consumption probabilities by approximately 1 percentage point at each level. Notably, after accounting for misreporting, the effect sizes increase by more than 50% at each level. Despite these shifts in consumption, our findings do not corroborate significant changes in reporting behavior.
The paper is organized as follows. Section 2 outlines the model setup, detailing the observation rule for self-reported outcomes and defining the distributional treatment effects of interest. Section 3 introduces the concept of distributional parallel trends, providing necessary and sufficient conditions for them to hold within a general threshold-crossing framework. Section 4 addresses point identification of model parameters when outcomes are correctly reported, starting with the scenario without covariates and then moving to a semiparametric model with unspecified error term distributions. Section 5 extends the framework to account for underreporting, establishing point identification in a corresponding semiparametric model. Section 6 is dedicated to point identification of distributional treatment effects on the treated, both with and without misreporting. Section 7 discusses both parametric and semiparametric estimation for these distributional treatment effects. Section 8 presents extensive Monte Carlo simulations to assess the finite sample behavior of the ML estimator in the scenario of underreported outcomes, covering various exclusion restriction scenarios and contrasting different index and error variances. Section 9 empirically investigates the effects of recreational marijuana legalization on the consumption behavior of 8th grade high-school students. Section 10 concludes. Finally, an alternative characterization of distributional parallel trends can be found in Appendix A, while Appendix B contains the proofs of all propositions in the paper. A proof for the result of Appendix A, details on the calculation of the standard errors, an alternative estimation procedure as well as additional simulation and empirical results can be found in the supplementary material.
2 Setup
Let $Y(1)$ and $Y(0)$ denote the potential outcomes for individuals that take on a finite number of $J+1$ values (in a pre-specified period of time) with and without treatment, respectively. In light of the empirical application of Section 9, we will refer to $Y(1)$ and $Y(0)$ as potential consumption levels in the following. Also, for simplicity, we will focus on the case of two groups and two time periods exclusively. The extension to several (pre-treatment) periods is immediate at the cost of more cumbersome notation. We define $G\in\{0,1\}$ as the group membership indicator (where group 1 is the treated group) and $T\in\{0,1\}$ as the time period indicator (where period 1 is after the treatment). The binary treatment indicator for individuals is given by $D=G\cdot T$, so that $D=1$ if $(G,T)=(1,1)$, and $D=0$ otherwise. To streamline the notation, we adopt the shorthand as in Athey and Imbens (2006):
$Y_{gt}(d)\stackrel{{\scriptstyle}}{{\sim}}Y(d)\mid G=g,T=t$
for each $d\in\{0,1\},g\in\{0,1\}$ and $t\in\{0,1\}$, where $\sim$ denotes “is distributed as”. [^2] We will adopt the same notation for all other random variables in this paper.
As mentioned in the introduction, a prevalent issue in analyzing sensitive behaviors, such as the consumption of “social bads,” is the presence of “false zeros” or, more broadly, underreporting. Individuals often conceal their actual participation in such behaviors when responding to retrospective surveys (Nguimkeu et al., 2019). This leads to a partially observable outcome model where a specific outcome value is observed in the data if and only if the actual outcome corresponds to that value and is reported truthfully (cf. Poirier, 1980). To formalize this idea, let $R(d)$ denote the potential “reporting intentions” for $d=0,1$. Importantly, note that this formulation allows reporting behavior to change in response to the treatment status $d$. For example, changes in the legalization of recreational marijuana usage for adults may alter the perceived harmfulness or stigma associated with its consumption, thereby also affecting individuals’ willingness to report truthfully. Using similar shorthand as before, we continue to denote
$R_{gt}(d)\stackrel{{\scriptstyle}}{{\sim}}R(d)\mid G=g,T=t$
for each $d\in\{0,1\},g\in\{0,1\}$ and $t\in\{0,1\}$. Here, we interpret $R_{gt}(d)$ as the (potential) level up to which an individual commits to reporting her consumption truthfully. The potential outcome of interest, $C_{gt}(d)$, $d=0,1$, is then given by
$C_{gt}(d)=\min\left\{Y_{gt}(d),R_{gt}(d)\right\}.$ (1)
Thus, respondents will truthfully report their consumption $Y_{gt}(d)$ if and only if it is less than or equal to the reporting intention $R_{gt}(d)$. For instance, if $Y_{gt}(0)=2$ and $R_{gt}(0)=1$, then, in the absence of treatment, the respondent would only report consumption at level 1 even though the actual level is 2. It is worth noting that this formulation of underreporting behavior in (1) allows for more flexibility than standard “false zero” reporting setups, where respondents may underreport a positive outcome level exclusively as a “false zero” (see Brown et al., 2018; Greene et al., 2018; Nguimkeu et al., 2019). Building on this framework and (1), we may also define the actual observed outcome $C_{gt}$ in the standard manner as a linear combination of the potential outcomes $C_{gt}(0)$ and $C_{gt}(1)$ as:
$C_{gt}=C_{gt}(1)D+C_{gt}(0)(1-D),$
where $C_{gt}(d)=\min\{Y_{gt}(d),R_{gt}(d)\}$ for $d\in\{0,1\}$. The above definition bears similarities with the binary selection setup in Klein et al. (2015). However, additional complications in the identification arise due to the partial observability of the outcome of interest, stemming from the unknown reporting decision (see also Poirier (1980)). That is, unlike a selection setup where the selection indicator is fully observed, $R_{gt}(d)$ can be only partially inferred for each $d$ from the observed joint outcome of individual consumption and reporting decisions. More specifically, both the actual consumption and reporting decisions are observable if and only if individuals have reported the highest consumption level.
Throughout the paper, our primary interest lies in identifying the distributional treatment effects on the treated (DTTs), focusing primarily on the impact of marijuana legalization on the drug consumption of adolescents residing in treated states. That is, in the presence of additional covariates $\mathbf{X}_{gt}$ (recall that the subscripts highlight the possible dependence on the group membership $G=g$ and time period $T=t$), we aim to identify:
$\tau_{11}^{c}(j\mid\mathbf{x}):=$ $\Pr\left(Y_{11}(1)=j\mid\mathbf{X}_{11}=\mathbf{x}\right)-\Pr\left(Y_{11}(0)=j\mid\mathbf{X}_{11}=\mathbf{x}\right),$ (2)
for different levels $j$ and values $\mathbf{x}$ in the support of $\mathbf{X}_{11}$, i.e. values from the covariate distribution in the treated group after the treatment. Alternatively, we may also be interested in the unconditional distributional treatment effect on the treated. This can, depending on whether variation in $\mathbf{X}_{gt}$ plays essential roles in the identification arguments (see Section 4), either be directly addressed with an unconditional version of (2) or be obtained by “marginalization” with respect to the covariate distribution in the treated group after the treatment:
[…]
where $\mathcal{X}_{gt}$ denotes the support of $\mathbf{X}_{gt}$ and ${F}_{\mathbf{X}_{gt}}(\cdot)$ the covariate distribution function for a given group $g$ and time period $t$. Another relevant policy parameter of interest is the distributional treatment effect for the treated on the reporting intention:
$\tau^{r}_{11}(j\mid\mathbf{z}):=$ $\Pr\left(R_{11}(1)=j\mid\mathbf{Z}_{11}=\mathbf{z}\right)-\Pr\left(R_{11}(0)=j\mid\mathbf{Z}_{11}=\mathbf{z}\right),$ (3)
for different levels $j$ and values $\mathbf{z}$ in the support of $\mathbf{Z}_{11}$, where $\mathbf{Z}_{11}$ denotes a covariate vector nonnested within $\mathbf{X}_{11}$, see Section 5. The parameter in (3) provides insights into the possible changes in the probabilities of reporting on individuals in response to the treatment. This may also shed light on how factors like stigma and shame influence the reporting behavior.
Regardless of the estimand of interest, note that the first quantity on the RHS of (2) is generally nonparametrically identified from the data when $Y_{11}$ is observable, whereas the second probability is a counterfactual distribution that cannot be identified without further assumptions. As is customary in the DiD or the synthetic control literature, we therefore focus on modeling the counterfactual outcome $Y_{11}(0)$. To do so, for each group and time period $(g,t)\in\{0,1\}$, we consider a general threshold-crossing model framework for the counterfactual outcomes and introduce a latent outcome variable $Y^{*}_{gt}(0)$:
$Y_{gt}(0)=\left\{\begin{array}[]{lll}0&\text{ if }\quad\kappa_{0}\leq Y^{*}_{gt}(0)\leq\kappa_{1}\\ 1&\text{ if }\quad\kappa_{1}<Y^{*}_{gt}(0)\leq\kappa_{2}\\ 2&\text{ if }\quad\kappa_{2}<Y^{*}_{gt}(0)\leq\kappa_{3},\end{array}\right.$
[…]
$R_{gt}(0)=\left\{\begin{array}[]{lll}0&\text{ if }\quad\iota_{0}\leq R^{*}_{gt}(0)\leq\iota_{1}\\ 1&\text{ if }\quad\iota_{1}<R^{*}_{gt}(0)\leq\iota_{2}\\ 2&\text{ if }\quad\iota_{2}<R^{*}_{gt}(0)\leq\iota_{3},\end{array}\right.$ (7)
where $\iota_{0}<\iota_{1}<\iota_{2}<\iota_{3}$, setting again $\iota_{0}=-\infty$ and $\iota_{3}=\infty$, and (w.l.o.g.) $\iota_{1}=0$ and $\iota_{2}=1$.
3 Characterizing Distributional Parallel Trends
We start by considering the setup in the absence of misreporting, i.e. $C_{gt}(d)=Y_{gt}(d)$ for $d\in\{0,1\}$ almost surely, and provide characterizations of the distributional parallel trends in a general threshold-crossing setup where the unobservable distribution is invariant over time within each group. Specifically, recall that $\mathbf{X}_{gt}$ denotes a vector of covariates for a specific group membership $g$ and time period $t$, and let $\mathcal{X}_{gt}$ and $d_{X_{gt}}$ denote the corresponding support and dimension of that vector, respectively. Then, letting +:=$\mathbb{R}_{+}:=\{x\in\mathbb{R}\mid x>0\}$, we define the latent consumption $Y_{gt}^{\ast}(0)$ to take the following form:
Assumption 1.
$Y^{*}_{gt}(0)=\eta_{gt}(\mathbf{X}_{gt})+\lambda_{gt}(\mathbf{X}_{gt})\varepsilon_{gt},$ (8)
where $\eta_{gt}:\mathbb{R}^{d_{X_{gt}}}\rightarrow\mathbb{R}$ and $\lambda_{gt}:\mathbb{R}^{d_{X_{gt}}}\rightarrow\mathbb{R}_{+}$ are unknown, real-valued functions, $\varepsilon_{gt}$ is strictly stationary within each group given $\mathbf{X}_{gt}$ with zero median (or mean) and unit variance, and has an unknown, continuous, and strictly monotonic conditional cumulative distribution function (CDF) $F_{\varepsilon_{gt}\mid\mathbf{X}_{gt}}$.
Here, $\eta_{gt}(\mathbf{X}_{gt})$ can be viewed as the functional median (mean) parameter of the conditional distribution $Y_{gt}^{\ast}(0)\mid\mathbf{X}_{gt}$, and $\lambda_{gt}(\mathbf{X}_{gt})$ as the functional scaling parameter of that distribution. [^4] To clarify the notation, the functions $\eta_{gt}(\cdot)$ and $\lambda_{gt}(\cdot)$ are assumed to be identical to all units within the group $g$ and in period $t$, while the random variables $Y^{*}_{i,gt}(0)$, $\mathbf{X}_{i,gt}$, $\varepsilon_{i,gt}$ can vary across units. We drop the subscript $i$ to ease the notation. Note that $\eta_{gt}$ and $\lambda_{gt}$ can also be treated as trivial functions if one is interested in unconditional distributions. While the representation of $Y^{*}_{gt}(0)$ in (8) is w.l.o.g. for any given $g\in\{0,1\}$ and $t\in\{0,1\}$ (provided minor regularity conditions on the moments), the strict stationarity on the other hand postulates that the conditional distribution of $\varepsilon_{gt}$ must not change over time within each group. The latter represents a conditional form of Assumption 3.3 in Athey and Imbens (2006) and lies at the heart of both the DiD and the CiC approaches. Finally, note that Assumption 1 also nests various (parametric and semiparametric) threshold crossing models (Klein and Sherman, 2002; Klein et al., 2015) as well as nonlinear DiD models (Kim and Lee, 2017; Wooldridge, 2023).
We next outline the definition of (conditional) distributional parallel trends for the latent outcome variable of interest, $Y^{*}_{11}(0)$, which will be used as key identification assumption throughout the paper:
Definition 1.
The (conditional) distributional parallel trends are said to hold if for every covariate value $\mathbf{x}\in\cap_{g,t}\mathcal{X}_{gt}$, there exists an open interval $I\subseteq\mathbb{R}$, such that
$0<F_{Y^{*}_{11}(0)\mid\mathbf{X}_{11}}(v\mid\mathbf{x})$ $=F_{Y^{*}_{10}(0)\mid\mathbf{X}_{10}}\left(F_{Y^{*}_{00}(0)\mid\mathbf{X}_{00}}^{-1}\left(F_{Y^{*}_{01}(0)\mid\mathbf{X}_{01}}(v\mid\mathbf{x})\mid\mathbf{x}\right)\mid\mathbf{x}\right)<1$ (9)
for all $v\in I$.
We provide two characterizations of the distributional parallel trends within our threshold-crossing framework in Proposition 1 below and Proposition A.1 in Appendix A. The formula in (9) is akin to the identification result of the (unconditional) continuous outcome CiC model of Athey and Imbens (2006), see Theorem 3.1 of their paper. Note that the unconditional distributional parallel trends can be defined similarly if the focus is on the unconditional distribution. Therefore, without considering additional covariates, Definition 1 is graphically illustrated in Figure 1. Here, the counterfactual distribution is constructed by adding the difference in the distributions for the control group over time at each level $q^{\prime}$ (corresponding to the quantile $y^{\ast}$) to the treatment distribution function at each level $q$ (corresponding to quantile $y^{\ast}$ in the treatment distribution before treatment). When $I=\mathbb{R}$, as argued in Athey and Imbens (2006), the quantiles of the counterfactual distribution $Y^{\ast}_{11}(0)$ may be constructed by inverting the equality in (9) for each level $q\in[0,1]$:
$F_{Y_{11}^{*}(0)}^{-1}(q)=F_{Y_{01}^{*}(0)}^{-1}\left(F_{Y_{00}^{*}(0)}\left(F_{Y_{10}^{*}(0)}^{-1}(q)\right)\right).$
Note that the class of models implied by the distributional parallel trends above is generally not nested within the parallel trends model set out in (2) of Roth and Sant’Anna (2023). Specifically, denoting $F_{Y_{gt}(0)}$ as the CDF of the potential outcome in the absence of the treatment, the latter relies on the identifying assumption:
$F_{Y_{11}(0)}(y)=F_{Y_{10}(0)}(y)+F_{Y_{01}(0)}(y)-F_{Y_{00}(0)}(y)$ (10)
for all $y\in\{0,1,2\}$. This direct extension of the linear DiD model for the mean to the entire outcome distribution, while intuitive, may give rise to predictions outside the unit interval as the right-hand side (RHS) is not restricted to lie between zero and one. On the other hand, one may attempt to place a similar “parallel trends”-type assumption on the CDFs of the latent outcome $Y_{gt}^{*}(0)$:
$F_{Y^{*}_{11}(0)}(y)=F_{Y^{*}_{10}(0)}(y)+F_{Y^{*}_{01}(0)}(y)-F_{Y^{*}_{00}(0)}(y)$
for all $y\in\mathbb{R}$. However, this may, in turn, incur undesirable restrictions within some standard location-scale families (see, e.g., Example 2 in Roth and Sant’Anna, 2023). [^5] Taking the family of normal distributions as an example, such “parallel trends”-type assumption for the CDFs holds only if $Y_{gt}^{*}(0)$ is either strictly stationary over time within each group or stable across groups in each period (but not a mixture of these cases or others). In our context, this would consequently restrict the (functional) location and scaling parameters $\eta_{gt}$ and $\lambda_{gt}$ to be identical either across groups or over time. As can be demonstrated in Proposition A.1 in Appendix A, our distributional parallel trends generally maintain considerable flexibility in how the time trend, the group membership, and the observed heterogeneity affect the counterfactual outcome. In this respect, we argue that this assumption is sensible for outcomes with discrete, ordered responses.
Figure 1: Stylized CiC Model
[…]
Note: The figure is a stylized example of a CiC model where the curves in blue denote the marginal distribution functions $F_{Y^{\ast}_{gt}(0)}(\cdot)$ of the counterfactual latent outcome $Y^{\ast}_{gt}(0)$, $g\in\{0,1\}$ and $t\in\{0,1\}$, with the dashed line denoting the imputed function $F_{Y^{\ast}_{11}(0)}(\cdot)$. The curve in red on the other hand denotes the marginal distribution function of the latent outcome $Y_{11}^{\ast}(1)$ under treatment.
Our first result characterizes the necessary and sufficient condition on the functional parameters for distributional parallel trends to hold:
Proposition 1.
Suppose Assumption 1 holds. Then, distributional parallel trends defined in Definition 1 hold if and only if
$\eta_{11}(\mathbf{x})-\frac{\lambda_{11}(\mathbf{x})}{\lambda_{10}(\mathbf{x})}\eta_{10}(\mathbf{x})=\eta_{01}(\mathbf{x})-\frac{\lambda_{01}(\mathbf{x})}{\lambda_{00}(\mathbf{x})}\eta_{00}(\mathbf{x})\quad\text{and}\quad\frac{\lambda_{11}(\mathbf{x})}{\lambda_{10}(\mathbf{x})}=\frac{\lambda_{01}(\mathbf{x})}{\lambda_{00}(\mathbf{x})},$ (11)
for all covariate values $\mathbf{x}\in\cap_{g,t}\mathcal{X}_{gt}$.
Within the general threshold-crossing model given by Assumption 1, Proposition 1 shows that the distributional parallel trends hold if and only if a weighted “parallel trends”-type assumption holds for the functional median (mean) parameters and a proportional “parallel trends”-type holds for the scaling parameters. We note that if the time exclusively affects the median (or mean) parameters, then the results align with the conventional parallel trends, thereby nesting the commonly used Logit or Probit DiD. More specifically, as can be seen in the first part of (11), the change in the median (or average) latent outcome for the treated group over time given the covariates, $\eta_{11}(\mathbf{x})-\eta_{10}(\mathbf{x})$, in the absence of treatment, is extrapolated by the change for the control group over the same period, $\eta_{01}(\mathbf{x})-\eta_{00}(\mathbf{x})$, after an appropriate adjustment through the scaling parameters, serving as tailored weighting factors. Here, the weights play the role of correcting the part of the common time trends that cannot be eliminated by the simple mean difference over time within groups. On the other hand, the second part of (11) states that the proportional change in the scaling parameters for the treated group over time, $\frac{\lambda_{11}(\mathbf{x})}{\lambda_{10}(\mathbf{x})}$, in the absence of treatment, must be identical to the proportional change for the treated group over the same periods, $\frac{\lambda_{01}(\mathbf{x})}{\lambda_{00}(\mathbf{x})}$. The equality in (11) will serve as a fundamental cornerstone for identifying the counterfactual distributions in Sections 4 and 5.
It is important to recognize that the results of Proposition 1 remain applicable even when strictly monotonic transformations are applied to the latent outcome $Y_{gt}^{\ast}(0)$, a characteristic inherent to the continuous CiC model (see Athey and Imbens, 2006). [^6] This differs from the invariance of parallel trends discussed in Roth and Sant’Anna (2023), which focuses on the average untreated potential outcomes. This aspect is particularly relevant for data with discrete ordered outcomes, which are often re-categorized in different ways to verify the “robustness” of empirical findings. For instance, in a $J+1$-level scenario (ranging from 00 to $J$), it is common to group the first $j_{1}$ (with $0\leq j_{1}<J$) levels into a new level 0, and then levels $j_{1}+1$ to $j_{2}$ (with $j_{1}+1\leq j_{2}<J$) into a new level 1. Such reclassification constitutes, in essence, a strictly monotonic transformation of $Y_{gt}^{\ast}(0)$, thereby preserving the applicability of Proposition 1.777Consider the model with $J+1$ levels: $Y_{gt}(0)=\sum_{j=0}^{J}\mathbb{I}\left[\kappa_{j}<Y^{*}_{gt}(0)\leq\kappa_{j+1}\right]$, where $\kappa_{j}<\kappa_{j+1}$ for all $0\leq j\leq J$, $\kappa_{0}=-\infty$, $\kappa_{J+1}=\infty$, and $\mathbb{I}\left[A\right]$ is the indicator function that equals 1 when $A$ is true and 0 otherwise. Reclassifying the first $j_{1}$ levels as level 0 and then levels $j_{1}+1$ to $j_{2}$ as level 1 implements the monotonic transformation $v\mapsto\frac{v-\kappa_{j_{1}}}{\kappa_{j_{2}}-\kappa_{j_{1}}}$ across all terms in the inequalities, setting $\kappa_{j_{1}}=0$ and $\kappa_{j_{2}}=1$.
Finally, notice that in Appendix A, we establish a second equivalence result to further characterize the class of models that generate the distributional parallel trends under Assumption 1. The result helps to clarify the link between the CiC models considered in Athey and Imbens (2006) and our approach. While being nonnested within the discrete CiC models with covariates in their paper, our model does share the features of an extended class of their continuous CiC model with modified assumptions accommodating additional covariates, see Appendix A for details.
4 Model Without Underreporting
4.1 No Covariates
In this subsection, we establish a point identification result for the counterfactual distribution of interest in the absence of misreporting and other covariates $\mathbf{X}_{gt}$. That is, setting $\eta_{gt}(\cdot)$ and $\lambda_{gt}(\cdot)$ to be constant functions in $\mathbf{X}_{gt}$ (almost surely), the latent consumption $Y_{gt}^{\ast}(0)$ takes the form:
$Y^{*}_{gt}(0)=\eta_{gt}+\lambda_{gt}\varepsilon_{gt},$ (12)
where $\eta_{gt}\in\mathbb{R}$ and $0<\lambda_{gt}<\infty$ by Assumption 1 (see also Yamauchi, 2020). We readily obtain the following impossibility result for the model without covariates and without “false zeros” in the data:
Proposition 2.
Suppose Assumption 1 holds and consider the model in (12). Also, suppose (unconditional) distributional parallel trends hold as in Definition 1. Then, $\eta_{gt}$ and $\lambda_{gt}$ are (point) identified if and only if $F_{\varepsilon_{gt}}^{-1}\left(\mathrm{Pr}\left(Y_{gt}=0\right)\right)$ and $F_{\varepsilon_{gt}}^{-1}\left(\mathrm{Pr}\left(Y_{gt}\leq 1\right)\right)$ are known.
Proposition 2 highlights the fact that in a setup with no further structure and no additional covariates, point identification of $\eta_{gt}$ and $\lambda_{gt}$ is impossible without the knowledge of the distribution function $F_{\varepsilon_{gt}}$. [^8] Unlike the if-and-only-if characterizations for distributional parallel trends in threshold-crossing models in this paper, Yamauchi (2020) establishes a point identification result for the model in (12) where $F_{\varepsilon_{gt}}$ is assumed to be known. The paper does, however, not consider the case where $F_{\varepsilon_{gt}}$ is unknown (see Subsection 4.2) or where the outcome is potentially misreported (see Section 5). The proposition can be straightforwardly generalized to conditional on covariates. We note that, with known $F_{\varepsilon_{gt}}$, the parameters are uniquely determined by the observable quantities
$\eta_{gt}=\frac{-F_{\varepsilon}^{-1}\left(\mathrm{Pr}\left(Y_{gt}=0\right)\right)}{F_{\varepsilon}^{-1}\left(\mathrm{Pr}\left(Y_{gt}\leq 1\right)\right)-F_{\varepsilon}^{-1}\left(\mathrm{Pr}\left(Y_{gt}=0\right)\right)}\quad\text{and}\quad\lambda_{gt}=\frac{1}{F_{\varepsilon}^{-1}\left(\mathrm{Pr}\left(Y_{gt}\leq 1\right)\right)-F_{\varepsilon}^{-1}\left(\mathrm{Pr}\left(Y_{gt}=0\right)\right)}$
for $(g,t)\neq(1,1)$. Moreover, Proposition 1 establishes that:
$\eta_{11}=\eta_{01}+\frac{\eta_{10}-\eta_{00}}{\lambda_{00}/\lambda_{01}}\quad\text{and}\quad\lambda_{11}=\frac{\lambda_{01}\lambda_{10}}{\lambda_{00}}.$
Nevertheless, employing a parametric stand naturally calls for caution regarding potential misspecification. Therefore, we will ease this distributional assumption in Subsection 4.2 below.
4.2 Including Covariates
Next, we consider a semiparametric version of our model for $Y_{gt}^{\ast}(0)$, where the distribution of $\varepsilon_{gt}$ remains unspecified and covariates $\mathbf{X}_{gt}$ enter the model. We will exploit variation in the latter to obtain a point identification result for the counterfactual $Y_{11}^{\ast}(0)$ distribution. At the same time, in order to facilitate the comparison with more standard ordered choice models from the literature and in light of the extension to misreporting in Section 5, we will place further structure on the functional form of $\eta_{gt}(\cdot)$ and $\lambda_{gt}(\cdot)$, respectively. Specifically, for every $(g,t)\in\{0,1\}$, we assume that, with probability one, $\eta_{gt}(\mathbf{X}_{gt})=\overline{\mathbf{X}}_{gt}^{\prime}\overline{\boldsymbol{\eta}}_{gt}$ with $\overline{\mathbf{X}}_{gt}=(1,\mathbf{X}_{gt}^{\prime})^{\prime}$ and $\overline{\boldsymbol{\eta}}_{gt}=(\eta_{c,gt},\boldsymbol{\eta}_{gt}^{\prime})^{\prime}$, while $\lambda_{gt}(\mathbf{X}_{gt})=\lambda_{gt}$. Thus:
$Y_{gt}^{\ast}(0)=\overline{\mathbf{X}}_{gt}^{\prime}\overline{\boldsymbol{\eta}}_{gt}+\lambda_{gt}\varepsilon_{gt}.$ (13)
Equation (13) is a standard functional specification for ordered choice models (Klein and Sherman, 2002), and allows the median of $Y_{gt}^{\ast}(0)$ to vary with the covariates $\mathbf{X}_{gt}$, but not the scaling parameter.999 Note that interpretation of $\overline{\mathbf{X}}_{gt}^{\prime}\overline{\boldsymbol{\eta}}_{gt}$ as conditional median will be important for the identification of the intercept parameter $\eta_{c,gt}$ (e.g., see Manski, 1988). Letting $\overline{\mathcal{X}}_{gt}$ the support of $\overline{\mathbf{X}}_{gt}$, we make the following assumptions:
Assumption 2.
For each $g\in\{0,1\}$ and $t\in\{0,1\}$, $\varepsilon_{gt}$ is independent of $\mathbf{X}_{gt}$.
Assumption 3.
For each $(g,t)\neq(1,1)$, there exists at least one $k\in\{1,\ldots$, $K\}$ such that $\eta_{k,gt}\neq 0$ and
$\operatorname{Pr}\left(X_{k,gt}\in\left(a_{1},a_{2}\right)\mid\overline{\mathbf{X}}_{-k,gt}\right)>0$
with probability one for all open intervals $\left(a_{1},a_{2}\right)\subseteq\mathbb{R}$, where $\overline{\mathbf{X}}_{-k,gt}$ denotes the subvector of $\overline{\mathbf{X}}_{gt}$ excluding the $k$th slope. W.l.o.g., let $k=1$.
Assumption 4.
For each $g\in\{0,1\}$ and $t\in\{0,1\}$, $\overline{\mathcal{X}}_{gt}$ is not contained in a proper linear subspace of $\mathbb{R}^{k+1}$.
Assumption 5.
It holds that $\mathcal{X}_{11}\subseteq\cap_{(g,t)\neq(1,1)}\mathcal{X}_{gt}$.
Assumption 2 states covariate independence. Although this is often challenging to justify in practice, it remains a standard condition in the semiparametric estimation and identification literature, playing a crucial role in the identification argument (cf. Athey and Imbens, 2006). Assumption 3 on the other hand requires the existence of a “special regressor” with large support, an assumption frequently used in models with discrete responses to ensure point identification of the parameters (see Manski, 1988; Lewbel, 2000b; Han and Vytlacil, 2017; Khan et al., 2021). [^10] In principle, other identifying strategies from the semiparametric literature, such as exploiting the differentiability with respect to the continuous variables (e.g., Klein and Spady, 1993), are also possible with more complicated arguments. Finally, Assumption 4 is a standard identification assumption ruling out perfect multicollinearity, while Assumption 5 postulates a common support of $\mathbf{X}_{11}$ and the other subgroups.
We have the following identification result:
Proposition 3.
Suppose Assumptions 1-5 hold and consider the model in (13). Also, suppose (unconditional) distributional parallel trends hold as in Definition 1. Then, for each $g\in\{0,1\}$, $t\in\{0,1\}$ and $v\in\mathbb{R}$, $\overline{\boldsymbol{\eta}}_{gt}$, $\lambda_{gt}$ and $F_{\varepsilon_{gt}}(v)$ are identified. In particular, it holds for the counterfactual median and scaling parameters that:
$\overline{\boldsymbol{\eta}}_{11}=\overline{\boldsymbol{\eta}}_{01}+\frac{\overline{\boldsymbol{\eta}}_{10}-\overline{\boldsymbol{\eta}}_{00}}{\lambda_{00}/\lambda_{01}}\quad\text{and}\quad\lambda_{11}=\frac{\lambda_{01}\lambda_{10}}{\lambda_{00}}.$
In contrasting the identification result of the counterfactual distribution in Proposition 3 with the point identification in Theorem 4.4 of Athey and Imbens (2006) for the discrete CiC model with covariates, we note that while Proposition 3 imposes more structural constraints such as a linear index, it offers a direct pathway to an estimation procedure that naturally accommodates the discreteness of the outcome, as we will outline in Section 7. This presents a clear estimation advantage compared to the methods proposed by Athey and Imbens (2006), which are not specifically tailored to discrete outcomes or the inclusion of covariates. However, it is important to stress that, akin to Theorem 4.4 in Athey and Imbens (2006), the identification result in Proposition 3 relies on strong assumptions that, while commonly invoked in the literature on semiparametric discrete outcome models, may sometimes be hard to justify.
5 Underreported Outcomes
This section addresses the scenario where outcomes may be underreported. It considers cases where, for $d\in\{0,1\}$, $C_{gt}(d)\neq Y_{gt}(d)$ with positive probability, a situation often referred to as “partial observability”. In nonlinear models, such discrepancies between true and reported outcomes generally lead to a failure of (point) identification. This issue persists even in fully parametric models unless additional exclusion restrictions are assumed (cf. Poirier, 1980).
Building upon the observation rule outlined in (1) and the latent outcome formulation of $R_{gt}(0)$ in (7), and drawing on the model for $Y_{gt}^{\ast}(0)$ presented in (13) from Section 4.2, we are now prepared to introduce a model for $R_{gt}^{\ast}(0)$. Similar in structure to the one for $Y_{gt}^{\ast}(0)$, this model incorporates additional covariates $\mathbf{Z}_{gt}$:
$R_{gt}^{\ast}(0)=\overline{\mathbf{Z}}_{gt}^{\prime}\overline{\boldsymbol{\pi}}_{gt}+\zeta_{gt}\nu_{gt},$ (14)
where $\overline{\mathbf{Z}}_{gt}=(1,\mathbf{Z}_{gt}^{\prime})^{\prime}$ and $\overline{\boldsymbol{\pi}}_{gt}=(\pi_{c,gt},\boldsymbol{\pi}_{gt}^{\prime})^{\prime}$. The covariate vectors $\mathbf{X}_{gt}$ and $\mathbf{Z}_{gt}$ are allowed to share common elements. It is required, though, that each vector includes at least one distinct continuous variable not present in the other. W.l.o.g., we will assume that all elements, except for the first, in both $\mathbf{X}_{gt}$ and $\mathbf{Z}_{gt}$ are common. Consequently, the augmented vectors $\overline{\mathbf{X}}_{gt}$ and $\overline{\mathbf{Z}}_{gt}$ are of the same dimension $k+1$.
Under Assumption 6 below, for each $j\in\{0,1,2\}$, the upper tail probability of consumption in the absence of treatment can be expressed as:
$\Pr\left(C_{gt}(0)\geq j\mid\mathbf{X}_{gt},\mathbf{Z}_{gt}\right)$ $=\Pr\left(Y_{gt}(0)\geq j,R_{gt}(0)\geq j\mid\mathbf{X}_{gt},\mathbf{Z}_{gt}\right)$
$=\Pr\left(\overline{\mathbf{X}}_{gt}^{\prime}\overline{\boldsymbol{\eta}}_{gt}+\lambda_{gt}\varepsilon_{gt}\geq\kappa_{j},\overline{\mathbf{Z}}_{gt}^{\prime}\overline{\boldsymbol{\pi}}_{gt}+\zeta_{gt}\nu_{gt}\geq\iota_{j}\right)$ (15)
$=\overline{C}_{gt}\left(F_{-{\varepsilon_{gt}}}\left(\frac{\overline{\mathbf{X}}_{gt}^{\prime}\overline{\boldsymbol{\eta}}_{gt}-\kappa_{j}}{\lambda_{gt}}\right),F_{-{\nu}_{gt}}\left(\frac{\overline{\mathbf{Z}}_{gt}^{\prime}\overline{\boldsymbol{\pi}}_{gt}-\iota_{j}}{\zeta_{gt}}\right)\right).$
where $F_{-\varepsilon_{gt}}$ and $F_{-\nu_{gt}}$ denote CDFs of $-\varepsilon_{gt}$ and $-\nu_{gt}$, respectively, and $\overline{C}_{gt}(\cdot,\cdot)$ denotes a nonparametric Copula function. As noted in Section 2, the above setup bears similarities with the distributional sample selection model proposed by Chernozhukov et al. (2023), as well as the bivariate Probit model with a binary endogenous regressor considered by Han and Vytlacil (2017). However, a crucial distinction of our model, represented by (15), lies in the fact that the reporting intention $R_{gt}(0)$ is not directly observable. As a result, the second argument in the stated expression cannot be separately identified and estimated. This suggests that the identification results presented in the aforementioned papers are not directly applicable to our context. Specifically, the approach of using a single excluded variable in the second argument, as employed in Theorem 2.1 by Chernozhukov et al. (2023) and Theorem 6.1 by Han and Vytlacil (2017), may not be sufficient for point identifying the parameters inside the marginal distribution in our model, nor for the marginal distribution itself. This insight is supported by the simulation results in Section 8. These simulations suggest that, even within the parametric framework and using fairly large sample sizes, some parameters can still be poorly estimated if only one exclusion restriction is in place.
We make the following assumption:
Assumption 6.
(i) $-\varepsilon_{gt}$ and $-\nu_{gt}$ are strictly stationary within each group with zero median and unit variance, jointly independent of $(\mathbf{X}_{gt}^{\prime},\mathbf{Z}_{gt}^{\prime})^{\prime}$, and have unknown, real analytic and strictly monotonic CDFs $F_{-\varepsilon_{gt}}$ and $F_{-\nu_{gt}}$, respectively.
(ii) For each $g\in\{0,1\}$ and $t\in\{0,1\}$, the copula function $\overline{C}_{gt}(\cdot,\cdot)$ defined in (15) is strictly monotonic and real analytic in each of its coordinates.
Assumption 7.
(i) For each $g\in\{0,1\}$ and $t\in\{0,1\}$, the variables $X_{1,gt}$ and $Z_{1,gt}$ are excluded from the vectors $\mathbf{Z}_{gt}$ and $\mathbf{X}_{gt}$, respectively, while the other elements of them are common.
(ii) For each $(g,t)\neq(1,1)$, $\eta_{1,gt}\neq 0$ and $\pi_{1,gt}\neq 0$ and there exists some Cartesian product of two open intervals $I_{1}\times I_{2}\subseteq\mathbb{R}\times\mathbb{R}$ such that
$\operatorname{Pr}\left(\left(X_{1,gt},Z_{1,gt}\right)\in I_{1}\times I_{2}\mid\overline{\mathbf{W}}_{gt}\right)>0$
with probability one, where $\mathbf{{W}}_{gt}$ denote the vector of regressors that are common to both $\mathbf{{X}}_{gt}$ and $\mathbf{{Z}}_{gt}$ and $\overline{\mathbf{W}}_{gt}=\left(1,\mathbf{W}_{gt}^{\prime}\right)^{\prime}$.
(iii) For each $g\in\{0,1\}$ and $t\in\{0,1\}$, $\mathcal{\overline{X}}_{gt}$ and $\mathcal{\overline{Z}}_{gt}$ are not contained in a proper linear subspace of $\mathbb{R}^{k+1}$.
Assumption 6(i) closely resembles Assumptions 1-2. While the covariate independence is restrictive, as previously discussed, it is important to note that we do not impose any constraints on the correlation between the unobservables $\varepsilon_{gt}$ and $\nu_{gt}$. This allowance caters to scenarios where certain types of “consumers” may be more or less likely to misreport than others. This aspect of our model is distinct from the literature on binary choice models with two-sided misreporting (or misclassification) that typically imposes (conditional) independence between the unobservables governing the outcome and reporting decisions (see Hausman et al., 1998; Lewbel, 2000a). On the other hand, the other parts of Assumption 6 impose further conditions on the joint and marginal distributions of the unobservables. In particular, the real analytic property will allow for identification of the counterfactual distribution based on an extrapolation argument, which has also been featured in the microeconometric (see Arellano and Bonhomme, 2017; Pereda-Fernández, 2023) and the Industrial Organization identification literature (see Iaria and Wang, 2022, and references therein). Specifically, the extrapolation is based on the fact that if two real analytic functions coincide on an open neighborhood, then they coincide everywhere, and in particular also in the limit (if it exists). Identification based on such extrapolation is in contrast to “identification at infinity” type arguments, which require the existence of excluded (instrumental) variables with a large support (e.g., Chamberlain, 1986; Lewbel, 2007; D’Haultfoeuille and Maurel, 2013). Moreover, note that the real analytic property is satisfied by several well-known Copula families such as the Gaussian, Frank, Clayton, Plackett, Joe, or Gumbel Copula families. Indeed, while Assumptions 7(i) spells out the exclusion restrictions, Assumptions 7(ii) “only” imposes mild conditions on continuity of the excluded variables $X_{1,gt}$ and $Z_{1,gt}$, without requiring a “special regressor” as often necessitated in the semiparametric identification literature on discrete response models. On the other hand, the need for an excluded (continuous) variable in each decision equation is common to other identification work dealing with multiple latent outcome equations, see for instance Lewbel et al. (2021). Finally, Assumption 7(iii) extends Assumption 5 to the multivariate case.
We obtain the following identification result:
Proposition 4.
Suppose Assumptions 6-7 hold and consider the models in (13)-(14). Also, suppose distributional parallel trends hold as in Definition 1 for $Y_{gt}^{*}(0)$ as well as $R_{gt}^{*}(0)$. Then, for each $g\in\{0,1\}$, $t\in\{0,1\}$ and $v\in\mathbb{R}$, $\overline{\boldsymbol{\eta}}_{gt}$, $\overline{\boldsymbol{\pi}}_{gt}$, $\lambda_{gt}$, $\zeta_{gt}$, $F_{-\varepsilon_{gt}}(v)$ and $F_{-\nu_{gt}}(v)$ are identified. In particular, it holds for the counterfactual quantities:
$\overline{\boldsymbol{\eta}}_{11}=\overline{\boldsymbol{\eta}}_{01}+\frac{\overline{\boldsymbol{\eta}}_{10}-\overline{\boldsymbol{\eta}}_{00}}{\lambda_{00}/\lambda_{01}}\quad\text{and}\quad\lambda_{11}=\frac{\lambda_{01}\lambda_{10}}{\lambda_{00}},$ (16)
and
$\overline{\boldsymbol{\pi}}_{11}=\overline{\boldsymbol{\pi}}_{01}+\frac{\overline{\boldsymbol{\pi}}_{10}-\overline{\boldsymbol{\pi}}_{00}}{\zeta_{00}/\zeta_{01}}\quad\text{and}\quad\zeta_{11}=\frac{\zeta_{01}\zeta_{10}}{\zeta_{00}}.$ (17)
It should be emphasized that the distributional parallel trends solely set constraints on the change of the marginal distributions of counterfactual outcomes over time. However, these assumptions do not place any constraints on the evolution of the dependence structure between the outcome and reporting decisions. Additionally, it is noteworthy that the identification result presented in Proposition 4 requires knowledge of neither the marginal distributions nor the Copula function, thereby classifying it as semiparametric. Nevertheless, when transitioning to estimation, one has to select specific parametric forms for the Copula function $\overline{C}_{gt}(\cdot,\cdot)$ as well as the distributions $F_{-\varepsilon_{gt}}$ and $F_{-\nu_{gt}}$. This is because the identification based on the real analytic assumption is not “constructive” and does not provide a direct pathway to construct estimators. As discussed earlier, however, due to the partially observable nature of the consumption and reporting decisions, it appears difficult to bypass this real analytic assumption without relying on excluded variables with large support.
Remark 1: Note that, under the assumption of index sufficiency and instrument validity, the above assumption setup not only allows for specification tests of the functional form, but also for the presence of misreporting. Specifically, starting with the test of correct functional form and focusing on a specific choice of $(g,t)\neq(1,1)$, let $M\left(j;\mathbf{X}_{gt}^{\prime}\widetilde{\boldsymbol{\eta}}_{gt},\mathbf{Z}_{gt}^{\prime}\widetilde{\boldsymbol{\pi}}_{gt}\right):=\Pr\left(C_{gt}\geq j\mid\mathbf{X}_{gt}^{\prime}\widetilde{\boldsymbol{\eta}}_{gt},\mathbf{Z}_{gt}^{\prime}\widetilde{\boldsymbol{\pi}}_{gt}\right)$ denote the nonparametric tail probability of observed consumption at level $j\in\{0,1,2\}$, where $\widetilde{\boldsymbol{\eta}}_{gt}$ and $\widetilde{\boldsymbol{\pi}}_{gt}$ are the slope coefficient vectors normalized by $\|\boldsymbol{\eta}_{gt}\|$ and $\|\boldsymbol{\pi}_{gt}\|$, respectively. Note that this tail probability is nonparametrically identified under Assumption 7 and index sufficiency, see for instance Donkers and Schafgans (2008) or Ahn et al. (2018). Then, under Assumptions 6 and 7, we may construct a test for the null hypothesis:
$H_{0}^{C}:\overline{C}_{gt}\left(F_{-{\varepsilon}_{gt}}\left(\frac{\overline{\mathbf{x}}^{\prime}\overline{\boldsymbol{\eta}}_{gt}-\kappa_{j}}{\lambda_{gt}}\right),F_{-{\nu}_{gt}}\left(\frac{\overline{\mathbf{z}}^{\prime}\overline{\boldsymbol{\pi}}_{gt}-\iota_{j}}{\zeta_{gt}}\right)\right)=M\left(j;\mathbf{x}^{\prime}\widetilde{\boldsymbol{\eta}}_{gt},\mathbf{z}^{\prime}\widetilde{\boldsymbol{\pi}}_{gt}\right)$
for all $j\in\{0,1,2\}$ and $(\mathbf{x},\mathbf{z})\in\mathcal{X}_{gt}\times\mathcal{Z}_{gt}$, where $\mathcal{X}_{gt}$ and $\mathcal{Z}_{gt}$ denote compact subsets of the support of $\mathbf{X}_{gt}$ and $\mathbf{Z}_{gt}$, respectively, vs.
$H_{A}^{C}:\overline{C}_{gt}\left(F_{-{\varepsilon}_{gt}}\left(\frac{\overline{\mathbf{x}}^{\prime}\overline{\boldsymbol{\eta}}_{gt}-\kappa_{j}}{\lambda_{gt}}\right),F_{-{\nu}_{gt}}\left(\frac{\overline{\mathbf{z}}^{\prime}\overline{\boldsymbol{\pi}}_{gt}-\iota_{j}}{\zeta_{gt}}\right)\right)\neq M\left(j;\mathbf{x}^{\prime}\widetilde{\boldsymbol{\eta}}_{gt},\mathbf{z}^{\prime}\widetilde{\boldsymbol{\pi}}_{gt}\right)$
for at least some $j\in\{0,1,2\}$ and $(\mathbf{x},\mathbf{z})\in\mathcal{X}_{gt}\times\mathcal{Z}_{gt}$. Similarly, we may also use the fact that $M\left(j;\mathbf{X}_{gt}^{\prime}\widetilde{\boldsymbol{\eta}}_{gt},\mathbf{Z}_{gt}^{\prime}\widetilde{\boldsymbol{\pi}}_{gt}\right)$ is nonparametrically identified under our assumptions to test for the presence of misreporting in our semiparametric model setup. That is, given index sufficiency and Assumption 7, we may test:
$H_{0}^{R}:M\left(j;\mathbf{x}^{\prime}\widetilde{\boldsymbol{\eta}}_{gt},\mathbf{z}^{\prime}\widetilde{\boldsymbol{\pi}}_{gt}\right)=\overline{M}\left(j;\mathbf{x}^{\prime}\widetilde{\boldsymbol{\eta}}_{gt}\right)$
for all $j\in\{0,1,2\}$ and all $(\mathbf{x},\mathbf{z})\in\mathcal{X}_{gt}\times\mathcal{Z}_{gt}$ vs.
$H_{0}^{R}:M\left(j;\mathbf{x}^{\prime}\widetilde{\boldsymbol{\eta}}_{gt},\mathbf{z}^{\prime}\widetilde{\boldsymbol{\pi}}_{gt}\right)\neq\overline{M}\left(j;\mathbf{x}^{\prime}\widetilde{\boldsymbol{\eta}}_{gt}\right)$
for at least some $j\in\{0,1,2\}$ and $(\mathbf{x},\mathbf{z})\in\mathcal{X}_{gt}\times\mathcal{Z}_{gt}$, where $\overline{M}\left(j;\mathbf{X}_{gt}^{\prime}\widetilde{\boldsymbol{\eta}}_{gt}\right):=\Pr\left(C_{gt}\geq j\mid\mathbf{X}_{gt}^{\prime}\widetilde{\boldsymbol{\eta}}_{gt}\right)$. We comment on suitable test statistics for the above tests in Remark 3 in Section 7.2.
6 Distributional Treatment Effects
To identify the DTTs, $\tau_{11}^{c}(j)$, $j\in\{0,1,2\}$, as outlined in (2), we start by considering the scenario where $Y_{gt}(d)=C_{gt}(d)$ almost surely, for $d\in\{0,1\}$. Although modeling $Y_{11}(1)$ is per se not essential in DiD settings, it becomes indispensable when addressing scenarios involving misreporting. Therefore, to simplify the notation and facilitate the transition to the model with misreporting, we also postulate an analogous model structure for $Y_{11}(1)$ as for $Y_{11}(0)$ without misreporting. Specifically, drawing on the model for $Y_{11}^{\ast}(0)$ in (13), we define the latent consumption in the presence of treatment as:
$Y_{11}^{\ast}(1)=\overline{\mathbf{X}}_{11}^{\prime}\overline{\boldsymbol{\eta}}_{11}^{(1)}+\lambda_{11}^{(1)}\varepsilon_{11}^{(1)}.$ (18)
where we introduced the superscript (1) to distinguish the parameters of the potential outcome under treatment from those under control. From this point forward, we will accordingly add the superscript $(0)$ to parameters in the absence of treatment. Given the similarity in the structure of $Y_{11}^{\ast}(1)$, we adopt assumptions similar to those for $Y_{11}^{\ast}(0)$ in Section 4.1 above:
Assumption 8.
$\varepsilon_{11}^{(1)}$ is strictly stationary within each group with zero median and unit variance, independent of $\mathbf{X}_{11}$, and has an unknown, continuous, and strictly monotonic CDF $F_{\varepsilon_{11}^{(1)}}$. Besides, $\eta_{1,11}^{(1)}\neq 0$.
We readily obtain the following proposition for identification of the conditional DTTs for consumption, in the absence of misreporting:
Proposition 5.
Suppose Assumptions 1-5 and 8 hold, and consider the model in (13) and (18). Also, suppose distributional parallel trends hold as in Definition 1 for $Y_{gt}^{*}(0)$. Then, $\tau_{11}^{c}(j\mid{\mathbf{x}})$ is identified for all $j\in\{0,1,2\}$ and $\mathbf{x}\in\mathcal{X}_{11}$ as:
$\tau_{11}^{c}(j\mid{\mathbf{x}})=$ $\left(F_{\varepsilon_{11}^{(1)}}\left(\frac{\kappa_{j+1}-\overline{\mathbf{x}}^{\prime}\overline{\boldsymbol{\eta}}^{(1)}_{11}}{\lambda^{(1)}_{11}}\right)-F_{\varepsilon_{11}^{(1)}}\left(\frac{\kappa_{j}-\overline{\mathbf{x}}^{\prime}\overline{\boldsymbol{\eta}}^{(1)}_{11}}{\lambda^{(1)}_{11}}\right)\right)$
$-\left(F_{\varepsilon_{11}^{(0)}}\left(\frac{\kappa_{j+1}-\overline{\mathbf{x}}^{\prime}\overline{\boldsymbol{\eta}}^{(0)}_{11}}{\lambda_{11}^{(0)}}\right)-F_{\varepsilon_{11}^{(0)}}\left(\frac{\kappa_{j}-\overline{\mathbf{x}}^{\prime}\overline{\boldsymbol{\eta}}_{11}^{(0)}}{\lambda_{11}^{(0)}}\right)\right).$
Next, we turn to the case where the outcome is subject to underreporting. Drawing on the model for $R_{11}^{\ast}(0)$ in (14), We define:
$R_{11}^{\ast}(1)=\overline{\mathbf{Z}}_{11}^{\prime}\overline{\boldsymbol{\pi}}_{11}^{(1)}+\zeta_{11}^{(1)}\nu_{11}^{(1)}.$ (19)
We adopt assumptions similar to those for $Y_{11}^{\ast}(0)$ in Section 4.2 above:
Assumption 9.
$-\varepsilon_{11}^{(1)}$ and $-\nu_{11}^{(1)}$ are strictly stationary within each group with zero median and unit variance, jointly independent of $(\mathbf{X}_{11}^{\prime},\mathbf{Z}_{11}^{\prime})^{\prime}$, and have unknown, real analytic and strictly monotonic CDFs $F_{-\varepsilon_{11}^{(1)}}$ and $F_{-\nu_{11}^{(1)}}$, respectively. Besides, $\eta_{1,11}^{(1)}\neq 0$ and $\pi_{1,11}^{(1)}\neq 0$ and there exists some Cartesian product of two open intervals $I_{1}\times I_{2}\subseteq\mathbb{R}\times\mathbb{R}$ such that
$\operatorname{Pr}\left(\left(X_{1,11},Z_{1,11}\right)\in I_{1}\times I_{2}\mid\overline{\mathbf{W}}_{11}\right)>0$
with probability one, where $\overline{\mathbf{W}}_{11}$ is defined as in Assumption 7. Also, the copula function $\overline{C}_{11}^{(1)}(\cdot,\cdot)$ defined below is strictly monotonic and real analytic in each of its coordinates:
$\Pr\left(C_{11}(1)\geq j\mid\mathbf{X}_{11},\mathbf{Z}_{11}\right)$ $=\overline{C}_{11}^{(1)}\left(F_{-{\varepsilon_{11}^{(1)}}}\left(\frac{\overline{\mathbf{X}}_{11}^{\prime}\overline{\boldsymbol{\eta}}_{11}^{(1)}-\kappa_{j}}{\lambda_{11}^{(1)}}\right),F_{-{\nu}_{11}^{(1)}}\left(\frac{\overline{\mathbf{Z}}_{11}^{\prime}\overline{\boldsymbol{\pi}}_{11}^{(1)}-\iota_{j}}{\zeta_{11}^{(1)}}\right)\right).$
We have the following result for identification of the conditional DTTs for consumption and reporting, respectively, in the presence of misreporting:
Proposition 6.
Suppose Assumptions 6-7 and 9 hold, and consider the model in (13)-(14) and (18)-(19). Also, suppose distributional parallel trends hold as in Definition 1 for $Y_{gt}^{*}(0)$ as well as $R_{gt}^{*}(0)$. Then, $\tau_{11}^{c}(j)$ and $\tau_{11}^{r}(j)$ are identified for all $j\in\{0,1,2\}$, $\mathbf{x}\in\mathcal{X}_{11}$ and $\mathbf{z}\in\mathcal{Z}_{11}$ as:
$\tau_{11}^{c}(j\mid{\mathbf{x}})=$ $\left(F_{-\varepsilon_{11}^{(1)}}\left(\frac{\overline{\mathbf{x}}^{\prime}\overline{\boldsymbol{\eta}}^{(1)}_{11}-\kappa_{j+1}}{\lambda^{(1)}_{11}}\right)-F_{-\varepsilon_{11}^{(1)}}\left(\frac{\overline{\mathbf{x}}^{\prime}\overline{\boldsymbol{\eta}}^{(1)}_{11}-\kappa_{j}}{\lambda^{(1)}_{11}}\right)\right)$
$-\left(F_{-\varepsilon_{11}^{(0)}}\left(\frac{\overline{\mathbf{x}}^{\prime}\overline{\boldsymbol{\eta}}^{(0)}_{11}-\kappa_{j+1}}{\lambda_{11}^{(0)}}\right)-F_{-\varepsilon_{11}^{(0)}}\left(\frac{\overline{\mathbf{x}}^{\prime}\overline{\boldsymbol{\eta}}^{(0)}_{11}-\kappa_{j+1}}{\lambda_{11}^{(0)}}\right)\right).$
$\tau_{11}^{r}(j\mid{\mathbf{z}})=$ $\left(F_{-\nu_{11}^{(1)}}\left(\frac{\overline{\mathbf{z}}^{\prime}\overline{\boldsymbol{\pi}}^{(1)}_{11}-\iota_{j+1}}{\zeta^{(1)}_{11}}\right)-F_{-\nu_{11}^{(1)}}\left(\frac{\overline{\mathbf{z}}^{\prime}\overline{\boldsymbol{\pi}}^{(1)}_{11}-\iota_{j}}{\zeta^{(1)}_{11}}\right)\right)$
$-\left(F_{-\nu_{11}^{(0)}}\left(\frac{\overline{\mathbf{z}}^{\prime}\overline{\boldsymbol{\pi}}^{(0)}_{11}-\iota_{j+1}}{\zeta_{11}^{(0)}}\right)-F_{-\nu_{11}^{(0)}}\left(\frac{\overline{\mathbf{z}}^{\prime}\overline{\boldsymbol{\pi}}^{(0)}_{11}-\iota_{j}}{\zeta_{11}^{(0)}}\right)\right).$
Interest may also extend to unconditional DTTs, which are straightforward to identify and to construct. Additionally, introducing stability conditions to the copula function can help identify the counterfactual copula function. This aspect is relevant for understanding the impact of treatment on the dependence structure between individual consumption decisions and reporting intentions. In Section LABEL:sec:copuladecomp of the supplementary material, we present a decomposition of another possible causal quantity. This effect is characterized by the difference between potential self-reported outcome distributions under treatment and control, where the counterfactual copula function is involved.
7 Estimation
In this section, we discuss the estimation of the DTTs $\tau_{11}^{c}(j)$. For each observational unit $i$, let $G_{i}$ and $T_{i}$ denote binary indicators of group membership (with $G_{i}=1$ indicating being in the treated group) and time period (with $T_{i}=1$ indicating being after the treatment), respectively. Following Sant’Anna and Zhao (2020), we write the observed outcomes and the covariates as $(C_{it},\mathbf{X}_{it}^{\prime},\mathbf{Z}_{it}^{\prime})^{\prime}$, where the subscript $t$ denotes the time period. This allows to accommodate various sampling frameworks, including repeated cross-section sampling, where observations are drawn conditional on a specific group $G_{i}=g$ and time period $T_{i}=t$, and panel sampling that draws observation pairs conditional on $G_{i}=g$, see Assumption 10 below. Consequently, we define $C_{i}=C_{i1}T_{i}+C_{i0}(1-T_{i})$ and $Y_{i}=Y_{i1}T_{i}+Y_{i0}(1-T_{i})$, with regressors defined in a similar manner. It is important to recall that in the absence of “false zeros”, we observe $Y_{it}=C_{it}$ almost surely.
7.1 Estimation Without Misreporting
We start with the estimation of the counterfactual distribution in the absence of misreporting, as described in Section 4, focusing initially on a parametric model without covariates detailed in Section 4.1. Recalling that $\kappa_{0}=-\infty$, $\kappa_{1}=0$, $\kappa_{2}=1$, $\kappa_{3}=+\infty$, we consider a sample $\{C_{i},G_{i},T_{i}\}_{i=1}^{n}$ with the properties specified in Assumption 10 below. Also, suppose Assumptions 1 and 8 hold and consider the models in (12) and (18) (no covariates for the latter), where the distributions of $\varepsilon_{gt}^{(d)}$ are known for all $d=g\cdot t$, $g\in\{0,1\}$ and $t\in\{0,1\}$. To avoid notational complexity, we assume that the error term distributions are invariant across groups, time and treatment statuses, i.e., $F_{\varepsilon_{gt}^{(d)}}=F_{\varepsilon}$. We define $\boldsymbol{\theta}^{(d)}_{gt}:=\left(\eta^{(d)}_{gt},\lambda^{(d)}_{gt}\right)^{\prime}$ and aggregate these into $\boldsymbol{\theta}:=\left(\boldsymbol{\theta}^{(0)^{\prime}}_{00},\boldsymbol{\theta}^{(0)^{\prime}}_{01},\boldsymbol{\theta}^{(0)^{\prime}}_{10},\boldsymbol{\theta}^{(1)^{\prime}}_{11}\right)^{\prime}$. The ML estimator of $\boldsymbol{\theta}$ is formulated as follows:
$\widehat{\boldsymbol{\theta}}:=$ $\arg\min_{\boldsymbol{\theta}\in\mathcal{B}}\sum_{i=1}^{n}\sum_{j=0}^{2}\mathbb{I}\left[Y_{i}=j\right]\left[\sum_{(g,t)\neq(1,1)}\mathbb{I}\left[G_{i}=g,T_{i}=t\right]\ln\left(F_{\varepsilon}\left(\frac{\kappa_{j+1}-\eta_{gt}^{(0)}}{\lambda_{gt}^{(0)}}\right)-F_{\varepsilon}\left(\frac{\kappa_{j}-\eta_{gt}^{(0)}}{\lambda_{gt}^{(0)}}\right)\right)\right.$
$+\left.\mathbb{I}\left[G_{i}=1,T_{i}=1\right]\ln\left(F_{\varepsilon}\left(\frac{\kappa_{j+1}-\eta^{(1)}_{11}}{\lambda^{(1)}_{11}}\right)-F_{\varepsilon}\left(\frac{\kappa_{j}-\eta^{(1)}_{11}}{\lambda^{(1)}_{11}}\right)\right)\right].$ (20)
where $\mathcal{B}$ denotes the parameter space, $\mathbb{I}\left[A\right]$ is the indicator function that equals 1 when $A$ is true and 0 otherwise.
This step allows for the construction of DTT estimates, $\tau_{11}^{c}(j)$ for $j\in\{0,1,2\}$:
$\widehat{\tau}^{c}_{11}(j):=\left[F_{\varepsilon}\left(\frac{\kappa_{j+1}-\widehat{\eta}^{(1)}_{11}}{\widehat{\lambda}^{(1)}_{11}}\right)-F_{\varepsilon}\left(\frac{\kappa_{j}-\widehat{\eta}^{(1)}_{11}}{\widehat{\lambda}^{(1)}_{11}}\right)\right]-\left[F_{\varepsilon}\left(\frac{\kappa_{j+1}-\widehat{\eta}^{(0)}_{11}}{\widehat{\lambda}^{(0)}_{11}}\right)-F_{\varepsilon}\left(\frac{\kappa_{j}-\widehat{\eta}^{(0)}_{11}}{\widehat{\lambda}^{(0)}_{11}}\right)\right]$ (21)
where
$\widehat{\eta}^{(0)}_{11}:=\widehat{\eta}^{(0)}_{01}+\frac{\widehat{\eta}^{(0)}_{10}-\widehat{\eta}^{(0)}_{00}}{\widehat{\lambda}^{(0)}_{00}/\widehat{\lambda}^{(0)}_{01}}\quad\text{and}\quad\widehat{\lambda}^{(0)}_{11}:=\frac{\widehat{\lambda}^{(0)}_{01}\widehat{\lambda}^{(0)}_{10}}{\widehat{\lambda}^{(0)}_{00}}.$
For the inference on the treatment parameters $\tau_{11}^{c}(j)$, it is crucial to derive the asymptotic distribution of the ML estimator. To do so, we impose the following assumptions:
Assumption 10.
It holds that either (i) conditional on $G_{i}=g$ and $T_{i}=t$ with $g,t\in\{0,1\}$, $\{C_{i}\}_{i=1}^{n}$ is an i.i.d. random sample from the corresponding subpopulation or (ii) conditional on $G_{i}=g$ with $g\in\{0,1\}$, $\{C_{i0},C_{i1}\}_{i=1}^{n}$ is an i.i.d. random sample from the corresponding subpopulation, either $g=0$ or $g=1$.
Assumption 11.
The ML estimator of $\boldsymbol{\theta}$ defined in (7.1) satisfies a Central Limit Theorem:
$\sqrt{n}\left(\widehat{\boldsymbol{\theta}}-\boldsymbol{\theta}\right)\stackrel{{\scriptstyle d}}{{\rightarrow}}N\left(\boldsymbol{0},\boldsymbol{\Sigma}\right),$
where $\boldsymbol{\Sigma}$ is a positive definite matrix.
The first part (i) of Assumption 10 addresses the case where repeated cross-section data is available, while part (ii) covers the case of panel data. While Assumption 10 only accommodates the “conditional” sampling scheme where group membership and time are fixed, the assumption could readily be extended to allow for other schemes where $G_{i}$ and $T_{i}$ are random (see Sant’Anna and Zhao, 2020). Assumption 11 on the other hand is a high-level condition, valid under the stated model assumptions as well as certain primitive conditions on the Hessian and the population shares of each group. In particular, it implies that $n_{gt}/n\rightarrow K$ (with $0<K<\infty$) as $n\rightarrow\infty$ for each $g\in\{0,1\}$, $t\in\{0,1\}$, where $n_{gt}$ denotes the number of observations in the group $g$ and time period $t$.
Proposition 7.
Suppose Assumptions 10-11 hold. Then, for each $j\in\{0,1,2\}$:
$\sqrt{n}\left(\widehat{\tau}^{c}_{11}(j)-\tau_{11}^{c}(j)\right)\stackrel{{\scriptstyle d}}{{\rightarrow}}N\left(\boldsymbol{0},\frac{\partial\tau_{11}^{c}(j)}{\partial{{\boldsymbol{\theta}}}}^{\prime}\boldsymbol{\Sigma}\frac{\partial\tau_{11}^{c}(j)}{\partial{{\boldsymbol{\theta}}}}\right),$
where the definition of $\frac{\partial\tau_{11}^{c}(j)}{\partial{{\boldsymbol{\theta}}}}$ is given in Section LABEL:sec:derivatives of the supplementary material.
The components of the asymptotic variance can be readily estimated. Specifically, the asymptotic variance matrix $\boldsymbol{\Sigma}$ can be replaced by a standard heteroskedasticity robust estimator, denoted as $\widehat{\boldsymbol{\Sigma}}_{R}$, while the derivatives $\frac{\partial\tau_{11}^{c}(j)}{\partial{{\boldsymbol{\theta}}}}$ can be substituted with a consistent estimator $\widehat{\partial\tau_{11}^{c}(j)/\partial{{\boldsymbol{\theta}}}}$.
Remark 2: Note that the identification argument for the semiparametric model from Subsection 4.2 can in principle be used to construct a plug-in estimator of the DTTs. Here, we provide an informal outline of the steps. Section LABEL:sec:proofEstimation of the supplementary material considers the estimation procedure for the DTTs more formally. Introducing notations for the rescaled and recentered quantities, let ${\boldsymbol{\eta}}_{gt}^{(d)^{*}}:=\frac{{\boldsymbol{\eta}}_{gt}^{(d)}}{\left\|{\boldsymbol{\eta}}_{gt}^{(d)}\right\|}$, $\kappa_{j,gt}^{(d)^{*}}:=\frac{\kappa_{j}}{\left\|{\boldsymbol{\eta}}_{gt}^{(d)}\right\|}$, $\widetilde{\varepsilon}_{gt}^{(d)}:=\eta_{c,gt}^{(d)}+\lambda_{gt}^{(d)}\varepsilon_{gt}^{(d)}$ and $\widetilde{\varepsilon}_{gt}^{(d)^{*}}:=\frac{\widetilde{\varepsilon}_{gt}^{(d)}}{\left\|{\boldsymbol{\eta}}_{gt}^{(d)}\right\|}$ for all $d=g\cdot t$, $g\in\{0,1\}$ and $t\in\{0,1\}$. For simplicity, we will assume that $\eta_{c,gt}^{(d)}=\eta_{c}$ and $\lambda_{gt}^{(d)}=\lambda$ for all $(g,t)\in\{0,1\}$ and $d\in\{0,1\}$, and comment on the case where these parameters differ at the end of this remark. Given the new notation, the model in (13) can thus be reformulated as:
$Y_{gt}(d)=$ $\sum_{j=0}^{2}j\cdot\mathbb{I}\left[\kappa_{j,gt}^{(d)^{*}}<{\mathbf{X}}_{gt}^{\prime}{\boldsymbol{\eta}}^{(d)^{*}}_{gt}+\widetilde{\varepsilon}^{(d)^{*}}_{gt}\leq\kappa_{j+1,gt}^{(d)^{*}}\right].$ (22)
The rescaled slope coefficient vector ${\boldsymbol{\eta}}^{(d)^{*}}_{gt}$ in (22) can be estimated using various rank estimators available in the literature (e.g., Sherman, 1993). Subsequently, the threshold parameter $\kappa_{2,gt}^{(d)^{*}}=1/\|{\boldsymbol{\eta}}^{(d)}_{gt}\|$ can be estimated using the $\sqrt{n}_{gt}$-consistent threshold estimator for semiparametric ordered choice models proposed by Klein and Sherman (2002). This yields an estimator for the scaling factor as
$\widehat{\|\boldsymbol{\eta}^{(d)}_{gt}\|}:=1/\widehat{\kappa_{2,gt}^{(d)^{*}}}.$
Consequently, the estimator for the unscaled parameter vector becomes:
$\widehat{\boldsymbol{\eta}}^{(d)}_{gt}:=\widehat{\boldsymbol{\eta}}_{gt}^{(d)^{*}}\widehat{\|\boldsymbol{\eta}^{(d)}_{gt}\|}.$
After estimating all parameters, we recover the distributions of $\widetilde{\varepsilon}_{gt}^{(d)}$ using, e.g., a standard nonparametric kernel estimator:
$\widehat{F}_{\widetilde{\varepsilon}_{gt}^{(d)}}(v):=\widehat{\mathrm{E}}\left[\mathbb{I}\left[Y_{i}\leq 0\right]\mid\mathbf{X}_{i}^{\prime}\widehat{\boldsymbol{\eta}}_{gt}^{(d)}=-v,G_{i}=g,T_{i}=t\right].$
With estimates for all groups and time periods at hand, and under strict stationarity for the counterfactual distribution from Assumption 1, we can construct the estimated conditional DTTs for $j\in\{0,1,2\}$ and $\mathbf{x}\in\mathcal{X}_{11}$ as outlined in Proposition 6:
$\widehat{\tau}^{c}_{11}(j\mid\mathbf{x})$
::=:= $\mathbb{I}\left[(\kappa_{j+1}-\mathbf{x}^{\prime}\widehat{\boldsymbol{\eta}}^{(1)}_{11}),(\kappa_{j}-\mathbf{x}^{\prime}\widehat{\boldsymbol{\eta}}^{(1)}_{11})\in\Psi_{11}^{(1)};\;(\kappa_{j+1}-\mathbf{x}^{\prime}\widehat{\boldsymbol{\eta}}^{(0)}_{11}),(\kappa_{j}-\mathbf{x}^{\prime}\widehat{\boldsymbol{\eta}}^{(0)}_{11})\in\Psi_{11}^{(0)}\right]$
×$\times\left(\left[\widehat{F}_{{\widetilde{\varepsilon}}_{11}^{(1)}}\left(\kappa_{j+1}-\mathbf{x}^{\prime}\widehat{\boldsymbol{\eta}}^{(1)}_{11}\right)-\widehat{F}_{{\widetilde{\varepsilon}}_{11}^{(1)}}\left(\kappa_{j}-\mathbf{x}^{\prime}\widehat{\boldsymbol{\eta}}^{(1)}_{11}\right)\right]-\left[\widehat{F}_{{\widetilde{\varepsilon}}_{11}^{(0)}}\left(\kappa_{j+1}-\mathbf{x}^{\prime}\widehat{\boldsymbol{\eta}}^{(0)}_{11}\right)-\widehat{F}_{{\widetilde{\varepsilon}}_{11}^{(0)}}\left(\kappa_{j}-\mathbf{x}^{\prime}\widehat{\boldsymbol{\eta}}^{(0)}_{11}\right)\right]\right),$
where $\widehat{\boldsymbol{\eta}}_{11}^{(0)}:=\widehat{\boldsymbol{\eta}}_{01}^{(0)}+\widehat{\boldsymbol{\eta}}_{10}^{(0)}-\widehat{\boldsymbol{\eta}}_{00}^{(0)}$ and $\widehat{F}_{{\widetilde{\varepsilon}}_{11}^{(0)}}(v):=\widehat{F}_{{\widetilde{\varepsilon}}_{10}^{(0)}}(v)$ for all $v\in\mathbb{R}$, while $\Psi_{11}^{(d)}$, $d\in\{0,1\}$, denote suitable trimming sets. Under regularity conditions on the different estimators, the above procedure yields consistent estimates of the DTTs of interest, see Section LABEL:sec:proofEstimation of the supplementary material for details. Finally, when the median (or mean) as well as the variance parameters actually differ, a slightly modified procedure can be adopted. In particular, we can construct estimators for the median (or mean) parameter $\eta_{c,gt}^{(d)}$ and the scaling parameter $\lambda_{gt}^{(d)}$ as follows:
[…]
for all $d=g\cdot t$, $g\in\{0,1\}$ and $t\in\{0,1\}$, where $\Psi^{(d)}_{n,gt}$ now denotes a trimming set on which the index density is strictly bounded away from zero and that slowly expands with the sample size (cf. Hansen, 2008). This leads to an alternative estimator for the unscaled error term distribution:
$\widehat{F}_{{\varepsilon}_{gt}^{(d)}}(v):=\widehat{F}_{\widetilde{\varepsilon}_{gt}^{(d)}}\left(\widehat{\eta}_{c,gt}^{(d)}+\widehat{\lambda}_{gt}^{(d)}v\right),$
for all $\left(\widehat{\eta}_{c,gt}^{(d)}+\widehat{\lambda}_{gt}^{(d)}v\right)\in\Psi^{(d)}_{n,gt}$, though deriving the asymptotic distribution in this case is more involved and beyond the scope of the paper.
7.2 Estimation With Underreported Outcomes
In this subsection, we turn to the estimation of the DTTs when outcomes are subject to underreporting, as elaborated in Section 5. Acknowledging the non-constructive nature of the identification argument presented therein, our subsequent exposition assumes that $\overline{C}_{gt}^{(d)}(\cdot,\cdot)$ is a parametric copula, specifically $\overline{C}_{gt}^{(d)}\left(\cdot,\cdot;\rho_{gt}^{(d)}\right)$, parameterized by a dependence parameter $\rho_{gt}^{(d)}$ for all $d=g\cdot t$, $g\in\{0,1\}$ and $t\in\{0,1\}$. Additionally, we postulate that the distributions of $-\varepsilon_{gt}^{(d)}$ and $-\nu_{gt}^{(d)}$ are known. For simplicity, we assume invariance across groups, time and treatment statuses in terms of copula functions and marginal distributions, i.e., $F_{\varepsilon_{gt}^{(d)}}=F_{\varepsilon}$, $F_{\nu_{gt}^{(d)}}=F_{\nu}$, and $\overline{C}_{gt}^{(d)}=\overline{C}$. This assumption can be relaxed straightforwardly if needed.
We define $\widetilde{\boldsymbol{\theta}}^{(d)}_{gt}:=\left(\overline{\boldsymbol{\eta}}^{(d)\prime}_{gt},\lambda^{(d)}_{gt},\overline{\boldsymbol{\pi}}^{(d)\prime}_{gt},\zeta^{(d)}_{gt},\rho^{(d)}_{gt}\right)^{\prime}$ for all $d=g\cdot t$, $g\in\{0,1\}$ and $t\in\{0,1\},$ and aggregate these into $\widetilde{\boldsymbol{\theta}}:=\left(\widetilde{\boldsymbol{\theta}}^{(0)\prime}_{00},\widetilde{\boldsymbol{\theta}}^{(0)\prime}_{01},\widetilde{\boldsymbol{\theta}}^{(0)\prime}_{10},\widetilde{\boldsymbol{\theta}}^{(1)\prime}_{11}\right)^{\prime}$. With this, the estimator in (7.1) is modified to:
$\widehat{\widetilde{\boldsymbol{\theta}}}:=$ $\arg\min_{\widetilde{\boldsymbol{\theta}}\in\mathcal{B}}\sum_{i=1}^{n}\sum_{j=0}^{2}\mathbb{I}\left[C_{i}=j\right]\left\{\sum_{(g,t)\neq(1,1)}\mathbb{I}\left[G_{i}=g,T_{i}=t\right]\times\right.$
$\left.\ln\left[\overline{C}\left(F_{-\varepsilon}\left(\frac{\overline{\mathbf{X}}_{i}^{\prime}\overline{\boldsymbol{\eta}}_{gt}^{(0)}-\kappa_{j+1}}{\lambda_{gt}^{(0)}}\right),F_{-\nu}\left(\frac{\overline{\mathbf{Z}}_{i}^{\prime}\overline{\pi}_{gt}^{(0)}-\iota_{j+1}}{\zeta_{gt}^{(0)}}\right);\rho_{gt}^{(0)}\right)-\overline{C}\left(F_{-\varepsilon}\left(\frac{\overline{\mathbf{X}}_{i}^{\prime}\overline{\boldsymbol{\eta}}_{gt}^{(0)}-\kappa_{j}}{\lambda_{gt}^{(0)}}\right),F_{-\nu}\left(\frac{\overline{\mathbf{Z}}_{i}^{\prime}\overline{\boldsymbol{\pi}}_{gt}^{(0)}-\iota_{j}}{\zeta_{gt}^{(0)}}\right);\rho_{gt}^{(0)}\right)\right]\right.$
$\left.+\mathbb{I}\left[G_{i}=1,T_{i}=1\right]\ln\left[\overline{C}\left(F_{-\varepsilon}\left(\frac{\overline{\mathbf{X}}_{i}^{\prime}\overline{\boldsymbol{\eta}}_{11}^{(1)}-\kappa_{j+1}}{\lambda_{11}^{(1)}}\right),F_{-\nu}\left(\frac{\overline{\mathbf{Z}}_{i}^{\prime}\overline{\pi}_{11}^{(1)}-\iota_{j+1}}{\zeta_{11}^{(1)}}\right);\rho_{11}^{(1)}\right)\right.\right.$ (23)
$\left.\left.-\overline{C}\left(F_{-\varepsilon}\left(\frac{\overline{\mathbf{X}}_{i}^{\prime}\overline{\boldsymbol{\eta}}_{11}^{(1)}-\kappa_{j}}{\lambda_{11}^{(1)}}\right),F_{-\nu}\left(\frac{\overline{\mathbf{Z}}_{i}^{\prime}\overline{\boldsymbol{\pi}}_{11}^{(1)}-\iota_{j}}{\zeta_{11}^{(1)}}\right);\rho_{11}^{(1)}\right)\right]\right\}$
This estimator is conceptually similar to the first step of the estimation procedure described in Nguimkeu et al. (2019), which deals with the estimation of linear models with a partially observed, endogenous regressor within a group and a time period. Using the estimates, we construct the estimated conditional DTTs as outlined in Proposition 6:
$\widehat{\tau}^{c}_{11}(j\mid\mathbf{x})$ $:=\left[F_{-\varepsilon}\left(\frac{\overline{\mathbf{x}}^{\prime}\widehat{\overline{\boldsymbol{\eta}}}^{(1)}_{11}-\kappa_{j+1}}{\widehat{\lambda}^{(1)}_{11}}\right)-F_{-\varepsilon}\left(\frac{\overline{\mathbf{x}}^{\prime}\widehat{\overline{\boldsymbol{\eta}}}^{(1)}_{11}-\kappa_{j}}{\widehat{\lambda}^{(1)}_{11}}\right)\right]-\left[F_{-\varepsilon}\left(\frac{\overline{\mathbf{x}}^{\prime}\widehat{\overline{\boldsymbol{\eta}}}^{(0)}_{11}-\kappa_{j+1}}{\widehat{\lambda}^{(0)}_{11}}\right)-F_{-\varepsilon}\left(\frac{\overline{\mathbf{x}}^{\prime}\widehat{\overline{\boldsymbol{\eta}}}^{(0)}_{11}-\kappa_{j}}{\widehat{\lambda}^{(0)}_{11}}\right)\right]$
$\widehat{\tau}^{r}_{11}(j\mid\mathbf{z})$ $:=\left[F_{-\nu}\left(\frac{\overline{\mathbf{z}}^{\prime}\widehat{\overline{\boldsymbol{\pi}}}^{(1)}_{11}-\iota_{j+1}}{\widehat{\zeta}^{(1)}_{11}}\right)-F_{-\nu}\left(\frac{\overline{\mathbf{z}}^{\prime}\widehat{\overline{\boldsymbol{\pi}}}^{(1)}_{11}-\iota_{j}}{\widehat{\zeta}^{(1)}_{11}}\right)\right]-\left[F_{-\nu}\left(\frac{\overline{\mathbf{z}}^{\prime}\widehat{\overline{\boldsymbol{\pi}}}^{(0)}_{11}-\iota_{j+1}}{\widehat{\zeta}^{(0)}_{11}}\right)-F_{-\nu}\left(\frac{\overline{\mathbf{z}}^{\prime}\widehat{\overline{\boldsymbol{\pi}}}^{(0)}_{11}-\iota_{j}}{\widehat{\zeta}^{(0)}_{11}}\right)\right]$
for $j\in\{0,1,2\}$, $\mathbf{x}\in\mathcal{X}_{11}$ and $\mathbf{z}\in\mathcal{Z}_{11}$, where
$\widehat{\overline{\boldsymbol{\eta}}}_{11}^{(0)}:=\widehat{\overline{\boldsymbol{\eta}}}_{01}^{(0)}+\frac{\widehat{\overline{\boldsymbol{\eta}}}_{10}^{(0)}-\widehat{\overline{\boldsymbol{\eta}}}_{00}^{(0)}}{\widehat{\lambda}_{00}^{(0)}/\widehat{\lambda}_{01}^{(0)}},\quad\widehat{\lambda}_{11}^{(0)}:=\frac{\widehat{\lambda}_{01}^{(0)}\widehat{\lambda}_{10}^{(0)}}{\widehat{\lambda}_{00}^{(0)}},\quad\widehat{\overline{\boldsymbol{\pi}}}_{11}^{(0)}:=\widehat{\overline{\boldsymbol{\pi}}}_{01}^{(0)}+\frac{\widehat{\overline{\boldsymbol{\pi}}}_{10}^{(0)}-\widehat{\overline{\boldsymbol{\pi}}}_{00}^{(0)}}{\widehat{\zeta}_{00}^{(0)}/\widehat{\zeta}_{01}^{(0)}}\quad\text{and}\quad\widehat{\zeta}_{11}^{(0)}:=\frac{\widehat{\zeta}_{01}^{(0)}\widehat{\zeta}_{10}^{(0)}}{\widehat{\zeta}_{00}^{(0)}}.$
We also explore the construction of “marginalized” DTTs for the treated group after treatment (see, e.g., Chernozhukov et al., 2013):
[…]
Here, $\widehat{F}_{\mathbf{X}_{11}}(\cdot)$ represents an estimator of the covariate distribution. For instance:
$\widehat{\tau}^{c}_{11}(j):=$ $\frac{1}{n_{11}}\sum_{i=1}^{n_{11}}\mathbb{I}\left[G_{i}=1,T_{i}=1\right]\mathbb{I}\left[\mathbf{X}_{i}\in\mathcal{X}_{11}\right]\left(F_{-\varepsilon}\left(\frac{\overline{\mathbf{X}}_{i}^{\prime}\widehat{\overline{\boldsymbol{\eta}}}^{(1)}_{11}-\kappa_{j+1}}{\widehat{\lambda}^{(1)}_{11}}\right)-F_{-\varepsilon}\left(\frac{\overline{\mathbf{X}}_{i}^{\prime}\widehat{\overline{\boldsymbol{\eta}}}^{(1)}_{11}-\kappa_{j}}{\widehat{\lambda}^{(1)}_{11}}\right)\right)$
$-\frac{1}{n_{11}}\sum_{i=1}^{n_{11}}\mathbb{I}\left[G_{i}=1,T_{i}=1\right]\mathbb{I}\left[\mathbf{X}_{i}\in\mathcal{X}_{11}\right]\left(F_{-\varepsilon}\left(\frac{\overline{\mathbf{X}}_{i}^{\prime}\widehat{\overline{\boldsymbol{\eta}}}^{(0)}_{11}-\kappa_{j+1}}{\widehat{\lambda}^{(0)}_{11}}\right)-F_{-\varepsilon}\left(\frac{\overline{\mathbf{X}}_{i}^{\prime}\widehat{\overline{\boldsymbol{\eta}}}^{(0)}_{11}-\kappa_{j}}{\widehat{\lambda}^{(0)}_{11}}\right)\right).$
The asymptotic distribution of these DTT variants can be derived similarly to the approach in Subsection 7.1, with standard errors constructed accordingly.
Remark 3: Test statistics for the hypotheses $H_{0}^{C}$ vs. $H_{A}^{C}$ or $H_{0}^{R}$ vs. $H_{A}^{R}$, as discussed in Remark 1, may be readily constructed using standard Cramer-von Mises measures of distance. Specifically, for testing $H_{0}^{C}$ vs. $H_{A}^{C}$, the following statistic can be employed for some $(g.t)\neq(1,1)$:
$\sum_{i=1}^{n}\sum_{j=1}^{2}\mathbb{I}\left[G_{i}=g,T_{i}=t\right]\mathbb{I}\left[(\mathbf{X}_{i},\mathbf{Z}_{i})\in\mathcal{X}_{gt}\times\mathcal{Z}_{gt}\right]$
×$\times\left[\overline{C}\left(F_{-{\varepsilon}}\left(\frac{\overline{\mathbf{X}}_{i}^{\prime}\widehat{\overline{\boldsymbol{\eta}}}_{gt}^{(0)}-\kappa_{j}}{\widehat{\lambda}_{gt}^{(0)}}\right),F_{-{\nu}}\left(\frac{\overline{\mathbf{Z}}_{i}^{\prime}\widehat{\overline{\boldsymbol{\pi}}}_{gt}^{(0)}-\iota_{j}}{\widehat{\zeta}_{gt}^{(0)}}\right);\widehat{\rho}_{gt}^{(0)}\right)-\widehat{M}\left(j;\mathbf{X}_{i}^{\prime}\widehat{\widetilde{\boldsymbol{\eta}}}_{gt},\mathbf{Z}_{i}^{\prime}\widehat{\widetilde{\boldsymbol{\pi}}}_{gt}\right)\right]^{2},$
where $\left(\widehat{\overline{\boldsymbol{\eta}}}_{gt}^{(0)^{\prime}},\widehat{\lambda}_{gt}^{(0)},\widehat{\overline{\boldsymbol{\pi}}}_{gt}^{(0)^{\prime}},\widehat{\zeta}^{(0)}_{gt},\widehat{\rho}_{gt}^{(0)}\right)$ are ML estimators discussed before, $\left(\widehat{\widetilde{\boldsymbol{\eta}}}_{gt}^{\prime},\widehat{\widetilde{\boldsymbol{\pi}}}_{gt}^{\prime}\right)$ denote multi-index estimators (Donkers and Schafgans, 2008; Ahn et al., 2018), and $\widehat{M}(j;\cdot,\cdot)$ is a suitable nonparametric estimator. A similar approach can be formulated for testing $H_{0}^{R}$ vs. $H_{A}^{R}$. The limiting distributions of both tests, while beyond the scope of the paper, are standard, and suitable resampling procedures can be used to construct asymptotically valid critical values.
8 Simulation
In this section, we present Monte Carlo simulation evidence to investigate the finite sample behavior of the ML estimator outlined in Subsection 7.2. Our focus is on assessing the finite sample bias and variance of the ML estimator, specifically in relation to the model parameters for marginal consumption and reporting distributions, as well as the dependence parameter in the copula function. The analysis spans various exclusion restriction scenarios, comparing cases with equal index and error variances to those where the index variance is less than the error variance.
The design of our observed outcome equation for simulation aligns with that in (1), but streamlined by not incorporating the treatment effects framework and simplifying the notation for the potential outcomes and the $gt$ subscript. Specifically, we model the consumption and reporting decisions as:
$C=\sum_{j=0}^{2}\mathbb{I}\left[\beta_{0}+\beta_{1}X+\beta_{2}W-\lambda\varepsilon>\kappa_{j}\right],$
$R=\sum_{j=0}^{2}\mathbb{I}\left[\pi_{0}+\pi_{1}Z+\pi_{2}W-\zeta\nu>\iota_{j}\right].$
Here, $W\sim\mathcal{N}(0,1)$ represents a common regressor, while $X$ and $Z$ are the “excluded” variables following a uniform distribution $\mathcal{U}\left(-\sqrt{3},\sqrt{3}\right)$ (detailed in Table 1). The joint distribution $(\varepsilon,\nu)^{\prime}$ is modeled using a Frank copula, denoted as $\overline{C}\left(F_{\varepsilon},F_{\nu};\rho\right)$, with the marginals being standard normal distributions. The Frank copula, representative of symmetric copulas, is defined as:
$\overline{C}(v_{1},v_{2};\rho)=-\frac{1}{\rho}\ln\left(1+\frac{(\exp(-\rho v_{1})-1)(\exp(-\rho v_{2})-1)}{\exp(-\rho)-1}\right)$
for $\rho\neq 0$. When $\rho=0$, it defaults to the independence copula.
For the simulation purposes, the dependence parameter $\rho$ is set such that the corresponding Spearman’s $\rho$ is 0.50.5-0.5- 0.5.111111Results for cases where the Spearman’s $\rho$ equals $-0.25$, yielding qualitatively similar outcomes, are available in Subsection LABEL:ssec:frank of the supplementary material. This approach is intended to facilitate comparability between dependence parameters across copulas, each of which can have distinct interpretations (see Han and Lee, 2019). Additionally, in Section LABEL:sec:simulations of the supplementary material, we extend our analysis to include simulations with the Clayton Copula, known for its asymmetric properties. These simulations yield findings that are qualitatively similar to those obtained using the Frank copula, thus highlighting the robustness of our study. We consider sample sizes $n=\{2000,5000,10000\}$, with each size undergoing 2000 Monte Carlo replications. To assess finite sample bias and variance, we focus on the Mean Bias (MB) and the Root Mean Square Error (RMSE).
Table 1 provides an overview of the various designs examined in our Monte Carlo study. We selected distributions and parameter values to ensure that the (unconditional) variances of the indices and the unobservables are equal to each other, with the exceptions of Cases 1-b and 2-b.
Table 1: Simulation Design Cases
Case No. of Exclusions Specification
0 No Exclusion $C=\sum_{j=1}^{2}\mathbb{I}\left[-1+2\cdot W-2\cdot\varepsilon>\kappa_{j}\right]$
$R=\sum_{j=1}^{2}\mathbb{I}\left[1.5-2\cdot W-2\cdot\nu>\iota_{j}\right]$
1-a 1 Exclusion $C=\sum_{j=1}^{2}\mathbb{I}\left[-1+2\cdot W-2\cdot\varepsilon>\kappa_{j}\right]$
$R=\sum_{j=1}^{2}\mathbb{I}\left[1.5-\sqrt{2}\cdot W-\sqrt{2}\cdot Z-2\cdot\nu>\iota_{j}\right]$
1-b 1 Exclusion $C=\sum_{j=1}^{2}\mathbb{I}\left[-1+\sqrt{2}\cdot W-2\cdot\varepsilon>\kappa_{j}\right]$
$R=\sum_{j=1}^{2}\mathbb{I}\left[1.5-W-Z-2\cdot\nu>\iota_{j}\right]$
2-a 2 Exclusions $C=\sum_{j=1}^{2}\mathbb{I}\left[-1+\sqrt{2}\cdot W+\sqrt{2}\cdot X-2\cdot\varepsilon>\kappa_{j}\right]$
$R=\sum_{j=1}^{2}\mathbb{I}\left[1.5-\sqrt{2}\cdot W-\sqrt{2}\cdot Z-2\cdot\nu>\iota_{j}\right]$
2-b 2 Exclusions $C=\sum_{j=1}^{2}\mathbb{I}\left[-1+W+X-2\cdot\varepsilon>\kappa_{j}\right]$
$R=\sum_{j=1}^{2}\mathbb{I}\left[1.5-W-Z-2\cdot\nu>\iota_{j}\right]$
Case 0 represents a scenario without exclusion restrictions, where identification is expected to fail. In contrast, Cases 1-a and 1-b focus on designs with a single exclusion restriction, differing in their (unconditional) index variances. Case 1-a considers equal (unconditional) index and error variances, whereas Case 1-b explores a setting with a smaller index variance relative to that of the error. Extending this framework, Cases 2-a and 2-b, akin to Cases 1-a and 1-b, respectively, incorporate two distinct exclusion restrictions in each equation.
Table 2: Frank Copula - Mean Bias
Case 0
Param. obs = 2000, $\rho$ = -0.5 obs = 5000, $\rho$ = -0.5 obs = 10000, $\rho$ = -0.5
$\beta_{0}$ $45.387$$10.271$$0.289$
$\beta_{1}$ $605.906$ $92.359$ $16.638$
$\pi_{0}$ $62.434$$12.121$$1.343$
$\pi_{1}$ $651.649$ $141.740$ $2.395$
$\lambda$ $719.680$ $79.700$ $16.190$
$\zeta$ $715.016$ $138.753$ $0.252$
$\rho$ $0.393$ $0.300$ $0.218$
Case 1-a
Param. obs = 2000, $\rho$ = -0.5 obs = 5000, $\rho$ = -0.5 obs = 10000, $\rho$ = -0.5
$\beta_{0}$$9.946$ $0.011$ $0.007$
$\beta_{1}$ $124.558$ $0.110$ $0.046$
$\pi_{0}$ $100.666$$0.229$$0.092$
$\pi_{1}$$77.552$ $0.110$ $0.040$
$\pi_{2}$$74.046$ $0.020$$0.002$
$\lambda$ $115.175$ $0.076$ $0.031$
$\zeta$ $91.241$$0.021$ $0.006$
$\rho$ $0.253$ $0.144$ $0.066$
Case 1-b
Param. obs = 2000, $\rho$ = -0.5 obs = 5000, $\rho$ = -0.5 obs = 10000, $\rho$ = -0.5
$\beta_{0}$$245.968$ $4.973$$0.011$
$\beta_{1}$ $1,902.489$ $15.519$ $0.128$
$\pi_{0}$ $252.689$$0.445$$0.341$
$\pi_{1}$$154.198$ $0.178$ $0.138$
$\pi_{2}$$149.381$ $0.034$ $0.040$
$\lambda$ $2,350.347$ $20.812$ $0.160$
$\zeta$ $270.014$$0.080$$0.078$
$\rho$ $0.364$ $0.273$ $0.190$
Case 2-a
Param. obs = 2000, $\rho$ = -0.5 obs = 5000, $\rho$ = -0.5 obs = 10000, $\rho$ = -0.5
$\beta_{0}$$0.035$$0.019$$0.005$
$\beta_{1}$ $0.114$ $0.046$ $0.016$
$\beta_{2}$ $0.118$ $0.046$ $0.017$
$\pi_{0}$ $13.961$ $0.043$ $0.028$
$\pi_{1}$$9.951$$0.040$$0.023$
$\pi_{2}$$11.006$$0.038$$0.024$
$\lambda$ $0.133$ $0.052$ $0.020$
$\zeta$ $13.324$ $0.043$ $0.029$
$\rho$ $0.026$ $0.007$ $0.001$
Case 2-b
Param. obs = 2000, $\rho$ = -0.5 obs = 5000, $\rho$ = -0.5 obs = 10000, $\rho$ = -0.5
$\beta_{0}$$280.768$$3.627$$0.014$
$\beta_{1}$ $603.093$ $9.119$ $0.055$
$\beta_{2}$ $603.862$ $8.056$ $0.050$
$\pi_{0}$ $238.591$ $0.152$ $0.063$
$\pi_{1}$$139.640$$0.088$$0.038$
$\pi_{2}$$147.751$$0.096$$0.034$
$\lambda$ $1,122.068$ $16.560$ $0.091$
$\zeta$ $285.927$ $0.167$ $0.071$
$\rho$ $0.063$ $0.017$ $0.004$
Notes: See Table 1 for the specific design of each case. All results are produced using $2000$ Monte Carlo replications.
Table 3: Frank Copula - RMSE
Case 0
Param. obs = 2000, $\rho$ = -0.5 obs = 5000, $\rho$ = -0.5 obs = 10000, $\rho$ = -0.5
$\beta_{0}$ $2,860.340$ $310.396$ $61.885$
$\beta_{1}$ $8,978.860$ $2,086.819$ $816.737$
$\pi_{0}$ $2,553.201$ $688.355$ $2.279$
$\pi_{1}$ $7,481.836$ $2,942.766$ $4.219$
$\lambda$ $7,083.483$ $1,754.262$ $714.109$
$\zeta$ $5,732.446$ $2,576.872$ $2.182$
$\rho$ $0.650$ $0.575$ $0.489$
Case 1-a
Param. obs = 2000, $\rho$ = -0.5 obs = 5000, $\rho$ = -0.5 obs = 10000, $\rho$ = -0.5
$\beta_{0}$ $684.210$ $0.190$ $0.126$
$\beta_{1}$ $3,169.837$ $0.471$ $0.285$
$\pi_{0}$ $2,328.631$ $1.059$ $0.736$
$\pi_{1}$ $1,702.444$ $0.685$ $0.474$
$\pi_{2}$ $1,661.837$ $0.433$ $0.293$
$\lambda$ $2,992.510$ $0.390$ $0.239$
$\zeta$ $1,925.182$ $0.607$ $0.428$
$\rho$ $0.488$ $0.362$ $0.233$
Case 1-b
Param. obs = 2000, $\rho$ = -0.5 obs = 5000, $\rho$ = -0.5 obs = 10000, $\rho$ = -0.5
$\beta_{0}$ $5,103.763$ $211.961$ $0.187$
$\beta_{1}$ $11,215.290$ $608.039$ $0.436$
$\pi_{0}$ $3,290.975$ $1.583$ $1.146$
$\pi_{1}$ $2,035.018$ $0.756$ $0.561$
$\pi_{2}$ $1,936.292$ $0.576$ $0.341$
$\lambda$ $13,803.720$ $833.790$ $0.522$
$\zeta$ $3,457.443$ $1.040$ $0.719$
rho $0.595$ $0.505$ $0.400$
Case 2-a
Param. obs = 2000, $\rho$ = -0.5 obs = 5000, $\rho$ = -0.5 obs = 10000, $\rho$ = -0.5
$\beta_{0}$ $0.323$ $0.165$ $0.110$
$\beta_{1}$ $0.621$ $0.265$ $0.175$
$\beta_{2}$ $0.549$ $0.233$ $0.150$
$\pi_{0}$ $616.635$ $0.476$ $0.335$
$\pi_{1}$ $438.369$ $0.359$ $0.249$
$\pi_{2}$ $484.978$ $0.308$ $0.215$
$\lambda$ $0.712$ $0.291$ $0.191$
$\zeta$ $587.512$ $0.415$ $0.288$
$\rho$ $0.171$ $0.103$ $0.072$
Case 2-b
Param. obs = 2000, $\rho$ = -0.5 obs = 5000, $\rho$ = -0.5 obs = 10000, $\rho$ = -0.5
$\beta_{0}$ $2,581.494$ $107.814$ $0.164$
$\beta_{1}$ $4,770.145$ $256.687$ $0.261$
$\beta_{2}$ $4,687.187$ $222.714$ $0.223$
$\pi_{0}$ $3,211.289$ $0.958$ $0.587$
$\pi_{1}$ $1,913.604$ $0.512$ $0.305$
$\pi_{2}$ $2,048.146$ $0.454$ $0.261$
$\lambda$ $8,605.971$ $459.471$ $0.410$
$\zeta$ $3,794.507$ $0.889$ $0.509$
$\rho$ $0.255$ $0.146$ $0.104$
Notes: See Table 1 for the specific design of each case. All results are produced using $2000$ Monte Carlo replications.
Analyzing the results presented in Table 2, we observe a pronounced MB across all scenarios for sample sizes of $n=2000$ and $n=5000$, which substantially diminishes at $n=10000$. This behavior aligns with the findings for the RMSE in Table 3. In Case 0, the absence of exclusion restrictions results in a persistent identification issue across all sample sizes, as evidenced by the excessive MB and RMSE. Notably, scenarios with a single (continuous) excluded variable in only one of the equations, specifically Cases 1-a and 1-b, also demonstrate substantial MB at $n=10000$, especially for intercepts, scaling parameters, and dependence parameters. Conversely, in Cases 2-a and 2-b, which feature two exclusion restrictions (with continuous excluded variables), MB and RMSE are substantially reduced compared to their counterparts in Cases 1-a and 1-b. Interestingly, the (unconditional) variance of the index emerges as a crucial factor, as demonstrated by the larger MB and RMSE in Cases 1-b and 2-b, where the index variance is substantially smaller than that of the error. [^12] The RMSE results for Cases 1-b and 2-b indicate that various parameters are very imprecisely estimated at smaller sample sizes, often attributed to outlier estimates. A detailed investigation reveals that this imprecision often originates from outlier estimates, further emphasizing the importance of a large index variance for precise estimation.
In summary, our findings from this section suggest that employing at least two exclusion restrictions (one in each equation) effectively reduces the finite sample MB and RMSE when estimating model parameters of the joint distribution in a partially observable framework. Furthermore, a larger index variance contributes to precise estimation.
9 Empirical Application
The intense debate surrounding the legalization of marijuana (cannabis) and its impact on individual consumption behavior, and more generally, on society, has been a pivotal policy issue in Western countries for decades. In many of these nations, the possession and consumption of marijuana are illegal, primarily due to concerns over its contribution to crime escalation and the associated societal and economic costs (Brinkman and Mok-Lamme, 2019). Additionally, there are health-related concerns, with evidence suggesting that early-age cannabis consumption can lead to lower educational attainment (Van Ours and Williams, 2009; Moore et al., 2007; Macleod et al., 2004) and cognitive deficits in verbal learning and memory tasks (Solowij et al., 2011). The potential of cannabis serving as a gateway to more harmful drug use has also been documented (see Pudney, 2010, and references therein).
On the contrary, various other countries have initiated a legalization process of the possession and consumption of certain categories of cannabis drugs for recreational and medicinal purposes, driven by aims to reduce drug-related crime and corresponding law enforcement costs, as well as to generate additional tax revenue. Notably, nearly one-quarter of the U.S. states have legalized recreational marijuana by now, regulating and taxing its sales (Brinkman and Mok-Lamme, 2019; Hansen et al., 2020).
In this section, we investigate the effects of recreational marijuana legalization for adults in several U.S. states on the consumption behavior of 8th-grade high-school students in those states in 2016. The focus on these minors, typically aged 13 to 14, is critical due to their heightened vulnerability to the health risks associated with marijuana consumption. Unlike existing studies, which predominantly examined legalization effects on average usage or consumption (see Anderson et al., 2015; Wen et al., 2015; Cerdá et al., 2017; Hollingworth et al., 2022), or on the timing of the first (early age) consumption (Williams and Bretteville-Jensen, 2014), we investigate the distributional effects of recreational marijuana legalization on the entire consumption distribution within the aforementioned high-school cohort in various states. [^13] Unlike Hollingworth et al. (2022), who consider mean effects for the legalization of recreational marijuana usage on underage adolescents among others, the survey data used in this analysis exclusively covers enrolled high-school students (see next paragraph). Thus, it does not provide insights into effects on high-school drop-outs. Additionally, we account for the possibility that consumption behavior may not always be reported truthfully, and more importantly, that reporting behavior itself may be affected by the legalization, possibly due to shifts in stigma, public perception, or attitudes towards marijuana consumption.
The data is composed of several cross-sectional waves of the “Monitoring the Future” survey, an annual survey conducted in the United States gathering data on attitudes, behaviors, and values of American adolescents currently enrolled in high school. [^14] The “Monitoring the Future” survey is funded by the National Institute on Drug Abuse (NIDA). For more information, see Monitoring the Future, Institute for Social Research, University of Michigan (https://monitoringthefuture.org/). Specifically, the survey for 8th graders, which is conducted annually, covers a wide range of topics including substance (ab)use such as the consumption of marijuana and variants thereof. The primary sample of our empirical analysis consists of 46,472 high-school students, sampled as repeated cross-sections from 2015 to 2018 across all contiguous U.S. states except Oregon, Washington, and Colorado. With this sample, our focus is on the recreational marijuana legalization for adults aged 21 or older in California, Nevada, Massachusetts, and Maine in November 2016, examining its impact on 8th-grade students’ (short-term) consumption behavior. [^15] These states had previously legalized medicinal marijuana use before the observation period. Moreover, even though private cultivation of marijuana for personal consumption became legal (almost) immediately, the first non-medical retail sales of marijuana did not start until some time later in most cases, (e.g. not before 2018 in Maine and California). We classify the observations from 2015-2016 into the pre-treatment period ($t=0$), and those from 2017-2018 into post-treatment period ($t=1$). On the other hand, we group the observations from California, Nevada, Massachusetts and Maine into the treated group ($g=1$), and the rest into the control group ($g=0$). This state grouping is mainly owed to data privacy rules imposed by the hosting institute, preventing a more disaggregate analysis. It is also important to note that no other major marijuana-related legalization took place in the U.S. during this observation period, either for medicinal or recreational purposes. Finally, given these surveys were conducted in springs, we consider the 2016 data to be sufficiently removed from the actual legalization to avoid major anticipation effects.
The outcome variable of interest (cons_30days), based on the survey responses to the question “On how many occasions have you used marijuana (weed, pot) or hashish (hash, hash oil) … in the last 30 days?”, is categorized into three levels: “0” (never used), “1” (used 1-2 times), and “2” (used more than 2 times). Similar to Hollingworth et al. (2022), we not only include individual-level covariates like gender, age, ethnicity, and parental education, but also certain state-level control variables such as unemployment rate and median income, fixed at pre-treatment levels. [^16] State-level unemployment rates as well as the state median income were constructed from data of the Bureau of Labor Statistics (https://www.bls.gov/). Finally, we employ state-level marijuana prices from 2015 as an instrument for the consumption (state_price_m), and a “survey cooperation” measure (trust) for the reporting, following Greene et al. (2018) and Brown et al. (2018). [^17] Marijuana price data source:
https://oxfordtreatment.com/substance-abuse/marijuana/average-cost-of-marijuana/ (last accessed: 12-10-2023).
Here, marijuana prices are presumed to exclusively influence consumption decisions. The plausibility of the exclusion restriction relies on the assumption that consumption behavior is sensitive to prices, while reporting behavior is not, given that we also control for other state-level variables. On the other hand, “survey cooperation”, defined as the proportion of missing survey responses (based solely on questions common to all questionnaire forms and not directly related to drug use), is assumed to provide an indication of the respondents’ willingness to cooperate in the survey, unrelated to marijuana consumption itself, as the reporting decision occurs post-consumption only.
Table 4 presents descriptive statistics of the outcome variable and other controls. As expected, marijuana consumption among 8th graders appears relatively low, with only 6% of the sample reporting usage over the past 30 days. Further, unreported results indicate that these characteristics remain largely stable across time within each group.
Table 4: Summary Statistics
Variable N Mean Std. Dev. Min Pctl. 25 Pctl. 75 Max
cons_30days 46472
… 1-2 occasions 1290 3%
… 3 or more occasions 1375 3%
… no use 43807 94%
male 46472 0.48 0.5 0 0 1 1
age_month 46472 169 6.2 122 165 173 231
white 46472 0.47 0.5 0 0 1 1
parent_edu_clg 46472 0.58 0.49 0 0 1 1
state_unemp 46472 5.3 0.78 2.9 4.8 5.9 6.8
state_median_income 46472 56552 7024 40037 51983 60925 75675
state_price_m 46472 240 35 159 220 265 325
trust 46472 0.96 0.072 0.28 0.96 1 1
treated 46472 0.19 0.39 0 0 0 1
after 46472 0.47 0.5 0 0 1 1
Notes: Data for 8th graders from 2015-2018, excluding the states Oregon, District of Columbia, Colorado and Washington. treated = 1 if the respondent is from California, Maine, Massachusetts and Nevada, = 0 otherwise. after = 1 if the respondent is surveyed in 2017-2018, = 0 otherwise. The variables male, white, parent_edu_clg are binary indicators for self-reported gender status (male), ethnicity (white), and parental education level (college) equal to one if the category was true, and zero otherwise. For the other variables, please refer to the text for the definition.
In Table 5, we present our initial findings from the unconditional analysis, as defined in (21), without misreporting and covariates. Our first set of results (gdid) suggests that $\tau^{c}_{11}(j)$ is not statistically different from zero for $j\in\{0,1,2\}$ at any conventional levels. These results align with the estimates from the more restrictive ordered (pdid_ordered) and binary choice (pdid_probit) DiD models, which include the binary time period and group indicators and their interaction as regressors.
Table 5: Unconditional Treatment Effects (No Misreporting)
pdid pdid (sd) gdid_ordered gdid_ordered (sd) gdid gdid (sd)
$\tau^{c}_{11}(0)$0.0071 0.0056 -0.0067 0.0056 -0.0071 0.0056
$\tau^{c}_{11}(1)$ 0.0071 0.0056 0.0027 0.0045 0.0047 0.0040
$\tau^{c}_{11}(2)$ NA NA 0.0039 0.0033 0.0025 0.0042
Notes: Data between 2015-2018 are used.  The table contains the estimated distributional treatment effects $\tau_{11}^{c}(j)$, $j\in\{0,1,2\}$ for the Probit DiD with binary outcomes (pdid_binary), the Probit DiD with ordered outcomes (pdid_ordered), and generalized DiD proposed in this paper (gdid). For the Probit DiD with binary outcomes, $\tau^{c}_{11}(1)$ indicates the effect for users with outcomes $\geq$ 1. Standard errors are computed via the Delta method and can be found in the corresponding columns next to the coefficient estimates.
Next, we move to specifications without misreporting but including covariates to account for differences across groups in terms of observed characteristics. We present estimates of $\tau_{11}^{c}(j)$ in Table 6 (first panel), directing readers to Section LABEL:sec:additionalempirical in the supplementary material for parameter estimates. Given nearly identical variance parameters across groups and time in Table 5, we constrain $\lambda_{gt}$ to be constant in this specification. [^18] Unreported results (available upon request) show that allowing for flexible variance parameters does not alter the estimates qualitatively. Notably, the estimates of $\widehat{\tau}_{11}^{c}(j)$ now appear to be larger in absolute terms, with a significant decrease in non-consumption probability by around 2 percentage points, and significant increases in consumption probabilities by approximately 1 percentage point. Compared to the unconditional results, conditioning on additional regressors seems to amplify the effect sizes in absolute terms.
In the final step, we estimate models with covariates incorporating misreporting according to Section 5 (see second panel of Table 6). As before, we only present estimates of the marginalized treatment effects $\tau_{11}^{c}(j)$ and $\tau_{11}^{r}(j)$ on consumption and misreporting, respectively, as defined in (24), with further details available in Section LABEL:sec:additionalempirical of the supplement. It is noteworthy to mention that the slope estimates for the consumption decision closely mirror those from the specification without misreporting. Both of the instruments are found to be significant at the 5 or 10% level across groups and time (except for the treatment group after treatment). [^19] The positive and significant effects for marijuana prices align with other studies (e.g. Greene et al., 2018), potentially reflecting drug quality and availability effects, in the absence of quality as a control variable). Notably, we see more pronounced effect sizes after accounting for misreporting (increased by more than 50% for all levels), with all $\widehat{\tau}_{11}^{c}(j)$ being significant at the 5% level. No significant treatment effects are found on the reporting behavior. These findings are robust to a change in the Copula function to the asymmetric Clayton Copula.
Table 6: Distributional Treatment Effects (Marginalized)
Without Misreporting
gdid gdid (sd)
$\tau_{11}^{c}(0)$0.0200 0.0082
$\tau_{11}^{c}(1)$ 0.0088 0.0036
$\tau_{11}^{c}(2)$ 0.0113 0.0047
With Misreporting
gdid_frank gdid_frank (sd) gdid_clayton gdid_clayton (sd)
$\tau^{c}_{11}(0)$0.0327 0.0137 -0.0327 0.0141
$\tau^{c}_{11}(1)$ 0.0138 0.0057 0.0138 0.0058
$\tau^{c}_{11}(2)$ 0.0189 0.0081 0.0189 0.0083
$\tau^{r}_{11}(0)$ 0.1199 0.1213 0.1242 0.1442
$\tau^{r}_{11}(1)$ 0.0125 0.0116 0.0179 0.0268
$\tau^{r}_{11}(2)$0.1324 0.1307 -0.1421 0.1698
Data between 2015-2018 are used.  The first panel of results (“Without Misreporting”) displays estimates of treatment effects $\tau_{11}(j)$, $j\in\{0,1,2\}$, for the model estimated without misreporting. The treatment effect estimates have been marginalized with respect to the distribution of $\mathbf{X}_{11}$ (see Equation (21)). The second panel of results (“With Misreporting”) displays estimates of treatment effects $\tau_{11}(j)$ and $\tau_{11}^{r}(j)$, $j\in\{0,1,2\}$, for the model with misreporting. The treatment effect estimates have also been marginalized with respect to the distribution of $\mathbf{X}_{11}$. The first two columns display results for a Frank Copula specification, the last two columns show corresponding results for a Clayton Copula specification. Standard errors are computed via the Delta method and can be found in the corresponding columns next to the coefficient estimates.
In summary, the empirical findings of this section reveal nuanced effects of recreational marijuana legalization on 8th-grade high-school students’ short-term consumption and reporting behavior. Our initial analysis without misreporting and covariates showed no significant changes in consumption probabilities. However, after introducing covariates, we observed a more pronounced impact, with a significant decrease in non-consumption probability and a significant increase in consumption probabilities. These effect sizes were further amplified and remain significant after accounting for misreporting, while no significant treatment effects are found on the reporting behavior. Nevertheless, it should be noted that, as a limitation of our study, the categorization of the outcome variable focuses on usage frequency and does not capture the doses consumed per occasion, potentially overlooking variations in consumption intensity.
10 Conclusion
This paper develops a generalized Difference-in-Difference model for discrete, possibly underreported outcomes in repeated cross-sectional or panel data. The models are based on the continuous CiC model of Athey and Imbens (2006). For the case without misreporting and covariates, we propose a parametric model that generalizes standard Probit or Logit DiD models which only accommodate mean shifts. When covariates are present (but outcomes are reported correctly), we relax the distributional assumption on the unobservables, and provide a point identification result for a semiparametric index model. The result is complementary to Athey and Imbens (2006) and offers a direct pathway to estimation of the structural model and corresponding treatment effects.
In the second step, we extend the CiC model for discrete outcomes to partially observable outcomes. Special features of this extended model are that it allows the underreporting behavior to be correlated with the outcome of interest and that underreporting may depend on the treatment status itself. We provide point identification conditions for a semiparametric version of the model based on a real analytic extrapolation assumption, and suggest a flexible estimator for a parametric version of the model.
A set of Monte Carlo simulations investigates the finite sample performance of the corresponding ML estimator under different reporting and exclusion restriction scenarios. The results suggest that a single (or no) exclusion restriction may yield poor finite sample performance in the partially observable outcome framework. We apply our methodology to investigate the distributional effects of the legalization of recreational marijuana usage for adults in the U.S. states of California, Nevada, Massachusetts, and Maine in November 2016 on the consumption behavior of 8th-grade high-school students. Our findings suggest that there are small, positive significant effects on the (short-term) consumption probabilities, but no significant changes in the reporting behavior as a result of the legalization. Importantly, correcting for misreporting increases effect sizes by more than 50%.
Appendix
Appendix A Alternative Characterization: Distributional Parallel Trends
While Proposition 1 outlines the relationships among the functional parameters across different groups and over time, the subsequent result provides a further characterization of how models that satisfy distributional parallel trends are generated:
Proposition A.1.
Suppose Assumption 1 holds. Then, distributional parallel trends from Definition 9 holds if and only if there exist real-valued functions $\alpha_{t}(\mathbf{x}),\mu_{g}(\mathbf{x}),\beta_{t}(\mathbf{x})$ and $\sigma_{g}(\mathbf{x})$ such that $\beta_{t}(\mathbf{x})\sigma_{g}(\mathbf{x})>0$, $g\in\{0,1\}$ and $t\in\{0,1\}$, such that
$\eta_{gt}(\mathbf{x})=\alpha_{t}(\mathbf{x})+\beta_{t}(\mathbf{x})\mu_{g}(\mathbf{x})\quad\text{and}\quad\lambda_{gt}(\mathbf{x})=\beta_{t}(\mathbf{x})\sigma_{g}(\mathbf{x})$ (25)
for all covariate values $\mathbf{x}\in\cap_{g,t}\mathcal{X}_{gt}$,
Proposition A.1 demonstrates that the distributional parallel trends hold if and only if the functional median (mean) and the scaling parameters admit a specific form of separability in group and time. Specifically, each functional median (mean) parameter $\eta_{gt}(\mathbf{x})$ can be expressed as a bilinear form comprising a time-trend component $(\alpha_{t}(\mathbf{x}),\beta_{t}(\mathbf{x}))^{\prime}$ and a group-specific component $(1,\mu_{g}(\mathbf{x}))^{\prime}$, and each scaling parameter $\lambda_{gt}(\mathbf{x})$ follows a similar structure involving $\beta_{t}(\mathbf{x})$ and $\sigma_{g}(\mathbf{x})$. We highlight that this kind of separability imposes no substantial limitations on how the time trend, the group membership, or the observed heterogeneity affect the outcome variable. Furthermore, given the representations in (25), it becomes evident that our model is nonnested within the discrete CiC models with covariates considered by Athey and Imbens (2006), where, when fixing the covariates at specific values, the group membership only allows to affect the outcome variable through the distribution of the unobservable (that is assumed to be within-group time invariant and independent of covariates). In contrast, the group membership can still alter the way the covariates affect the outcome variable in our case.
Provided identification of the functional median (mean) and scaling parameters, Corollary A.2 identifies the underlying structure in the representations (25) up to normalization:
Corollary A.2.
Suppose that the functional parameters $\eta_{gt}(\mathbf{x})$ and $\lambda_{gt}(\mathbf{x})$ are identified for all covariate values $\mathbf{x}\in\cap_{g,t}\mathcal{X}_{gt}$, $g\in\{0,1\}$ and $t\in\{0,1\}$. If we adopt $\beta_{0}(\mathbf{x})=1$ as an normalization assumption, then $\beta_{1}(\mathbf{x})$ and $\sigma_{g}(\mathbf{x})$ are identified as
$\beta_{1}(\mathbf{x})=\frac{\lambda_{01}(\mathbf{x})}{\lambda_{00}(\mathbf{x})}\quad\text{and}\quad\sigma_{g}(\mathbf{x})=\lambda_{g0}(\mathbf{x}).$
for $g\in\{0,1\}$. If we further adopt $\alpha_{0}(\mathbf{x})=0$ as an normalization assumption, then $\alpha_{1}(\mathbf{x})$ and $\mu_{g}(\mathbf{x})$ are identified as
$\alpha_{1}(\mathbf{x})=\eta_{01}(\mathbf{x})-\eta_{00}(\mathbf{x})\frac{\lambda_{01}(\mathbf{x})}{\lambda_{00}(\mathbf{x})}\quad\text{and}\quad\mu_{g}(\mathbf{x})=\eta_{g0}(\mathbf{x})$
for $g\in\{0,1\}$.
It should be noted that the choice of normalizations is arbitrary. It merely alters the interpretation of the functional parameters but places no additional restrictions. Here, we have chosen to normalize the common time trend in the pre-treatment period to facilitate the interpretation of heterogeneity among the groups. Alternatively, a different choice of normalizations could be setting $\mu_{0}(\mathbf{x})=0$ and $\sigma_{0}(\mathbf{x})=1$.
In light of the structure in (25), we construct a model for the latent outcome $Y^{\ast}_{gt}(0)$, paralleling the features in the assumptions for the continuous CiC model of Athey and Imbens (2006) without covariates, as an example that satisfies the distributional parallel trends in our model with covariates. We define
$Y^{*}_{gt}(0)=h_{t}(\mathbf{X}_{gt},\varepsilon_{gt}),$ (26)
where $h_{t}(\cdot,\cdot)$ denotes a real-valued function that possibly changes over time. We make the following assumptions:
Assumption A.1.
For each $t\in\{0,1\},\mathbf{x}\in\mathbb{R}^{k}$ and $u\in\mathbb{R}$:
$h_{t}(\mathbf{x},u)=\alpha_{t}(\mathbf{x})+\beta_{t}(\mathbf{x})u,$
where $\alpha_{t}:\mathbb{R}^{d_{X_{gt}}}\rightarrow\mathbb{R}$ and $\beta_{t}:\mathbb{R}^{d_{X_{gt}}}\rightarrow\mathbb{R}_{+}$ are real-valued functions.
Assumption A.2.
$\varepsilon_{gt}$ is strictly stationary within each group given $\mathbf{X}_{gt}$.
Given Assumptions A.1-A.2 and denoting the conditional expectation and variance of $\varepsilon_{gt}$ given $\mathbf{X}_{gt}$ as $\mu_{g}(\mathbf{X}_{gt})$ and $\sigma^{2}_{g}(\mathbf{X}_{gt})$, respectively, we obtain
$\mathrm{E}\left[Y_{gt}^{*}(0)\mid\mathbf{X}_{gt}=\mathbf{x}\right]$ $=\alpha_{t}(\mathbf{x})+\beta_{t}(\mathbf{x})\mu_{g}(\mathbf{x})=:\eta_{gt}(\mathbf{x}),$
$\mathrm{Var}\left[Y_{gt}^{*}(0)\mid\mathbf{X}_{gt}=\mathbf{x}\right]$ $=\left(\beta_{t}(\mathbf{x})\sigma_{g}(\mathbf{x})\right)^{2}=:\lambda^{2}_{gt}(\mathbf{x}),$
for all $\mathbf{x}\in\mathcal{X}_{gt}$, which thus satisfies the distributional parallel trends in our model with covariates.
Appendix B Proofs
Proof of Proposition 1.
For the “if” part, under Assumption 1
$F_{Y^{*}_{11}(0)\mathbf{X}_{11}}(v\mid\mathbf{x})$ $=F_{\varepsilon_{11}\mid\mathbf{X}_{11}}\left(\frac{v-\eta_{11}(\mathbf{x})}{\lambda_{11}(\mathbf{x})}\mid\mathbf{x}\right)$
$=F_{\varepsilon_{11}\mid\mathbf{X}_{11}}\left(\frac{v-\eta_{01}(\mathbf{x})-\frac{\eta_{10}(\mathbf{x})-\eta_{00}(\mathbf{x})}{\lambda_{00}(\mathbf{x})/\lambda_{01}(\mathbf{x})}}{\frac{\lambda_{01}(\mathbf{x})\lambda_{10}(\mathbf{x})}{\lambda_{00}(\mathbf{x})}}\mid\mathbf{x}\right)$
$=F_{\varepsilon_{10}\mid\mathbf{X}_{10}}\left(\frac{\eta_{00}(\mathbf{x})+\lambda_{00}(\mathbf{x})F_{\varepsilon_{00}\mid\mathbf{X}_{00}}^{-1}\left(F_{\varepsilon_{01}\mid\mathbf{X}_{01}}\left(\frac{v-\eta_{01}(\mathbf{x})}{\lambda_{01}(\mathbf{x})}\mid\mathbf{x}\right)\mid\mathbf{x}\right)-\eta_{10}(\mathbf{x})}{\lambda_{10}(\mathbf{x})}\mid\mathbf{x}\right)$
$=F_{Y^{*}_{10}(0)\mid\mathbf{X}_{10}}\left(F_{Y^{*}_{00}(0)\mid\mathbf{X}_{00}}^{-1}\left(F_{Y^{*}_{01}(0)\mid\mathbf{X}_{01}}(v\mid\mathbf{x})\mid\mathbf{x}\right)\mid\mathbf{x}\right),$
for all $v\in\mathbb{R}$ and for all covariate values $\mathbf{x}\in\cap_{g,t}\mathcal{X}_{gt}$, where we used strict stationarity $\varepsilon_{g1}\mid X_{g1}\sim\varepsilon_{g0}\mid X_{g0}$ in the third equality, and $F_{Y^{*}_{gt}(0)\mid\mathbf{X}_{gt}}^{-1}(q\mid\mathbf{x})=\eta_{gt}(\mathbf{x})+\lambda_{gt}(\mathbf{x})F_{\varepsilon_{gt}}^{-1}(q\mid\mathbf{x})$ in the last equality.
For the “only if” part, let $\mathbf{x}\in\cap_{g,t}\mathcal{X}_{gt}$. Given Assumption 1 and the distributional parallel trends, there exists an open interval $I$ such that
$F_{Y^{*}_{11}(0)\mid\mathbf{X}_{11}}(v\mid\mathbf{x})$ $=F_{\varepsilon_{11}\mid\mathbf{X}_{11}}\left(\frac{v-\eta_{11}(\mathbf{x})}{\lambda_{11}(\mathbf{x})}\mid\mathbf{x}\right),$
which is equal to
$\ F_{Y^{*}_{10}(0)\mid\mathbf{X}_{10}}\left(F_{Y^{*}_{00}(0)\mid\mathbf{X}_{00}}^{-1}\left(F_{Y^{*}_{01}(0)\mid\mathbf{X}_{01}}(v\mid\mathbf{x})\mid\mathbf{x}\right)\mid\mathbf{x}\right)$
$=F_{\varepsilon_{10}\mid\mathbf{X}_{10}}\left(\frac{\eta_{00}(\mathbf{x})+\lambda_{00}(\mathbf{x})F_{\varepsilon_{00}\mid\mathbf{X}_{00}}^{-1}\left(F_{\varepsilon_{01}\mid\mathbf{X}_{01}}\left(\frac{v-\eta_{01}(\mathbf{x})}{\lambda_{01}(\mathbf{x})}\mid\mathbf{x}\right)\mid\mathbf{x}\right)-\eta_{10}(\mathbf{x})}{\lambda_{10}(\mathbf{x})}\mid\mathbf{x}\right)$
for all $v\in I$. The strict stationarity $\varepsilon_{g1}\mid\mathbf{X}_{g1}\sim\varepsilon_{g0}\mid\mathbf{X}_{g0}$ further implies a polynomial in $v$ on the open interval $I$:
$-\frac{\eta_{11}(\mathbf{x})}{\lambda_{11}(\mathbf{x})}-\frac{\eta_{00}(\mathbf{x})-\lambda_{00}(\mathbf{x})\frac{\eta_{01}(\mathbf{x})}{\lambda_{01}(\mathbf{x})}-\eta_{10}(\mathbf{x})}{\lambda_{10}(\mathbf{x})}+\left(\frac{1}{\lambda_{11}(\mathbf{x})}-\frac{\lambda_{00}(\mathbf{x})}{\lambda_{01}(\mathbf{x})\lambda_{10}(\mathbf{x})}\right)v=0.$
Note that the polynomial is well defined as $\lambda_{gt}(\mathbf{x})>0$ for all $(g,t)\in\{0,1\}$. However, a polynomial on an open interval gives rise to the zero function if and only if it is the zero polynomial, that is,
$-\frac{\eta_{11}(\mathbf{x})}{\lambda_{11}(\mathbf{x})}-\frac{\eta_{00}(\mathbf{x})-\lambda_{00}(\mathbf{x})\frac{\eta_{01}(\mathbf{x})}{\lambda_{01}(\mathbf{x})}-\eta_{10}(\mathbf{x})}{\lambda_{10}(\mathbf{x})}=0\quad\text{and}\quad\frac{1}{\lambda_{11}(\mathbf{x})}-\frac{\lambda_{00}(\mathbf{x})}{\lambda_{01}(\mathbf{x})\lambda_{10}(\mathbf{x})}=0,$
which implies
$\eta_{11}(\mathbf{x})=\eta_{01}(\mathbf{x})+\frac{\eta_{10}(\mathbf{x})-\eta_{00}(\mathbf{x})}{\lambda_{00}(\mathbf{x})/\lambda_{01}(\mathbf{x})}\quad\text{and}\quad\lambda_{11}(\mathbf{x})=\frac{\lambda_{01}(\mathbf{x})\lambda_{10}(\mathbf{x})}{\lambda_{00}(\mathbf{x})}.$
Proof of Proposition 2.
Let $(g,t)\neq(1,1)$. Since $\kappa_{1}$ and $\kappa_{2}$ are normalized to 0 and 1, we can write the observed fitted probabilities:
$\mathrm{Pr}\left(Y_{gt}=0\right)=F_{\varepsilon_{gt}}\left(-\frac{\eta_{gt}}{\lambda_{gt}}\right)\quad\text{and}\quad\mathrm{Pr}\left(Y_{gt}\leq 1\right)=F_{\varepsilon_{gt}}\left(\frac{1-\eta_{gt}}{\lambda_{gt}}\right).$
Given $F_{\varepsilon_{gt}}$ is invertible and the probabilities are nonzero by Assumptions 1, we build the system of equations:
$a+F_{\varepsilon_{gt}}^{-1}\left(\mathrm{Pr}\left(Y_{gt}=0\right)\right)b=0$
$a+F_{\varepsilon_{gt}}^{-1}\left(\mathrm{Pr}\left(Y_{gt}\leq 1\right)\right)b=1.$
If $F_{\varepsilon_{gt}}^{-1}\left(\mathrm{Pr}\left(Y_{gt}=0\right)\right)$ and $F_{\varepsilon_{gt}}^{-1}\left(\mathrm{Pr}\left(Y_{gt}\leq 1\right)\right)$ are known, then $(\eta_{gt},\lambda_{gt})$ is uniquely recovered by
$\left(\begin{array}[]{ll}\eta_{gt}\\ \lambda_{gt}\end{array}\right)$ $=\left(\begin{array}[]{ll}1&F_{\varepsilon_{gt}}^{-1}\left(\mathrm{Pr}\left(Y_{gt}=0\right)\right)\\ 1&F_{\varepsilon_{gt}}^{-1}\left(\mathrm{Pr}\left(Y_{gt}\leq 1\right)\right)\end{array}\right)^{-1}\left(\begin{array}[]{ll}0\\ 1\end{array}\right)$
$=\left(\begin{array}[]{ll}\frac{-F_{\varepsilon_{gt}}^{-1}\left(\mathrm{Pr}\left(Y_{gt}=0\right)\right)}{F_{\varepsilon_{gt}}^{-1}\left(\mathrm{Pr}\left(Y_{gt}\leq 1\right)\right)-F_{\varepsilon_{gt}}^{-1}\left(\mathrm{Pr}\left(Y_{gt}=0\right)\right)}\\ \frac{1}{F_{\varepsilon_{gt}}^{-1}\left(\mathrm{Pr}\left(Y_{gt}\leq 1\right)\right)-F_{\varepsilon_{gt}}^{-1}\left(\mathrm{Pr}\left(Y_{gt}=0\right)\right)}\end{array}\right),$
where the expressions on the RHS are well defined by Assumption 1. Consequently, given the distributional parallel trends, $\eta_{11}$ and $\lambda_{11}$ are identified by Proposition 1 as
$\eta_{11}=\eta_{01}+\frac{\eta_{10}-\eta_{00}}{\lambda_{00}/\lambda_{01}}\quad\text{ and }\quad\lambda_{11}=\frac{\lambda_{01}\lambda_{10}}{\lambda_{00}}.$
However, if $F_{\varepsilon_{gt}}^{-1}\left(\mathrm{Pr}\left(Y_{gt}=0\right)\right)$ and $F_{\varepsilon_{gt}}^{-1}\left(\mathrm{Pr}\left(Y_{gt}\leq 1\right)\right)$ are unknowns for $(g,t)\neq(1,1)$, then the system becomes a nonlinear system with four unknowns. In such a case, there are infinitely many solutions and so infinitely many observationally equivalent $(\eta_{gt},\lambda_{gt})$ contradicting point identification. ∎
Proof of Proposition 3.
Denoting $\overline{\boldsymbol{\eta}}_{gt}^{*}:=\frac{\overline{\boldsymbol{\eta}}_{gt}}{\left\|\overline{\boldsymbol{\eta}}_{gt}\right\|}$, $\kappa_{j,gt}^{*}=\frac{\kappa_{j}}{\left\|\overline{\boldsymbol{\eta}}_{gt}\right\|}$, $\widetilde{\varepsilon}_{gt}:=\lambda_{gt}\varepsilon_{gt}$ and $\widetilde{\varepsilon}_{gt}^{*}:=\frac{\widetilde{\varepsilon}_{gt}}{\left\|\overline{\boldsymbol{\eta}}_{gt}\right\|}$, the model can be reformulated as:
$Y_{gt}(0)=$ $\sum_{j=0}^{2}j\cdot\mathbb{I}\left[\kappa_{j,gt}^{*}\leq\overline{\mathbf{X}}_{gt}^{\prime}\overline{\boldsymbol{\eta}}^{*}_{gt}+\widetilde{\varepsilon}^{*}_{gt}\leq\kappa_{j+1,gt}^{*}\right].$
Let $(g,t)\neq(1,1)$. Since Assumptions 2-4 imply conditions SI, X3, and X5 in Manski (1988), we have by Corollary 5 therein that the parameter vector $\overline{\boldsymbol{\eta}}_{gt}^{*}$ is identified. As a consequence, the distribution function $F_{\widetilde{\varepsilon}_{gt}^{*}}$ of the rescaled error term $\widetilde{\varepsilon}_{gt}^{*}$, defined by $F_{\widetilde{\varepsilon}_{gt}^{*}}(v)=F_{\varepsilon_{gt}}\left(\frac{\left\|\overline{\boldsymbol{\eta}}_{gt}\right\|}{\lambda_{gt}}v\right)$, is identified on $\mathbb{R}$ by
$F_{\widetilde{\varepsilon}_{gt}^{*}}\left(-\overline{\mathbf{X}}_{gt}^{\prime}\overline{\boldsymbol{\eta}}_{gt}^{*}\right)=$ $\mathrm{E}\left[\Pr\left(Y_{gt}\leq 0\mid\overline{\mathbf{X}}_{gt}\right)\mid\overline{\mathbf{X}}_{gt}^{\prime}\overline{\boldsymbol{\eta}}_{gt}^{*}\right].$
Given that the inverse function $F_{\widetilde{\varepsilon}_{gt}^{*}}^{-1}$ exists under Assumption 1, we identify
$\left\|\overline{\boldsymbol{\eta}}_{gt}\right\|=\frac{1}{\kappa_{2,gt}^{*}}=\frac{1}{\mathrm{E}\left[F_{\widetilde{\varepsilon}_{gt}^{*}}^{-1}\left(\mathrm{E}\left[\Pr\left(Y_{gt}\leq 1\mid\mathbf{X}_{gt}\right)\mid\overline{\mathbf{X}}_{gt}^{\prime}\overline{\boldsymbol{\eta}}_{gt}^{*}\right]\right)+\overline{\mathbf{X}}_{gt}^{\prime}\overline{\boldsymbol{\eta}}_{gt}^{*}\right]}.$
As an immediate consequence, $\overline{\boldsymbol{\eta}}_{gt}$ is identified. The distribution function $F_{\widetilde{\varepsilon}_{gt}}$ of the rescaled error term $\widetilde{\varepsilon}_{gt}$, defined by $F_{\widetilde{\varepsilon}_{gt}}\left(v\right):=F_{\varepsilon_{gt}}\left(\frac{v}{\lambda_{gt}}\right)$, is identified by a similar argument from above. In fact, given identification of the rescaled error term distribution, $\lambda_{gt}$ is also straightforwardly identified as the standard deviation of that distribution, as well as $F_{{\varepsilon}_{gt}}$.
Now, focusing on $(g,t)=(1,1)$, identification of $\overline{\mathbf{x}}^{\prime}\overline{\boldsymbol{\eta}}_{11}$ and $\lambda_{11}$ directly follow from Proposition 1:
$\overline{\mathbf{x}}^{\prime}\overline{\boldsymbol{\eta}}_{11}=\overline{\mathbf{x}}^{\prime}\left(\overline{\boldsymbol{\eta}}_{01}+\frac{\overline{\boldsymbol{\eta}}_{10}-\overline{\boldsymbol{\eta}}_{00}}{\lambda_{00}/\lambda_{01}}\right)\quad\text{and}\quad\lambda_{11}=\frac{\lambda_{01}\lambda_{10}}{\lambda_{00}},$
for all covariate values $\mathbf{x}\in\mathcal{X}_{11}$. As $\overline{\mathcal{X}}_{11}$ does not lie in a proper linear subspace of $\mathbb{R}^{k+1}$ by Assumption 5, $\overline{\boldsymbol{\eta}}_{11}$ is identified as
$\overline{\boldsymbol{\eta}}_{11}=\overline{\boldsymbol{\eta}}_{01}+\frac{\overline{\boldsymbol{\eta}}_{10}-\overline{\boldsymbol{\eta}}_{00}}{\lambda_{00}/\lambda_{01}}.$
The identification of $F_{{\varepsilon}_{11}}$ directly follows from the strict stationarity in Assumption 1. ∎
Proof of Proposition 4.
Denote $\widetilde{\varepsilon}_{gt}:=\lambda_{gt}\varepsilon_{gt}$ and $\widetilde{\nu}_{gt}:=\zeta_{gt}\nu_{gt}$, as well as the composite distribution functions $F_{-\widetilde{\varepsilon}_{gt}}$ and $F_{-\widetilde{\nu}_{gt}}$ by
$F_{-\widetilde{\varepsilon}_{gt}}\left(v\right):=F_{-\varepsilon_{gt}}\left(\frac{v}{\lambda_{gt}}\right)\quad\text{and}\quad F_{-\widetilde{\nu}_{gt}}\left(v\right):=F_{-\nu_{gt}}\left(\frac{v}{\zeta_{gt}}\right).$
for any $v\in\mathbb{R}$. Under Assumption 6, it holds that
$\Pr\left(C_{gt}(0)\geq j\mid\mathbf{X}_{gt},\mathbf{Z}_{gt}\right)$ $=\Pr\left(Y_{gt}(0)\geq j,R_{gt}(0)\geq j\mid\mathbf{X}_{gt},\mathbf{Z}_{gt}\right)$
$=\Pr\left(\overline{\mathbf{X}}_{gt}^{\prime}\overline{\boldsymbol{\eta}}_{gt}+\lambda_{gt}\varepsilon_{gt}\geq\kappa_{j},\overline{\mathbf{Z}}_{gt}^{\prime}\overline{\boldsymbol{\pi}}_{gt}+\zeta_{gt}\nu_{gt}\geq\iota_{j}\right)$
$=\overline{C}_{gt}\left(F_{-\widetilde{\varepsilon}_{gt}}\left(\overline{\mathbf{X}}_{gt}^{\prime}\overline{\boldsymbol{\eta}}_{gt}-\kappa_{j}\right),F_{-\widetilde{\nu}_{gt}}\left(\overline{\mathbf{Z}}_{gt}^{\prime}\overline{\boldsymbol{\pi}}_{gt}-\iota_{j}\right)\right).$
Let $(g,t)\neq(1,1)$. By Assumptions 6 and 7, the signs of $\eta_{1,gt}$ and $\pi_{1,gt}$ are nonzero and identified as:
$\operatorname{sgn}\left(\mathrm{E}\left[\frac{\partial}{\partial Z_{1,gt}}\Pr\left(C_{gt}\geq j|\mathbf{W}_{gt},X_{1,gt},Z_{1,gt}\right)\right]\right)$
and
$\operatorname{sgn}\left(\mathrm{E}\left[\frac{\partial}{\partial X_{1,gt}}\Pr\left(C_{gt}\geq j|\mathbf{W}_{gt},X_{1,gt},Z_{1,gt}\right)\right]\right),$
respectively, for some $1\leq j\leq 2$, where $\operatorname{sgn}(x)=1\{x>0\}-1\{x<0\}$ is the sign function. W.l.o.g. let $\pi_{1,gt}>0$ for the rest of the proof, since otherwise we could consider $-\pi_{1,gt}$. Also, as the below exposition applies to each $1\leq j\leq 2$, we focus on a given $j$ level for simplicity. Under Assumption 6, given any $\mathbf{w}\in\mathbb{R}^{k-1}$, the function
$(x,z)\mapsto{c}_{\mathbf{w},gt}(x,z):=\overline{C}_{gt}\left(F_{-\widetilde{\varepsilon}_{gt}}\left(\eta_{1,gt}x+\overline{\mathbf{w}}^{\prime}\overline{\boldsymbol{\eta}}_{-1,gt}-\kappa_{j}\right),F_{-\widetilde{\nu}_{gt}}\left(\pi_{1,gt}z+\overline{\mathbf{w}}^{\prime}\overline{\boldsymbol{\pi}}_{-1,gt}-\iota_{j}\right)\right).$
is real analytic in each of its coordinates on $\mathbb{R}$, since the elementary functions are real analytic and the composition of real analytic functions are real analytic. Let $\eta^{*}_{1,gt},\lambda^{*}_{gt},\pi^{*}_{1,gt},\zeta^{*}_{gt}$ be arbitrary parameters in $\mathbb{R}$ with $\eta^{*}_{1,gt},\pi^{*}_{1,gt}>0$ and $\lambda^{*}_{gt},\zeta^{*}_{gt}>0$, $\overline{C}^{*}_{gt}$ an arbitrary copula function, and $F^{*}_{1,gt}$, $F^{*}_{2,gt}$ arbitrary distribution function such that $F^{*}_{l,gt}(x)=0.5$ if and only if $x=0$ for $l=1,2$, and
$\Pr\left(C_{gt}\geq j\mid\mathbf{X}_{gt},\mathbf{z}_{gt}\right)$ $=\overline{C}_{gt}^{*}\left(F^{*}_{1,gt}\left(\overline{\mathbf{X}}_{gt}^{\prime}\overline{\boldsymbol{\eta}}^{*}_{gt}-\kappa_{j}\right),F^{*}_{2,gt}\left(\overline{\mathbf{Z}}_{gt}^{\prime}\overline{\boldsymbol{\pi}}^{*}_{gt}-\iota_{j}\right)\right).$
Furthermore, given any $\mathbf{w}\in\mathbb{R}^{k-1}$, the composite function
$(x,z)\mapsto{c}^{*}_{\mathbf{w},gt}(x,z):={\overline{C}}^{*}_{gt}\left(F^{*}_{1,gt}\left({{\eta}^{*}_{1,gt}x+\overline{\mathbf{w}}^{\prime}\overline{{\boldsymbol{\eta}}}^{*}_{-1,gt}-\kappa_{j}}\right),F^{*}_{2,gt}\left({\pi}^{*}_{1,gt}z+\overline{\mathbf{w}}^{\prime}\overline{{\boldsymbol{\pi}}}^{*}_{-1,gt}-\iota_{j}\right)\right)$
is real analytic in each of its coordinates on $\mathbb{R}$. Now, fix $\mathbf{w}\in\mathcal{W}_{gt}$. Then, the composite function
$(x,z)\mapsto{c}_{\mathbf{w},gt}(x,z)-{c}^{*}_{\mathbf{w},gt}(x,z)$
is real analytic in each of its coordinates, and, by Assumption 7, is zero on the Cartesian product of two open intervals $I_{1}\times I_{2}$. This in turn implies that for any $x_{0}\in I_{1}$ and $z_{0}\in I_{2}$, the functions
$z\mapsto{c}_{\mathbf{w},gt}(x_{0},z)-{c}^{*}_{\mathbf{w},gt}(x_{0},z),$
$x\mapsto{c}_{\mathbf{w},gt}(x,z_{0})-{c}^{*}_{\mathbf{w},gt}(x,z_{0}),$
are real analytic and zero on the open intervals $I_{2}$ and $I_{1}$, respectively. Focusing on the former function for identification of $\boldsymbol{\eta}_{gt}$, the identity theorem for real analytic functions guarantees that ${c}_{\mathbf{w},gt}(x_{0},\cdot)$ and ${c}^{*}_{\mathbf{w},gt}(x_{0},\cdot)$ coincide everywhere on $\mathbb{R}$. In fact, the limits $\lim_{z\rightarrow\infty}{c}_{\mathbf{w},gt}(x_{0},z)$ and $\lim_{z\rightarrow\infty}{c}^{*}_{\mathbf{w},gt}(x_{0},z)$ exist, are unique and coincide:
$\lim_{z\rightarrow\infty}{c}_{\mathbf{w},gt}(x_{0},z)=$ $\lim_{z\rightarrow\infty}\overline{C}_{gt}\left(F_{-\widetilde{\varepsilon}_{gt}}\left(\eta_{1,gt}x_{0}+\overline{\mathbf{w}}^{\prime}\overline{\boldsymbol{\eta}}_{-1,gt}-\kappa_{j}\right),F_{-\widetilde{\nu}_{gt}}\left(\pi_{1,gt}z+\overline{\mathbf{w}}^{\prime}\overline{\boldsymbol{\pi}}_{-1,gt}-\iota_{j}\right)\right)$
$=$ $F_{-\widetilde{\varepsilon}_{gt}}\left(\eta_{1,gt}x_{0}+\overline{\mathbf{w}}^{\prime}\overline{\boldsymbol{\eta}}_{-1,gt}-\kappa_{j}\right)$
$=\lim_{z\rightarrow\infty}{c}^{*}_{\mathbf{w},gt}(x_{0},z)=$ $\lim_{z\rightarrow\infty}\overline{C}^{*}_{gt}\left(F^{*}_{1,gt}\left(\eta^{*}_{1,gt}x_{0}+\overline{\mathbf{w}}^{\prime}\overline{\boldsymbol{\eta}}^{*}_{-1,gt}-\kappa_{j}\right),F^{*}_{2,gt}\left(\pi^{*}_{1,gt}z+\overline{\mathbf{w}}^{\prime}\overline{\boldsymbol{\pi}}^{*}_{-1,gt}-\iota_{j}\right)\right)$
$=$ $F^{*}_{1,gt}\left(\eta^{*}_{1,gt}x_{0}+\overline{\mathbf{w}}^{\prime}\overline{\boldsymbol{\eta}}^{*}_{-1,gt}-\kappa_{j}\right).$ (27)
Because such identity holds for every $x_{0}\in I_{1}$, given Assumption 7 and by applying again the identity theorem for real analytic functions, the real analytic composite function
$x\mapsto F_{-\widetilde{\varepsilon}_{gt}}\left(\eta_{1,gt}x+\overline{\mathbf{w}}^{\prime}\overline{\boldsymbol{\eta}}_{-1,gt}-\kappa_{j}\right)-F^{*}_{1,gt}\left(\eta^{*}_{1,gt}x+\overline{\mathbf{w}}^{\prime}\overline{\boldsymbol{\eta}}^{*}_{-1,gt}-\kappa_{j}\right)$
is then zero on $\mathbb{R}$. In fact, as $\eta_{1,gt}\neq 0$, the individual functions also coincide at the point $x^{0}=\frac{\kappa_{j}-\overline{\mathbf{w}}^{\prime}\overline{\boldsymbol{\eta}}_{-1,gt}}{\eta_{1,gt}}$, such that $F_{-\widetilde{\varepsilon}_{gt}}$ is evaluated at its median:
$0.5=F_{-\widetilde{\varepsilon}_{gt}}\left(0\right)=F^{*}_{1,gt}\left(\eta^{*}_{1,gt}\left(\frac{\kappa_{j}-\overline{\mathbf{w}}^{\prime}\overline{\boldsymbol{\eta}}_{-1,gt}}{\eta_{1,gt}}\right)+\overline{\mathbf{w}}^{\prime}\overline{\boldsymbol{\eta}}^{*}_{-1,gt}-\kappa_{j}\right),$
However, it must be the case that $F^{*}_{1,gt}$ is evaluated at its median, too:
$\eta^{*}_{1,gt}\left(\frac{\kappa_{j}-\overline{\mathbf{w}}^{\prime}\overline{\boldsymbol{\eta}}_{-1,gt}}{\eta_{1,gt}}\right)+\overline{\mathbf{w}}^{\prime}\overline{\boldsymbol{\eta}}^{*}_{-1,gt}-\kappa_{j}=0.$
Rearrange the terms:
$\left(\kappa_{j}-\overline{\mathbf{w}}^{\prime}\overline{{\boldsymbol{\eta}}}_{-1,gt}\right)\eta^{*}_{1,gt}+\eta_{1,gt}\overline{\mathbf{w}}^{\prime}\overline{\boldsymbol{\eta}}^{*}_{-1,gt}=\kappa_{j}\eta_{1,gt}$
As the above exposition applies to each $1\leq j\leq 2$, we can build a system of linear equations:
$-\overline{\mathbf{w}}^{\prime}\overline{{\boldsymbol{\eta}}}_{-1,gt}a+\eta_{1,gt}b$ $=0,$
$\left(1-\overline{\mathbf{w}}^{\prime}\overline{{\boldsymbol{\eta}}}_{-1,gt}\right)a+\eta_{1,gt}b$ $=\eta_{1,gt}.$
This system necessarily has full rank:
$\operatorname{rank}\left(\left(\begin{array}[]{ll}-\overline{\mathbf{w}}^{\prime}\overline{{\boldsymbol{\eta}}}_{-1,gt}&\eta_{1,gt}\\ 1-\overline{\mathbf{w}}^{\prime}\overline{{\boldsymbol{\eta}}}_{-1,gt}&\eta_{1,gt}\end{array}\right)\right)=2.$
Thus, $\left(\eta^{*}_{1,gt},\overline{\mathbf{w}}^{\prime}\overline{\boldsymbol{\eta}}^{*}_{-1,gt}\right)$ is uniquely determined as the solution in $(a,b)$ to the above system. Solving the system, we obtain
$\left(\begin{array}[]{ll}\eta^{*}_{1,gt}\\ \overline{\mathbf{w}}^{\prime}\overline{\boldsymbol{\eta}}^{*}_{-1,gt}\end{array}\right)$ $=\left(\begin{array}[]{ll}-\overline{\mathbf{w}}^{\prime}\overline{{\boldsymbol{\eta}}}_{-1,gt}&\eta_{1,gt}\\ 1-\overline{\mathbf{w}}^{\prime}\overline{{\boldsymbol{\eta}}}_{-1,gt}&\eta_{1,gt}\end{array}\right)^{-1}\left(\begin{array}[]{ll}0\\ {\eta}_{1,gt}\end{array}\right)$
$=-\frac{1}{\eta_{1,gt}}\left(\begin{array}[]{ll}\eta_{1,gt}&-\eta_{1,gt}\\ \overline{\mathbf{w}}^{\prime}\overline{{\boldsymbol{\eta}}}_{-1,gt}-1&-\overline{\mathbf{w}}^{\prime}\overline{{\boldsymbol{\eta}}}_{-1,gt}\end{array}\right)\left(\begin{array}[]{ll}0\\ {\eta}_{1,gt}\end{array}\right)$
$=\left(\begin{array}[]{ll}\eta_{1,gt}\\ \overline{\mathbf{w}}^{\prime}\overline{{\boldsymbol{\eta}}}_{-1,gt}\end{array}\right).$
As a result, $\eta_{1,gt}$ and $\overline{\mathbf{w}}^{\prime}\overline{{\boldsymbol{\eta}}}_{-1,gt}$ are identified. Furthermore, as the above exposition applies to every $w\in\mathcal{W}_{gt}$, and $\overline{\mathcal{W}}_{gt}$ does not lie in a proper linear subspace of $\mathbb{R}^{k}$ by Assumption 7, $\overline{\boldsymbol{\eta}}_{-1,gt}$ is identified. This implies in turn that
$F_{-{\widetilde{\varepsilon}}_{gt}}\left(v\right)=F^{*}_{1,gt}\left(v\right),$
for all $v\in\mathbb{R}$, establishing identification of $F_{-{\widetilde{\varepsilon}}_{gt}}$. Consequently, $\lambda_{gt}$ is identified as the standard deviation of this distribution, as well as $F_{{\varepsilon}_{gt}}$. Identification of $\overline{\boldsymbol{\pi}}_{gt}$, $\zeta_{gt}$ and $F_{-{\nu}_{gt}}$ is analogous.
Now, focusing on $(g,t)=(1,1)$, given the distributional parallel trends, identification of $\overline{\mathbf{x}}^{\prime}\overline{\boldsymbol{\eta}}_{11}$ and $\lambda_{11}$ directly follow from Proposition 1:
$\overline{\mathbf{x}}^{\prime}\overline{\boldsymbol{\eta}}_{11}=\overline{\mathbf{x}}^{\prime}\left(\overline{\boldsymbol{\eta}}_{01}+\frac{\overline{\boldsymbol{\eta}}_{10}-\overline{\boldsymbol{\eta}}_{00}}{\lambda_{00}/\lambda_{01}}\right)\quad\text{and}\quad\lambda_{11}=\frac{\lambda_{01}\lambda_{10}}{\lambda_{00}},$
for all covariate values $\mathbf{x}\in\mathcal{X}_{11}$. As $\overline{\mathcal{X}}_{11}$ does not lie in a proper linear subspace of $\mathbb{R}^{k+1}$ by Assumption 5, $\overline{\boldsymbol{\eta}}_{11}$ is identified as
$\overline{\boldsymbol{\eta}}_{11}=\overline{\boldsymbol{\eta}}_{01}+\frac{\overline{\boldsymbol{\eta}}_{10}-\overline{\boldsymbol{\eta}}_{00}}{\lambda_{00}/\lambda_{01}}.$
Identification of $\overline{\boldsymbol{\pi}}_{11}$ and $\zeta_{gt}$ is analogous:
$\overline{\boldsymbol{\pi}}_{11}=\overline{\boldsymbol{\pi}}_{01}+\frac{\overline{\boldsymbol{\pi}}_{10}-\overline{\boldsymbol{\pi}}_{00}}{\zeta_{00}/\zeta_{01}}\quad\text{ and }\quad\zeta_{11}=\frac{\zeta_{01}\zeta_{10}}{\zeta_{00}}.$
Proof of Proposition 5.
The result follows by applying Proposition 3 for the identification of the distribution of $Y_{11}^{\ast}(0)$, and Assumption 8. ∎
Proof of Proposition 6.
The result follows by applying Proposition 4 for the identification of the marginal distribution of $Y_{11}^{\ast}(0)$, and by using Assumption 9 together with arguments from the proof of Proposition 4 to identify the marginal distribution of $Y_{11}^{\ast}(1)$. ∎
Proof of Proposition 7.
The proof follows from Assumptions 1, 10, and 11 as well as an application of the Delta Method. ∎

### v3

@@
-GL1-02-06-24-2-supplement
+GL1-22-11-24-2-supplement
 Abstract
 In this paper, we develop a Difference-in-Differences model for discrete, ordered outcomes, building upon elements from a continuous Changes-in-Changes model.
-We focus on outcomes derived from self-reported survey data eliciting socially undesirable, illegal, or stigmatized behaviors like tax evasion, substance abuse, or domestic violence, where too many “false zeros”, or more broadly, underreporting are likely.
-We provide characterizations for distributional parallel trends, a concept central to our approach, within a general threshold-crossing model framework.
-In cases where outcomes are assumed to be reported correctly, we propose a framework for identifying and estimating treatment effects across the entire distribution.
-This framework is then extended to modeling underreported outcomes, allowing the reporting decision to depend on treatment status.
-A simulation study documents the finite sample performance of the estimators.
+We focus on outcomes derived from self-reported survey data eliciting socially undesirable, illegal, or stigmatized behaviors like tax evasion or substance abuse, where too many “false zeros”, or more broadly, underreporting are likely.
+We start by providing a characterization for parallel trends within a general threshold-crossing model.
+We then propose a partial and point identification framework for different distributional treatment effects when the outcome is subject to underreporting.
 Applying our methodology, we investigate the impact of recreational marijuana legalization for adults in several U.S.
@@
 A prominent example of this more comprehensive approach is the Changes-in-Changes (CiC) model developed by Athey and Imbens (2006).
-This model is particularly valued for its invariance to transformations of the outcome and its intuitive interpretation, and has gained popularity in recent applied (e.g., Gregory and Zierahn, 2022; Assunçao et al., 2023) and theoretical (e.g., Huber et al., 2022; Ghanem et al., 2023) literature.
-The key assumption is that the distribution of unobservables governing counterfactual outcomes remains stable over time within the treated and comparison groups.
-This paper presents a generalization of the the commonly used Logit or Probit DiD method to discrete, ordered outcomes, building upon elements from a continuous CiC model.
-Central to our approach is the concept of distributional parallel trends, which we leverage for the latent outcome variable within a general threshold-crossing model class, using it as our key identification assumption.
-Our modeling approach is distinct from and nonnested within the discrete CiC model of Athey and Imbens (2006), even though it does share the features of the continuous CiC approach.
-As a second generalization, we consider discrete outcomes that are potentially underreported, nesting the “misreporting-to-zero” or “false-zero” scenario as a special case.
-This phenomenon of underreporting is prevalent in survey data on discrete ordered outcomes of count, ordinal, or binary nature, especially when socially undesirable, illegal, or stigmatized behaviors and actions are elicited.
+This paper presents a CiC model for discrete, ordered outcomes that are potentially underreported, nesting the “misreporting-to-zero” or “false-zero” scenario as a special case.
+Underreporting is prevalent in survey data on discrete ordered outcomes of count, ordinal, or binary nature, especially when socially undesirable, illegal, or stigmatized behaviors and actions are elicited, posing significant challenges for both model identification as well as estimation and inference.
 Common examples include the individual consumption of “social bads” such as marijuana (cannabis) and alcohol, and sensitive issues like abortion or domestic violence incidents.
-These behaviors or events often lead to a significant proportion of “false zeros” in survey data, due to individuals’ reluctance to report such incidents truthfully.
 [^1] For example, Brown et al.
@@
 (2019) for other examples.
-Such partial observability, without additional constraints on the Data Generating Process, poses significant challenges for both (point) identification and estimation.
 The paper makes several contributions to the literature.
-Firstly, for the case where outcomes are assumed to be reported correctly, we propose a framework for identification and estimation of their causal effects across the entire distribution.
-Specifically, we consider a general ordered choice model with strictly stationary error terms, and demonstrate that the distributional parallel trends assumption is a natural choice for discrete outcomes.
-We provide two characterizations of distributional parallel trends on the latent outcomes within this threshold-crossing framework: we establish that distributional parallel trends hold if and only if a “parallel trends”-type assumption holds for the functional median (mean) and the scaling parameters.
-The latter holds if and only if these two functional parameters admit a distinct form of separability in group and time.
-These characterizations are of independent interest and provide a direct pathway to two key point identification results of the paper.
-In particular, we show that distributional parallel trends point identifies the counterfactual distribution of the latent outcome in a parametric ordered choice setup when no other covariates are present.
-We also demonstrate that, when additional covariates are present, the counterfactual distribution are point identified even in the absence of distributional assumptions on the error term, i.e., when the model is semiparametric.
-Our results nest the commonly used Logit or Probit DiD for discrete responses (see Puhani, 2012; Wooldridge, 2023), which feature in the applied literature either as main specification for binary (Meyer and Rosenbaum, 2001; Finkelstein, 2007; Mayer et al., 2014) and ordered choice outcomes (e.g., Sequeira, 2016), or as “robustness” specification (e.g., Marie and Zölitz, 2017; Hvide and Jones, 2018; Hoekstra and Sloan, 2022).
-These models typically draw on the conventional parallel trends that exclusively affect the mean parameters.
-Moreover, when additional covariates are present, our identification result can directly be translated into an estimation procedure that accommodates the discrete nature of the outcome.
-This is in contrast to Athey and Imbens (2006), where the proposed estimators are not directly tailored to these types of outcomes and do not allow for structural estimation.
-As a second contribution, this paper extends the DiD framework to account for underreporting, a scenario leading to partially observable data, where the true outcomes are correctly recorded if and only if they are both actual and reported (see Poirier, 1980).
-We generalize the conventional ”false-zero” reporting by allowing survey respondents to report values that are below the actual level, not limited to zero only.
-We impose restrictions neither on the dependence structure between the unobservables governing the outcome and reporting the decisions, nor on their evolution across groups and over time.
-In this respect, our approach is distinct from the literature on binary choice models with two-sided misreporting (or misclassification) that typically assume the independence between the unobservables (see Hausman et al., 1998; Lewbel, 2000a).
-Addressing the identification and estimation challenges posed by such partially observed outcome models, we propose a set of sufficient conditions for the identification of causal parameters.
-Our setup, though akin to the parametric bivariate choice model with fully observed outcomes (e.g., Mourifié and Méango, 2014; Han and Vytlacil, 2017; Han and Lee, 2019), is notably more complex in its analysis due to the partial observability.
-We demonstrate that having an exclusion restriction in each equation (outcome and reporting), where the excluded variables are continuous, along with the real analytic condition on the joint distribution of unobservables, is sufficient for point identification of the distributional treatment effects on the treated.
+Firstly, we consider a general ordered choice model with underreporting and with strictly stationary error terms along the lines of the continuous CiC model, and provide a characterization of parallel trends on the latent outcomes within a threshold-crossing model framework: we establish that this type of “Changes-in-Changes” setup holds if and only if a parallel trends-type assumption holds for the functional mean (median) and the scaling parameters.
+In the binary case and under distributional assumptions, this characterization can nest the commonly used Logit or Probit DiD for binary responses (see Puhani, 2012; Wooldridge, 2023) as these models typically draw on the conventional parallel trends that exclusively affect the mean parameters.
+As a second contribution, we derive sharp, nonparametric bounds for the (counter)factual outcome distribution of the treatment group after treatment in the presence of underreporting.
+We show that, without further restrictions, the bounds on the counterfactual distribution remain uninformative even if an instrumental exclusion restriction for the consumption is available.
+By contrast, leveraging the CiC structure alongside with an instrumental exclusion restriction and an upper bound on the cumulative misreporting probability, we demonstrate that more informative, tight bounds can potentially be achieved.
+Our bounds extend the partial identification results of Athey and Imbens (2006) for the discrete outcome case to underreporting.
+Moreover, the results also complement recent results of Mondal and Wang (2024) for binary outcome models under misreporting by adapting their findings to an ordered choice setup.
+A suitable estimation and inference procedure for the nonparametric bounds is discussed in the supplementary material.
+As a third contribution, we complement the above findings with a set of point identification results for a more structural approach that explicitly models the consumption and reporting behavior, respectively.
+Specifically, this model generalizes the conventional “false-zero” reporting by allowing survey respondents to report values that are below the actual level, not limited to zero only.
+We impose restrictions neither on the dependence structure between the unobservables governing the outcome and reporting decisions, nor on their evolution across groups and over time.
+In this respect, our approach is distinct from the literature on binary choice models with two-sided misreporting (or misclassification) that typically assume the independence between the unobservables (see Hausman et al., 1998; Lewbel, 2000).
+Our approach also accounts for the potential endogeneity of reporting behavior with respect to treatment status, a frequent concern in observational studies (Arulampalam et al., 2017).
+We address the identification and estimation challenges posed by the partially observed outcome model by proposing a set of sufficient conditions for the identification of differnt causal parameters of interest for consumption and reporting behavior.
+Our setup, though akin to the parametric bivariate choice model with fully observed outcomes (e.g., Mourifié and Méango, 2014; Han and Vytlacil, 2017), is notably more complex in its analysis due to the partial observability.
+We demonstrate that having an exclusion restriction in either the reporting or the consumption equation that exhibits some continuous variation, along with a real analytic condition on the joint distribution of unobservables, is sufficient for point identification of the distributional treatment effects on the treated for consumption or the reporting behavior, respectively.
 This real analytic assumption, a strong form of smoothness imposed on the copula and the marginal distribution functions of the unobservables, has been used for identification in various contexts including quantile sample selection (Arellano and Bonhomme, 2017), triangular models with endogenous binary treatment (Pereda-Fernández, 2023), and random coefficient discrete choice models (Fox et al., 2012; Fox and Gandhi, 2016).
-While not being “constructive” and lacking a direct pathway for estimation, it permits the use of a wide range of well-established marginal distribution and copula functions, thereby offering practical flexibility.
-On the other hand, two exclusion restrictions are required to identify the parameters in both outcome and reporting equations.
-This is common in multi-index outcome models (e.g., Donkers and Schafgans, 2008; Lewbel and Lin, 2022), but may be hard to find in some applications.
-However, we note that this requirement for two exclusion restrictions can be relaxed if the focus is solely on the parameters of one of these equations.
-In addition, we discuss possibilities to device specification tests that allow to check the functional specification in practice.
-Crucially, our approach accounts for the potential endogeneity of reporting behavior with respect to treatment status, a frequent concern in observational studies (e.g., Arulampalam et al., 2017).
-This aspect is particularly pertinent in our empirical application, where we examine how reporting behavior may shift in response to treatment.
-The theoretical identification results are complemented by an extensive Monte Carlo study that investigates the finite sample performance across various exclusion restriction scenarios and distributional choices related to the real analytic copula function.
-Corroborating that parameters of the partially observable outcome model remain unidentified without exclusion restrictions as postulated by Poirier (1980), our study reveals a relatively sizable and slow-decreasing bias in the Maximum Likelihood (ML) estimator for certain parameters, when a single continuous excluded variable is employed in only one of the equations.
-This pattern persists across both Frank and Clayton copula functions.
-Crucially, the inclusion of two excluded variables, one per equation, substantially improves finite sample performance.
-This leads to a rapid decline in the bias of the ML estimator for all model parameters, offering practical guidance for applying these methods in real-world scenarios.
-Finally, as an empirical contribution, we investigate the impact of recreational marijuana legalization for adults in several U.S.
+Finally, as an empirical application, we investigate the impact of recreational marijuana legalization for adults in several U.S.
 states on the short-term consumption behavior of 8th-grade high-school students.
@@
 The paper is organized as follows.
-Section 2 outlines the model setup, detailing the observation rule for self-reported outcomes and defining the distributional treatment effects of interest.
+Section 2 outlines the model setup.
 Section 3 introduces the concept of distributional parallel trends, providing necessary and sufficient conditions for them to hold within a general threshold-crossing framework.
-Section 4 addresses point identification of model parameters when outcomes are correctly reported, starting with the scenario without covariates and then moving to a semiparametric model with unspecified error term distributions.
-Section 5 extends the framework to account for underreporting, establishing point identification in a corresponding semiparametric model.
-Section 6 is dedicated to point identification of distributional treatment effects on the treated, both with and without misreporting.
-Section 7 discusses both parametric and semiparametric estimation for these distributional treatment effects.
-Section 8 presents extensive Monte Carlo simulations to assess the finite sample behavior of the ML estimator in the scenario of underreported outcomes, covering various exclusion restriction scenarios and contrasting different index and error variances.
+Section 4 extends the framework to account for underreporting, establishing our main identification results.
 Section 5 empirically investigates the effects of recreational marijuana legalization on the consumption behavior of 8th grade high-school students.
 Section 6 concludes.
-Finally, an alternative characterization of distributional parallel trends can be found in Appendix A, while Appendix B contains the proofs of all propositions in the paper.
-A proof for the result of Appendix A, details on the calculation of the standard errors, an alternative estimation procedure as well as additional simulation and empirical results can be found in the supplementary material.
+The supplementary material contains auxiliary results for the partial identification model (Section S.1), the proofs of all formal results (Section S.2), an outline of estimation and inference for the nonparametric bounds as well as for the semiparametric approach modeling the reporting behavior explicitly (Section S.3), a Monte Carlo study where we examine the finite sample behavior of the estimators under different sets of exclusion restrictions (Section S.4), additional empirical results (Section S.5).
+Finally, the supplement also provides a different Copula decomposition as a basis to construct different treatment effects (Section S.6).
 2 Setup
+In order to introduce the CiC framework for the ordered choice model and the parallel trends characterization in the next section, we start by considering the setup in the absence of underreporting.
 Let $Y(1)$ and $Y(0)$ denote the potential outcomes for individuals that take on a finite number of $J+1$ values (in a pre-specified period of time) with and without treatment, respectively.
-In light of the empirical application of Section 9, we will refer to $Y(1)$ and $Y(0)$ as potential consumption levels in the following.
-Also, for simplicity, we will focus on the case of two groups and two time periods exclusively.
+In light of the empirical application of Section 5, we will refer to $Y(1)$ and $Y(0)$ as “true” potential consumption levels in the following.
+For simplicity, we will focus on the case of two groups and two time periods exclusively.
 The extension to several (pre-treatment) periods is immediate at the cost of more cumbersome notation.
 We define $G\in\{0,1\}$ as the group membership indicator (where group 1 is the treated group) and $T\in\{0,1\}$ as the time period indicator (where period 1 is after the treatment).
-The binary treatment indicator for individuals is given by $D=G\cdot T$, so that $D=1$ if $(G,T)=(1,1)$, and $D=0$ otherwise.
-To streamline the notation, we adopt the shorthand as in Athey and Imbens (2006):
+The binary treatment indicator for individuals is given by $D=G\cdot T$, so that $D=1$ if and only if $(G,T)=(1,1)$, and $D=0$ otherwise.
+Our primary goal lies in identifying distributional treatment effects on the treated (DTTs) for the outcome of interest, such as the effects of marijuana legalization on different levels of drug consumption among adolescents residing in treated states.
+That is, in the presence of additional covariates $\mathbf{X}$, we aim to identify:
+[…]
+for different levels $j$ and values $\mathbf{x}$ in the support of $\mathbf{X}$ (conditional on $G=1$ and $T=1$), i.e., values from the covariate distribution in the treated group after the treatment.
+To distinguish the current focus from alternative treatment effects on reporting which will be introduced in Section 4, we use the superscript “c” to indicate “consumption”.
+Furthermore, we may also be interested in the unconditional distributional treatment effect on the treated.
+This can be obtained by “marginalization” with respect to the covariate distribution in the treated group after the treatment:
+[…]
+where $\mathcal{X}_{11}$ denotes the support of $\mathbf{X}$ given $G=1$ and $T=1$, and ${F}_{\mathbf{X}|D=1}(\cdot)$ the covariate distribution function for the treated group after treatment.
+In the absence of underreporting, the first quantity on the right-hand side RHS of (1) is generally nonparametrically identified from the data, while the second term represents a counterfactual distribution that cannot be identified without further assumptions.
+However, this condition is rarely satisfied when survey data on sensitive behaviors are involved, as such data are often subject to underreporting.
+Consequently, the presence of underreporting further complicates the identification of both quantities.
+Given our primary interest is modeling self-reported ordered outcomes, we adopt a general threshold-crossing model framework for the counterfactual outcomes by introducing a latent outcome variable $Y^{*}(d)$, $d\in\{0,1\}$:
+$Y(d)=\left\{\begin{array}[]{lll}0&\text{ if }\quad\kappa_{0}\leq Y^{*}(d)\leq\kappa_{1}\\ 1&\text{ if }\quad\kappa_{1}<Y^{*}(d)\leq\kappa_{2}\\ 2&\text{ if }\quad\kappa_{2}<Y^{*}(d)\leq\kappa_{3},\end{array}\right.$
+[…]
+3 Parallel Trends for Ordered Outcomes
+As is customary in the DiD or the synthetic control literature, we start by modeling the potential outcomes $Y(0)$.
+To streamline the notation, we adopt the shorthand from Athey and Imbens (2006) in what follows and write:
 $Y_{gt}(d)\stackrel{{\scriptstyle}}{{\sim}}Y(d)\mid G=g,T=t$
-for each $d\in\{0,1\},g\in\{0,1\}$ and $t\in\{0,1\}$, where $\sim$ denotes “is distributed as”. [^2] We will adopt the same notation for all other random variables in this paper.
-As mentioned in the introduction, a prevalent issue in analyzing sensitive behaviors, such as the consumption of “social bads,” is the presence of “false zeros” or, more broadly, underreporting.
-Individuals often conceal their actual participation in such behaviors when responding to retrospective surveys (Nguimkeu et al., 2019).
-This leads to a partially observable outcome model where a specific outcome value is observed in the data if and only if the actual outcome corresponds to that value and is reported truthfully (cf.
-Poirier, 1980).
-To formalize this idea, let $R(d)$ denote the potential “reporting intentions” for $d=0,1$.
-Importantly, note that this formulation allows reporting behavior to change in response to the treatment status $d$.
-For example, changes in the legalization of recreational marijuana usage for adults may alter the perceived harmfulness or stigma associated with its consumption, thereby also affecting individuals’ willingness to report truthfully.
-Using similar shorthand as before, we continue to denote
-$R_{gt}(d)\stackrel{{\scriptstyle}}{{\sim}}R(d)\mid G=g,T=t$
-for each $d\in\{0,1\},g\in\{0,1\}$ and $t\in\{0,1\}$.
-Here, we interpret $R_{gt}(d)$ as the (potential) level up to which an individual commits to reporting her consumption truthfully.
-The potential outcome of interest, $C_{gt}(d)$, $d=0,1$, is then given by
-$C_{gt}(d)=\min\left\{Y_{gt}(d),R_{gt}(d)\right\}.$ (1)
-Thus, respondents will truthfully report their consumption $Y_{gt}(d)$ if and only if it is less than or equal to the reporting intention $R_{gt}(d)$.
-For instance, if $Y_{gt}(0)=2$ and $R_{gt}(0)=1$, then, in the absence of treatment, the respondent would only report consumption at level 1 even though the actual level is 2.
-It is worth noting that this formulation of underreporting behavior in (1) allows for more flexibility than standard “false zero” reporting setups, where respondents may underreport a positive outcome level exclusively as a “false zero” (see Brown et al., 2018; Greene et al., 2018; Nguimkeu et al., 2019).
-Building on this framework and (1), we may also define the actual observed outcome $C_{gt}$ in the standard manner as a linear combination of the potential outcomes $C_{gt}(0)$ and $C_{gt}(1)$ as:
-$C_{gt}=C_{gt}(1)D+C_{gt}(0)(1-D),$
-where $C_{gt}(d)=\min\{Y_{gt}(d),R_{gt}(d)\}$ for $d\in\{0,1\}$.
-The above definition bears similarities with the binary selection setup in Klein et al.
-(2015).
-However, additional complications in the identification arise due to the partial observability of the outcome of interest, stemming from the unknown reporting decision (see also Poirier (1980)).
-That is, unlike a selection setup where the selection indicator is fully observed, $R_{gt}(d)$ can be only partially inferred for each $d$ from the observed joint outcome of individual consumption and reporting decisions.
-More specifically, both the actual consumption and reporting decisions are observable if and only if individuals have reported the highest consumption level.
-Throughout the paper, our primary interest lies in identifying the distributional treatment effects on the treated (DTTs), focusing primarily on the impact of marijuana legalization on the drug consumption of adolescents residing in treated states.
-That is, in the presence of additional covariates $\mathbf{X}_{gt}$ (recall that the subscripts highlight the possible dependence on the group membership $G=g$ and time period $T=t$), we aim to identify:
-$\tau_{11}^{c}(j\mid\mathbf{x}):=$ $\Pr\left(Y_{11}(1)=j\mid\mathbf{X}_{11}=\mathbf{x}\right)-\Pr\left(Y_{11}(0)=j\mid\mathbf{X}_{11}=\mathbf{x}\right),$ (2)
-for different levels $j$ and values $\mathbf{x}$ in the support of $\mathbf{X}_{11}$, i.e.
-values from the covariate distribution in the treated group after the treatment.
-Alternatively, we may also be interested in the unconditional distributional treatment effect on the treated.
-This can, depending on whether variation in $\mathbf{X}_{gt}$ plays essential roles in the identification arguments (see Section 4), either be directly addressed with an unconditional version of (2) or be obtained by “marginalization” with respect to the covariate distribution in the treated group after the treatment:
-[…]
-where $\mathcal{X}_{gt}$ denotes the support of $\mathbf{X}_{gt}$ and ${F}_{\mathbf{X}_{gt}}(\cdot)$ the covariate distribution function for a given group $g$ and time period $t$.
-Another relevant policy parameter of interest is the distributional treatment effect for the treated on the reporting intention:
-$\tau^{r}_{11}(j\mid\mathbf{z}):=$ $\Pr\left(R_{11}(1)=j\mid\mathbf{Z}_{11}=\mathbf{z}\right)-\Pr\left(R_{11}(0)=j\mid\mathbf{Z}_{11}=\mathbf{z}\right),$ (3)
-for different levels $j$ and values $\mathbf{z}$ in the support of $\mathbf{Z}_{11}$, where $\mathbf{Z}_{11}$ denotes a covariate vector nonnested within $\mathbf{X}_{11}$, see Section 5.
-The parameter in (3) provides insights into the possible changes in the probabilities of reporting on individuals in response to the treatment.
-This may also shed light on how factors like stigma and shame influence the reporting behavior.
-Regardless of the estimand of interest, note that the first quantity on the RHS of (2) is generally nonparametrically identified from the data when $Y_{11}$ is observable, whereas the second probability is a counterfactual distribution that cannot be identified without further assumptions.
-As is customary in the DiD or the synthetic control literature, we therefore focus on modeling the counterfactual outcome $Y_{11}(0)$.
-To do so, for each group and time period $(g,t)\in\{0,1\}$, we consider a general threshold-crossing model framework for the counterfactual outcomes and introduce a latent outcome variable $Y^{*}_{gt}(0)$:
-$Y_{gt}(0)=\left\{\begin{array}[]{lll}0&\text{ if }\quad\kappa_{0}\leq Y^{*}_{gt}(0)\leq\kappa_{1}\\ 1&\text{ if }\quad\kappa_{1}<Y^{*}_{gt}(0)\leq\kappa_{2}\\ 2&\text{ if }\quad\kappa_{2}<Y^{*}_{gt}(0)\leq\kappa_{3},\end{array}\right.$
-[…]
-$R_{gt}(0)=\left\{\begin{array}[]{lll}0&\text{ if }\quad\iota_{0}\leq R^{*}_{gt}(0)\leq\iota_{1}\\ 1&\text{ if }\quad\iota_{1}<R^{*}_{gt}(0)\leq\iota_{2}\\ 2&\text{ if }\quad\iota_{2}<R^{*}_{gt}(0)\leq\iota_{3},\end{array}\right.$ (7)
-where $\iota_{0}<\iota_{1}<\iota_{2}<\iota_{3}$, setting again $\iota_{0}=-\infty$ and $\iota_{3}=\infty$, and (w.l.o.g.) $\iota_{1}=0$ and $\iota_{2}=1$.
-3 Characterizing Distributional Parallel Trends
-We start by considering the setup in the absence of misreporting, i.e.
-$C_{gt}(d)=Y_{gt}(d)$ for $d\in\{0,1\}$ almost surely, and provide characterizations of the distributional parallel trends in a general threshold-crossing setup where the unobservable distribution is invariant over time within each group.
-Specifically, recall that $\mathbf{X}_{gt}$ denotes a vector of covariates for a specific group membership $g$ and time period $t$, and let $\mathcal{X}_{gt}$ and $d_{X_{gt}}$ denote the corresponding support and dimension of that vector, respectively.
+for each $d\in\{0,1\},g\in\{0,1\}$ and $t\in\{0,1\}$, where $\sim$ denotes “is distributed as”.
+We will adopt the same notation for all other random variables as well.
+Thus, let $\mathbf{X}_{gt}$ denote a vector of covariates for a specific group membership $g$ and time period $t$, and let $\mathcal{X}_{gt}$ and $d_{X_{gt}}$ denote the corresponding support and dimension of that vector, respectively.
 Then, letting +:=$\mathbb{R}_{+}:=\{x\in\mathbb{R}\mid x>0\}$, we define the latent consumption $Y_{gt}^{\ast}(0)$ to take the following form:
 Assumption 1.
+For every $(g,t)\in\{0,1\}\times\{0,1\}$:
 $Y^{*}_{gt}(0)=\eta_{gt}(\mathbf{X}_{gt})+\lambda_{gt}(\mathbf{X}_{gt})\varepsilon_{gt},$ (2)
-where $\eta_{gt}:\mathbb{R}^{d_{X_{gt}}}\rightarrow\mathbb{R}$ and $\lambda_{gt}:\mathbb{R}^{d_{X_{gt}}}\rightarrow\mathbb{R}_{+}$ are unknown, real-valued functions, $\varepsilon_{gt}$ is strictly stationary within each group given $\mathbf{X}_{gt}$ with zero median (or mean) and unit variance, and has an unknown, continuous, and strictly monotonic conditional cumulative distribution function (CDF) $F_{\varepsilon_{gt}\mid\mathbf{X}_{gt}}$.
-Here, $\eta_{gt}(\mathbf{X}_{gt})$ can be viewed as the functional median (mean) parameter of the conditional distribution $Y_{gt}^{\ast}(0)\mid\mathbf{X}_{gt}$, and $\lambda_{gt}(\mathbf{X}_{gt})$ as the functional scaling parameter of that distribution.
+where $\eta_{gt}:\mathbb{R}^{d_{X_{gt}}}\rightarrow\mathbb{R}$ and $\lambda_{gt}:\mathbb{R}^{d_{X_{gt}}}\rightarrow\mathbb{R}_{+}$ are unknown, real-valued functions, $\varepsilon_{gt}$ is continuously distributed and strictly stationary within each group given $\mathbf{X}_{gt}$.
+Here, when one w.l.o.g.
+normalizes the mean (median) and variance of $\varepsilon_{gt}$ to zero and one, then $\eta_{gt}(\mathbf{X}_{gt})$ can be viewed as the functional mean (median) parameter of the conditional distribution $Y_{gt}^{\ast}(0)\mid\mathbf{X}_{gt}$, and $\lambda_{gt}(\mathbf{X}_{gt})$ as the functional scaling parameter of that distribution.
 [^4] To clarify the notation, the functions $\eta_{gt}(\cdot)$ and $\lambda_{gt}(\cdot)$ are assumed to be identical to all units within the group $g$ and in period $t$, while the random variables $Y^{*}_{i,gt}(0)$, $\mathbf{X}_{i,gt}$, $\varepsilon_{i,gt}$ can vary across units.
@@
 for any given $g\in\{0,1\}$ and $t\in\{0,1\}$ (provided minor regularity conditions on the moments), the strict stationarity on the other hand postulates that the conditional distribution of $\varepsilon_{gt}$ must not change over time within each group.
-The latter represents a conditional form of Assumption 3.3 in Athey and Imbens (2006) and lies at the heart of both the DiD and the CiC approaches.
-Finally, note that Assumption 1 also nests various (parametric and semiparametric) threshold crossing models (Klein and Sherman, 2002; Klein et al., 2015) as well as nonlinear DiD models (Kim and Lee, 2017; Wooldridge, 2023).
-We next outline the definition of (conditional) distributional parallel trends for the latent outcome variable of interest, $Y^{*}_{11}(0)$, which will be used as key identification assumption throughout the paper:
+The latter represents a conditional form of Assumption 3.3 in Athey and Imbens (2006) and lies at the heart of the CiC approach, which is particularly valued for its invariance to transformations of the outcome and its intuitive interpretation (e.g., Assunçao et al., 2023; Huber et al., 2022).
+Finally, note that Assumption 1 encompasses various (parametric and semiparametric) threshold crossing models (Klein and Sherman, 2002; Klein et al., 2015) and is closely related to nonlinear DiD models (Kim and Lee, 2017; Wooldridge, 2023).
+We next outline the definition of a CiC model for the latent outcome variable of interest, $Y^{*}_{11}(0)$, which will be used as key identification assumption throughout the paper.
+For a generic cumulative distribution function $F:\mathbb{R}\rightarrow[0,1]$ of a random variable $Y$ with support $\mathcal{Y}$, define its “generalized inverse” by
+$F^{-1}:[0,1]\rightarrow\mathbb{R}\cup\left\{-\infty,\infty\right\},\ q\mapsto F^{-1}\left(q\right)$ $:=\inf\left\{y\in\mathcal{Y}\mid q\leq F\left(y\right)\right\},$
+where we adopt the convention $\inf\{\emptyset\}=\infty$ and $\inf\{\mathbb{R}\}=-\infty$.
 Definition 1.
-The (conditional) distributional parallel trends are said to hold if for every covariate value $\mathbf{x}\in\cap_{g,t}\mathcal{X}_{gt}$, there exists an open interval $I\subseteq\mathbb{R}$, such that
-$0<F_{Y^{*}_{11}(0)\mid\mathbf{X}_{11}}(v\mid\mathbf{x})$ $=F_{Y^{*}_{10}(0)\mid\mathbf{X}_{10}}\left(F_{Y^{*}_{00}(0)\mid\mathbf{X}_{00}}^{-1}\left(F_{Y^{*}_{01}(0)\mid\mathbf{X}_{01}}(v\mid\mathbf{x})\mid\mathbf{x}\right)\mid\mathbf{x}\right)<1$ (9)
-for all $v\in I$.
-We provide two characterizations of the distributional parallel trends within our threshold-crossing framework in Proposition 1 below and Proposition A.1 in Appendix A.
-The formula in (9) is akin to the identification result of the (unconditional) continuous outcome CiC model of Athey and Imbens (2006), see Theorem 3.1 of their paper.
-Note that the unconditional distributional parallel trends can be defined similarly if the focus is on the unconditional distribution.
-Therefore, without considering additional covariates, Definition 1 is graphically illustrated in Figure 1.
-Here, the counterfactual distribution is constructed by adding the difference in the distributions for the control group over time at each level $q^{\prime}$ (corresponding to the quantile $y^{\ast}$) to the treatment distribution function at each level $q$ (corresponding to quantile $y^{\ast}$ in the treatment distribution before treatment).
-When $I=\mathbb{R}$, as argued in Athey and Imbens (2006), the quantiles of the counterfactual distribution $Y^{\ast}_{11}(0)$ may be constructed by inverting the equality in (9) for each level $q\in[0,1]$:
-$F_{Y_{11}^{*}(0)}^{-1}(q)=F_{Y_{01}^{*}(0)}^{-1}\left(F_{Y_{00}^{*}(0)}\left(F_{Y_{10}^{*}(0)}^{-1}(q)\right)\right).$
-Note that the class of models implied by the distributional parallel trends above is generally not nested within the parallel trends model set out in (2) of Roth and Sant’Anna (2023).
-Specifically, denoting $F_{Y_{gt}(0)}$ as the CDF of the potential outcome in the absence of the treatment, the latter relies on the identifying assumption:
-$F_{Y_{11}(0)}(y)=F_{Y_{10}(0)}(y)+F_{Y_{01}(0)}(y)-F_{Y_{00}(0)}(y)$ (10)
-for all $y\in\{0,1,2\}$.
-This direct extension of the linear DiD model for the mean to the entire outcome distribution, while intuitive, may give rise to predictions outside the unit interval as the right-hand side (RHS) is not restricted to lie between zero and one.
-On the other hand, one may attempt to place a similar “parallel trends”-type assumption on the CDFs of the latent outcome $Y_{gt}^{*}(0)$:
+$Y^{*}_{gt}(0)$ is said to be generated by a CiC model if for every $\mathbf{x}\in\cap_{g,t}\mathcal{X}_{gt}$, there exists an open interval $I_{\mathbf{x}}\subseteq\mathcal{Y}^{*}_{\mathbf{x},11}(0)\cap\mathcal{Y}^{*}_{\mathbf{x},01}(0)$ such that for all $v\in I_{\mathbf{x}}$:
+$F_{Y^{*}_{10}(0)\mid\mathbf{X}_{10}}^{-1}\left(F_{Y^{*}_{11}(0)\mid\mathbf{X}_{11}}(v\mid\mathbf{x})\mid\mathbf{x}\right)=F_{Y^{*}_{00}(0)\mid\mathbf{X}_{00}}^{-1}\left(F_{Y^{*}_{01}(0)\mid\mathbf{X}_{01}}(v\mid\mathbf{x})\mid\mathbf{x}\right).$ (3)
+We provide two characterizations of this CiC model within our threshold-crossing framework in Proposition 1 below and in Proposition LABEL:PROPS2 of the supplementary material, respectively.
+The formula in (3) closely resembles the unconditional identification result for the CiC model with continuous outcomes by Athey and Imbens (2006), see Theorem 3.1 of their paper.
+A similar formula can also be defined for the unconditional counterfactual distribution when the focus shifts to an unconditional analysis.
+To illustrate Definition 1 without additional covariates, Figure 1 graphically demonstrates the construction of the counterfactual distribution.
+Specifically, the counterfactual distribution is derived by adding the difference in the distributions for the control group over time at each level $q^{\prime}$ (corresponding to the quantile $y^{\ast}$) to the treatment distribution function at each level $q$ (corresponding to quantile $y^{\ast}$ in the treatment distribution before treatment).
+Note that (3) has often been compared to other “parallel trends”-type assumptions on the CDFs of the counterfactual distributions (see, e.g., Remark 4 in Roth and Sant’Anna, 2023).
+One such alternative involves imposing a “parallel trends”-type assumption directly on the CDFs of the latent outcome $Y_{gt}^{*}(0)$:
 $F_{Y^{*}_{11}(0)}(y)=F_{Y^{*}_{10}(0)}(y)+F_{Y^{*}_{01}(0)}(y)-F_{Y^{*}_{00}(0)}(y)$
 for all $y\in\mathbb{R}$.
-However, this may, in turn, incur undesirable restrictions within some standard location-scale families (see, e.g., Example 2 in Roth and Sant’Anna, 2023).
-[^5] Taking the family of normal distributions as an example, such “parallel trends”-type assumption for the CDFs holds only if $Y_{gt}^{*}(0)$ is either strictly stationary over time within each group or stable across groups in each period (but not a mixture of these cases or others).
-In our context, this would consequently restrict the (functional) location and scaling parameters $\eta_{gt}$ and $\lambda_{gt}$ to be identical either across groups or over time.
-As can be demonstrated in Proposition A.1 in Appendix A, our distributional parallel trends generally maintain considerable flexibility in how the time trend, the group membership, and the observed heterogeneity affect the counterfactual outcome.
-In this respect, we argue that this assumption is sensible for outcomes with discrete, ordered responses.
+However, this approach introduces undesirable restrictions within some standard location-scale families (see, e.g., Example 2 in Roth and Sant’Anna, 2023), making it less suitable for modeling discrete outcome.
+For instance, in the family of normal distributions, such a “parallel trends”-type assumption for the CDFs holds only under restrictive conditions - either $Y_{gt}^{*}(0)$ must exhibit strict stationarity over time within each group, or it must remain stable across groups in each period (but not a mixture of these cases or others).
+This would consequently restrict the (functional) location and scaling parameters $\eta_{gt}$ and $\lambda_{gt}$ to be identical either across groups or over time.
+In contrast, our CiC framework for the latent outcome is more flexible and does not impose homogeneity on the model parameters.
+Moreover, it accommodates standard parametric Probit and Logit DiD models with specified distribution for the error term.
+[^5] For example, consider a model $Y^{*}(0)=\mathbf{X}^{\prime}\boldsymbol{\eta}+\alpha T+\beta G+\varepsilon$, where $\varepsilon\sim N(0,1)$.
+It can be readily shown that (3) is satisfied.
+Given the widespread use of of Probit and Logit DiDs in applied work (e.g.
+Mayer et al., 2014; Sequeira, 2016; Hoekstra and Sloan, 2022), this flexibility represents a significant advantage of the CiC framework.
 Figure 1: Stylized CiC Model
@@
 The curve in red on the other hand denotes the marginal distribution function of the latent outcome $Y_{11}^{\ast}(1)$ under treatment.
-Our first result characterizes the necessary and sufficient condition on the functional parameters for distributional parallel trends to hold:
+Before stating the first identification result, we impose a further (conditional) support condition, which, together with the strict stationarity condition on $\varepsilon_{gt}$, implies $\mathcal{Y}_{\mathbf{x},11}\subseteq\mathcal{Y}_{\mathbf{x},01}$:
+Assumption 2.
+For every $\mathbf{x}\in\cap_{g,t}\mathcal{X}_{gt}$ and and $(g,t)\in\{0,1\}\times\{0,1\}$, $\mathcal{Y}^{*}_{\mathbf{x},11}(0)\subseteq\mathcal{Y}^{*}_{\mathbf{x},10}(0),$ and $\mathrm{int}\left(\mathcal{Y}^{*}_{\mathbf{x},11}(0)\right)$ is nonempty.
+The first proposition characterizes the necessary and sufficient condition on the functional parameters for distributional parallel trends to hold:
 Proposition 1.
-Suppose Assumption 1 holds. Then, distributional parallel trends defined in Definition 1 hold if and only if
+Suppose Assumptions 1 and 2 hold. Then, $Y^{*}_{gt}(0)$ is generated by a CiC model as defined in Definition 1 if and only if
 $\eta_{11}(\mathbf{x})-\frac{\lambda_{11}(\mathbf{x})}{\lambda_{10}(\mathbf{x})}\eta_{10}(\mathbf{x})=\eta_{01}(\mathbf{x})-\frac{\lambda_{01}(\mathbf{x})}{\lambda_{00}(\mathbf{x})}\eta_{00}(\mathbf{x})\quad\text{and}\quad\frac{\lambda_{11}(\mathbf{x})}{\lambda_{10}(\mathbf{x})}=\frac{\lambda_{01}(\mathbf{x})}{\lambda_{00}(\mathbf{x})},$ (4)
-for all covariate values $\mathbf{x}\in\cap_{g,t}\mathcal{X}_{gt}$.
-Within the general threshold-crossing model given by Assumption 1, Proposition 1 shows that the distributional parallel trends hold if and only if a weighted “parallel trends”-type assumption holds for the functional median (mean) parameters and a proportional “parallel trends”-type holds for the scaling parameters.
-We note that if the time exclusively affects the median (or mean) parameters, then the results align with the conventional parallel trends, thereby nesting the commonly used Logit or Probit DiD.
-More specifically, as can be seen in the first part of (11), the change in the median (or average) latent outcome for the treated group over time given the covariates, $\eta_{11}(\mathbf{x})-\eta_{10}(\mathbf{x})$, in the absence of treatment, is extrapolated by the change for the control group over the same period, $\eta_{01}(\mathbf{x})-\eta_{00}(\mathbf{x})$, after an appropriate adjustment through the scaling parameters, serving as tailored weighting factors.
+for all $\mathbf{x}\in\cap_{g,t}\mathcal{X}_{gt}$.
+Within the general threshold-crossing model given by Assumption 1, Proposition 1 shows that $Y^{*}_{gt}(0)$ is generated by a CiC model if and only if a weighted “parallel trends”-type assumption holds for the functional mean (median) parameters and a proportional “parallel trends”-type holds for the scaling parameters.
+We note that if time trends exclusively affect the mean (or median) parameters (so $\lambda_{g0}(\mathbf{x})=\lambda_{g1}(\mathbf{x})$), then the results align with the conventional parallel trends, thereby nesting the commonly used Logit or Probit DiD models.
+More specifically, as can be seen in the first part of (4), the change in the mean (or median) latent outcome for the treated group over time given the covariates, $\eta_{11}(\mathbf{x})-\eta_{10}(\mathbf{x})$, in the absence of treatment, is extrapolated by the change for the control group over the same period, $\eta_{01}(\mathbf{x})-\eta_{00}(\mathbf{x})$, after an appropriate adjustment through the scaling parameters, serving as tailored weighting factors.
 Here, the weights play the role of correcting the part of the common time trends that cannot be eliminated by the simple mean difference over time within groups.
-On the other hand, the second part of (11) states that the proportional change in the scaling parameters for the treated group over time, $\frac{\lambda_{11}(\mathbf{x})}{\lambda_{10}(\mathbf{x})}$, in the absence of treatment, must be identical to the proportional change for the treated group over the same periods, $\frac{\lambda_{01}(\mathbf{x})}{\lambda_{00}(\mathbf{x})}$.
-The equality in (11) will serve as a fundamental cornerstone for identifying the counterfactual distributions in Sections 4 and 5.
-It is important to recognize that the results of Proposition 1 remain applicable even when strictly monotonic transformations are applied to the latent outcome $Y_{gt}^{\ast}(0)$, a characteristic inherent to the continuous CiC model (see Athey and Imbens, 2006).
-[^6] This differs from the invariance of parallel trends discussed in Roth and Sant’Anna (2023), which focuses on the average untreated potential outcomes.
-This aspect is particularly relevant for data with discrete ordered outcomes, which are often re-categorized in different ways to verify the “robustness” of empirical findings.
-For instance, in a $J+1$-level scenario (ranging from 00 to $J$), it is common to group the first $j_{1}$ (with $0\leq j_{1}<J$) levels into a new level 0, and then levels $j_{1}+1$ to $j_{2}$ (with $j_{1}+1\leq j_{2}<J$) into a new level 1.
-Such reclassification constitutes, in essence, a strictly monotonic transformation of $Y_{gt}^{\ast}(0)$, thereby preserving the applicability of Proposition 1.777Consider the model with $J+1$ levels: $Y_{gt}(0)=\sum_{j=0}^{J}\mathbb{I}\left[\kappa_{j}<Y^{*}_{gt}(0)\leq\kappa_{j+1}\right]$, where $\kappa_{j}<\kappa_{j+1}$ for all $0\leq j\leq J$, $\kappa_{0}=-\infty$, $\kappa_{J+1}=\infty$, and $\mathbb{I}\left[A\right]$ is the indicator function that equals 1 when $A$ is true and 0 otherwise.
-Reclassifying the first $j_{1}$ levels as level 0 and then levels $j_{1}+1$ to $j_{2}$ as level 1 implements the monotonic transformation $v\mapsto\frac{v-\kappa_{j_{1}}}{\kappa_{j_{2}}-\kappa_{j_{1}}}$ across all terms in the inequalities, setting $\kappa_{j_{1}}=0$ and $\kappa_{j_{2}}=1$.
-Finally, notice that in Appendix A, we establish a second equivalence result to further characterize the class of models that generate the distributional parallel trends under Assumption 1.
+The equality in (4) will serve as a fundamental cornerstone for identifying the counterfactual distributions in Section 4.
+Finally, in the supplementary material, we establish another equivalence result to further characterize the class of CiC models in Definition 1.
 The result helps to clarify the link between the CiC models considered in Athey and Imbens (2006) and our approach.
-While being nonnested within the discrete CiC models with covariates in their paper, our model does share the features of an extended class of their continuous CiC model with modified assumptions accommodating additional covariates, see Appendix A for details.
-4 Model Without Underreporting
-4.1 No Covariates
-In this subsection, we establish a point identification result for the counterfactual distribution of interest in the absence of misreporting and other covariates $\mathbf{X}_{gt}$.
-That is, setting $\eta_{gt}(\cdot)$ and $\lambda_{gt}(\cdot)$ to be constant functions in $\mathbf{X}_{gt}$ (almost surely), the latent consumption $Y_{gt}^{\ast}(0)$ takes the form:
-$Y^{*}_{gt}(0)=\eta_{gt}+\lambda_{gt}\varepsilon_{gt},$ (12)
-where $\eta_{gt}\in\mathbb{R}$ and $0<\lambda_{gt}<\infty$ by Assumption 1 (see also Yamauchi, 2020).
-We readily obtain the following impossibility result for the model without covariates and without “false zeros” in the data:
+4 Underreported Outcomes
+As emphasized in the introduction, a persistent challenge in analyzing sensitive behaviors, such as the consumption of “social bads”, is the presence of “false zeros” or, more broadly, underreporting.
+Individuals often conceal their actual participation in such behaviors when responding to retrospective surveys (Nguimkeu et al., 2019).
+This leads to a partially observable outcome model where a specific outcome value is observed in the data if and only if the actual outcome corresponds to that value and is reported truthfully (cf.
+Poirier, 1980).
+To account for this issue, we introduce potential “reported” consumption, denoted by $C(d)$, $d\in\{0,1\}$, which allows for the possibility that $C(d)\neq Y(d)$ with positive probability.
+The observability rule is then given by $C=D\cdot C(1)+(1-D)\cdot C(0)$.
+It is well established that the CiC model with discrete outcomes, even in the absence of misreporting, does not achieve point identification for the counterfactual distribution $F_{Y_{11}(0)}$ (cf.
+Athey and Imbens, 2006).
+When underreporting is introduced alongside the discrete nature of the outcome, the identification problem becomes even more complex.
+In this section, we address the challenges posed by underreporting in two distinct ways.
+First, in Subsection 4.1, we derive nonparametric bounds for the counterfactual distribution of the treatment group after treatment by leveraging parallel trends and an exclusion restriction.
+However, as we will demonstrate, the expressions for the bounds of the counterfactual distribution are only informative when paired with a relatively small upper bound on the cumulative underreporting probability.
+A different route is taken in Subsection 4.2, where we introduce a semiparametric CiC model for both consumption and reporting decisions, and achieve point identification.
+4.1 Partial Identification
+In order to derive informative and tight bounds for the potential outcome distribution $F_{Y_{11}(d)}$, $d\in\{0,1\}$, we firstly leverage an instrumental exclusion restriction that does not affect consumption behavior directly.
+Assumption 3.
+For any $j\in\{0,1\}$ and $(g,t)\in\{0,1\}\times\{0,1\}$, $d=g\cdot t$, $\mathbf{x}\in\mathcal{X}_{gt}$ and $z\in\mathcal{Z}_{\mathbf{x},gt}$:
+$\operatorname{Pr}\left(Y_{gt}(d)\leq j\mid\mathbf{X}_{gt}=\mathbf{x},Z_{gt}=z\right)=\operatorname{Pr}\left(Y_{gt}(d)\leq j\mid\mathbf{X}_{gt}=\mathbf{x}\right).$
+An example for an instrumental exclusion $Z_{gt}$ in our context is a “survey cooperation” indicator, which measures respondents’ willingness to cooperate during the survey when consumption behavior is elicited.
+This indicator is assumed to have no direct effect on consumption $Y_{gt}(0)$, conditional on all other covariates.
+As is common with instrumental exclusion restrictions, the condition is untestable and the validity of the instrument is subject to the assumptions of the researcher.
+Assumption 4.
+For any $(g,t)\in\{0,1\}\times\{0,1\}$, $d=g\cdot t$, $\mathbf{x}\in\mathcal{X}_{gt}$ and $z\in\mathcal{Z}_{\mathbf{x},gt}$:
+$10<\operatorname{Pr}\left(C_{gt}(d)\leq 0\mid\mathbf{X}_{gt}=\mathbf{x},Z_{gt}=z\right)<\operatorname{Pr}\left(C_{gt}(d)\leq 1\mid\mathbf{X}_{gt}=\mathbf{x},Z_{gt}=z\right)<$ 1
+Assumption 4 is a non-degeneracy condition on the observed potential outcome $C_{gt}(d)$, $d\in\{0,1\}$, ensuring that all levels $j\in\{0,1,2\}$ lie within the conditional support of reported consumption.
+Finally, we combine the exclusion restriction from Assumption 3 with a direct bound on the cumulative misreporting probability.
+The latter becomes necessary because the exclusion restriction alone is insufficient to provide informative bounds for the counterfactual distribution, as demonstrated in Proposition 3 below.
+Assumption 5.
+For any $j\in\{0,1\}$, $(g,t)\in\{0,1\}\times\{0,1\}$, $d=g\cdot t$, $\mathbf{x}\in\mathcal{X}_{gt}$:
+$\operatorname{Pr}\left(C_{gt}(d)\geq j+1\mid Y_{gt}(d)\leq j,\mathbf{X}_{gt}=\mathbf{x}\right)$ $=0,$
+$\operatorname{Pr}\left(C_{gt}(d)\leq j\mid Y_{gt}(d)\geq j+1,\mathbf{X}_{gt}=\mathbf{x}\right)$ $\leq\alpha,$
+where $\alpha\in[0,1]$ is a known constant.
+Assumption 5 explicitly characterizes one-sided misreporting (i.e., underreporting).
+The first part of Assumption 5 states that all kinds of “overreporting probabilities” are zero.
+Specifically, if the true potential consumption level is less than or equal to $j$, individuals cannot overstate their consumption to any level greater than $j$ in their reports.
+The second part of Assumption 5 introduces an explicit upper bound on the “underreporting probabilities”.
+This bound can reflect the researcher’s institutional knowledge or serve as a sensitivity parameter.
+For $\alpha=1$, the assumption imposes no restriction on underreporting, whereas $\alpha=0$ prohibits any underreporting.
+Note that these assumptions generalize the concept of standard misreporting probabilities for binary outcomes (see, e.g., Mondal and Wang, 2024) to cumulative distribution functions in the context of ordered outcomes.
+To state the bounds, we introduce an alternative generalized inverse distribution function, denoted by $F^{(-1)}$ (see Athey and Imbens, 2006).
+Given a generic cumulative distribution function $F:\mathbb{R}\rightarrow[0,1]$ of a random variable $Y$ with support $\mathcal{Y}$, we define
+$F^{(-1)}:[0,1]\rightarrow\mathbb{R}\cup\{-\infty,\infty\},\ q\mapsto F^{(-1)}\left(q\right)$ $:=\sup\left\{y\in\mathcal{Y}\mid q\geq F\left(y\right)\right\},$
+where we adopt the convention $\sup\{\emptyset\}=-\infty$ and $\sup\{\mathbb{R}\}=\infty$. We obtain the following result:
 Proposition 2.
-Suppose Assumption 1 holds and consider the model in (12).
-Also, suppose (unconditional) distributional parallel trends hold as in Definition 1.
-Then, $\eta_{gt}$ and $\lambda_{gt}$ are (point) identified if and only if $F_{\varepsilon_{gt}}^{-1}\left(\mathrm{Pr}\left(Y_{gt}=0\right)\right)$ and $F_{\varepsilon_{gt}}^{-1}\left(\mathrm{Pr}\left(Y_{gt}\leq 1\right)\right)$ are known.
-Proposition 2 highlights the fact that in a setup with no further structure and no additional covariates, point identification of $\eta_{gt}$ and $\lambda_{gt}$ is impossible without the knowledge of the distribution function $F_{\varepsilon_{gt}}$.
-[^8] Unlike the if-and-only-if characterizations for distributional parallel trends in threshold-crossing models in this paper, Yamauchi (2020) establishes a point identification result for the model in (12) where $F_{\varepsilon_{gt}}$ is assumed to be known.
-The paper does, however, not consider the case where $F_{\varepsilon_{gt}}$ is unknown (see Subsection 4.2) or where the outcome is potentially misreported (see Section 5).
-The proposition can be straightforwardly generalized to conditional on covariates.
-We note that, with known $F_{\varepsilon_{gt}}$, the parameters are uniquely determined by the observable quantities
-$\eta_{gt}=\frac{-F_{\varepsilon}^{-1}\left(\mathrm{Pr}\left(Y_{gt}=0\right)\right)}{F_{\varepsilon}^{-1}\left(\mathrm{Pr}\left(Y_{gt}\leq 1\right)\right)-F_{\varepsilon}^{-1}\left(\mathrm{Pr}\left(Y_{gt}=0\right)\right)}\quad\text{and}\quad\lambda_{gt}=\frac{1}{F_{\varepsilon}^{-1}\left(\mathrm{Pr}\left(Y_{gt}\leq 1\right)\right)-F_{\varepsilon}^{-1}\left(\mathrm{Pr}\left(Y_{gt}=0\right)\right)}$
-for $(g,t)\neq(1,1)$. Moreover, Proposition 1 establishes that:
-$\eta_{11}=\eta_{01}+\frac{\eta_{10}-\eta_{00}}{\lambda_{00}/\lambda_{01}}\quad\text{and}\quad\lambda_{11}=\frac{\lambda_{01}\lambda_{10}}{\lambda_{00}}.$
-Nevertheless, employing a parametric stand naturally calls for caution regarding potential misspecification. Therefore, we will ease this distributional assumption in Subsection 4.2 below.
-4.2 Including Covariates
-Next, we consider a semiparametric version of our model for $Y_{gt}^{\ast}(0)$, where the distribution of $\varepsilon_{gt}$ remains unspecified and covariates $\mathbf{X}_{gt}$ enter the model.
-We will exploit variation in the latter to obtain a point identification result for the counterfactual $Y_{11}^{\ast}(0)$ distribution.
-At the same time, in order to facilitate the comparison with more standard ordered choice models from the literature and in light of the extension to misreporting in Section 5, we will place further structure on the functional form of $\eta_{gt}(\cdot)$ and $\lambda_{gt}(\cdot)$, respectively.
-Specifically, for every $(g,t)\in\{0,1\}$, we assume that, with probability one, $\eta_{gt}(\mathbf{X}_{gt})=\overline{\mathbf{X}}_{gt}^{\prime}\overline{\boldsymbol{\eta}}_{gt}$ with $\overline{\mathbf{X}}_{gt}=(1,\mathbf{X}_{gt}^{\prime})^{\prime}$ and $\overline{\boldsymbol{\eta}}_{gt}=(\eta_{c,gt},\boldsymbol{\eta}_{gt}^{\prime})^{\prime}$, while $\lambda_{gt}(\mathbf{X}_{gt})=\lambda_{gt}$.
+Suppose $Y^{*}_{gt}(0)$ is generated by a CiC model, Assumptions 1-5 hold, and $\mathcal{Y}_{\mathbf{x},gt}=\{0,1,2\}$ for $(g,t)\neq(1,1)$.
+Suppose $\alpha\geq\max_{z\in\mathcal{Z}_{gt,\mathbf{x}},(g,t)\neq(1,1)}\left\{\frac{F_{C_{gt}\mid\mathbf{X}_{gt}}\left(j\mid\mathbf{x}\right)-F_{C_{gt}\mid\mathbf{X}_{gt},Z_{gt}}\left(j\mid\mathbf{x},z\right)}{1-F_{C_{gt}\mid\mathbf{X}_{gt},Z_{gt}}\left(j\mid\mathbf{x},z\right)}\right\}$.
+Then, for each $j\in\{0,1,2\}$ and $\mathbf{x}\in\cap_{g,t}\mathcal{X}_{gt}$:
+$F_{Y_{11}(1)\mid\mathbf{X}_{11}}\left(j\mid\mathbf{x}\right)\in\left[L_{11,\alpha}\left(j\mid\mathbf{x}\right),U_{11}\left(j\mid\mathbf{x}\right)\right],$
+$F_{Y_{11}(0)\mid\mathbf{X}_{11}}\left(j\mid\mathbf{x}\right)\in\left[L^{(0)}_{11,\alpha}\left(j\mid\mathbf{x}\right),U^{(0)}_{11,\alpha}\left(j\mid\mathbf{x}\right)\right],$
+where
+$L^{(0)}_{11,\alpha}\left(j\mid\mathbf{x}\right)$ $:=L_{10,\alpha}\left(\left(U_{00}\right)^{(-1)}\left(L_{01,\alpha}\left(j\mid\mathbf{x}\right)\mid\mathbf{x}\right)\mid\mathbf{x}\right),$
+$U^{(0)}_{11,\alpha}\left(j\mid\mathbf{x}\right)$ $:=U_{10}\left(\left(L_{00,\alpha}\right)^{-1}\left(U_{01}\left(j\mid\mathbf{x}\right)\mid\mathbf{x}\right)\mid\mathbf{x}\right),$
+and
+$L_{gt,\alpha}\left(j\mid\mathbf{x}\right)$ $:=\frac{F_{C_{gt}\mid\mathbf{X}_{gt}}\left(j\mid\mathbf{x}\right)-\min\left\{F_{C_{gt}\mid\mathbf{X}_{gt}}\left(j\mid\mathbf{x}\right),\alpha\right\}}{1-\min\left\{F_{C_{gt}\mid\mathbf{X}_{gt}}\left(j\mid\mathbf{x}\right),\alpha\right\}},$ (5)
+$U_{gt}\left(j\mid\mathbf{x}\right)$ $:=\min_{z\in\mathcal{Z}_{gt,\mathbf{x}}}F_{C_{gt}\mid\mathbf{X}_{gt},Z_{gt}}\left(j\mid\mathbf{x},z\right)$ (6)
+for $j\in\{0,1\}$, and $L_{gt,\alpha}\left(j\mid\mathbf{x}\right)=U_{gt}\left(j\mid\mathbf{x}\right):=1$ for $j=2$.
+Moreover, the bounds are sharp when $\max_{z\in\mathcal{Z}_{gt,\mathbf{x}},(g,t)\in\{0,1\}\times\{0,1\}}\left\{\frac{F_{C_{gt}\mid\mathbf{X}_{gt}}\left(j\mid\mathbf{x}\right)-F_{C_{gt}\mid\mathbf{X}_{gt},Z_{gt}}\left(j\mid\mathbf{x},z\right)}{1-F_{C_{gt}\mid\mathbf{X}_{gt},Z_{gt}}\left(j\mid\mathbf{x},z\right)}\right\}\leq\alpha<\min_{(g,t)\in\{0,1\}\times\{0,1\}}F_{C_{gt}\mid\mathbf{X}_{gt}}\left(j\mid\mathbf{x}\right).$
+In addition, if $\alpha<\max_{z\in\mathcal{Z}_{gt,\mathbf{x}},(g,t)\in\{0,1\}\times\{0,1\}}\left\{\frac{F_{C_{gt}\mid\mathbf{X}_{gt}}\left(j\mid\mathbf{x}\right)-F_{C_{gt}\mid\mathbf{X}_{gt},Z_{gt}}\left(j\mid\mathbf{x},z\right)}{1-F_{C_{gt}\mid\mathbf{X}_{gt},Z_{gt}}\left(j\mid\mathbf{x},z\right)}\right\}$, then the set of admissible models is empty.
+The above proposition derives valid bounds for the conditional cumulative distribution function of the counterfactual $Y_{11}(0)$ and the factual $Y_{11}(1)$.
+In addition, note that the bounds in Equation (5) and (6) are also valid bounds for the conditional distribution $Y(0)$ of each sub-group with $(g,t)\neq(1,1)$.
+In fact, when $\alpha=1$ and no bound is placed on underreporting, it is easy to see that
+$\operatorname{Pr}\left(Y_{gt}(0)\leq j\mid\mathbf{X}_{gt}=\mathbf{x}\right)\in\left[0,\min_{z\in\mathcal{Z}_{gt,\mathbf{x}}}F_{C_{gt}\mid\mathbf{X}_{gt},Z_{gt}}\left(j\mid\mathbf{x},z\right)\right]$
+for $(g,t)\neq(1,1)$ and $j\in\{0,1\}$, which in turn implies that the bounds for the counterfactual distribution become
+$F_{Y_{11}(0)\mid\mathbf{X}_{11}}\left(j\mid\mathbf{x}\right)\in\left[0,1\right]$
+for all $\mathbf{x}\in\cap_{g,t}\mathcal{X}_{gt}$.
+On the other hand, observe that Proposition 2 does also not rule out the case of no underreporting when the researcher sets $\alpha=0$.
+To see this, observe that when $\alpha=0$, it holds that:
+$\operatorname{Pr}\left(C_{gt}(d)\leq j\mid Y_{gt}(d)\geq j+1,\mathbf{X}_{gt}=\mathbf{x},Z_{gt}=z\right)=0$
+for all $j\in\{0,1\}$. Under Assumption 3 and the first part of Assumption 4, we can write:
+$F_{C_{gt}\mid\mathbf{X}_{gt},Z_{gt}}(j\mid\mathbf{x},z)$
+$=$ $\operatorname{Pr}\left(Y_{gt}(0)\leq j\mid\mathbf{X}_{gt}=\mathbf{x}\right)$
+$\quad+\underbrace{\operatorname{Pr}\left(C_{gt}(d)\leq j\mid Y_{gt}(d)\geq j+1,\mathbf{X}_{gt}=\mathbf{x},Z_{gt}=z\right)}_{=0}\times\left(1-\operatorname{Pr}\left(Y_{gt}(0)\leq j\mid\mathbf{X}_{gt}=\mathbf{x}\right)\right).$
+However, under Assumption 3 and $\alpha=0$, this equality holds if and only if $F_{C_{gt}\mid\mathbf{X}_{gt},Z_{gt}}(j\mid\mathbf{x},z)=F_{C_{gt}\mid\mathbf{X}_{gt}}(j\mid\mathbf{x})$ for all $\mathbf{x}\in\cap_{g,t}\mathcal{X}_{gt}$ and $j\in\{0,1\}$, which in turn restricts the set of admissible, observable distributions.
+[^6] Note that in this case $\alpha=\max_{z\in\mathcal{Z}_{gt,\mathbf{x}},(g,t)\in\{0,1\}\times\{0,1\}}\left\{\frac{F_{C_{gt}\mid\mathbf{X}_{gt}}\left(j\mid\mathbf{x}\right)-F_{C_{gt}\mid\mathbf{X}_{gt},Z_{gt}}\left(j\mid\mathbf{x},z\right)}{1-F_{C_{gt}\mid\mathbf{X}_{gt},Z_{gt}}\left(j\mid\mathbf{x},z\right)}\right\}=0$ for all $\mathbf{x}\in\cap_{g,t}\mathcal{X}_{gt}$.
+Finally, in Section S.3.1 of the supplementary material, we comment on the possibility to estimate these bounds nonparametrically and to carry out inference via the nonparametric i.i.d.
+bootstrap following the ideas of Fang and Santos (2019) and Poirier and Masten (2020).
+4.2 Point Identification
+As the expressions for the bounds of the counterfactual distribution $Y_{11}(0)$ are only informative when paired with a relatively small upper bound on the cumulative underreporting probability (even the bounds may still be wide in practice), we take a different route in this subsection.
+Specifically, we introduce a semiparametric CiC model that jointly accounts for both consumption and reporting decisions.
+While this approach is more restrictive, as it imposes additional structure on the model, it will allow us to define and to identify new treatment parameters of interest, such as the DTT on the reporting intention.
+To formalize the approach, let $R(d)$, $d\in\{0,1\}$ denote the potential “reporting intention”, analogous to $Y(d)$, with $R(d)$ taking on $J+1$ discrete value values.
+Here, we interpret $R(d)$ as the (potential) level up to which an individual commits to reporting her consumption truthfully.
+The potential outcome of interest, $C(d)$, $d=0,1$, is then given by
+$C(d)=\min\left\{Y(d),R(d)\right\}.$ (7)
+Thus, respondents will truthfully report their consumption $Y(d)$ if and only if it is less than or equal to the reporting intention $R(d)$.
+For instance, if $Y(0)=2$ and $R(0)=1$, then, in the absence of treatment, the respondent would only report consumption at level 1 even though the actual level is 2.
+It is worth noting that this formulation of underreporting behavior in (7) allows for more flexibility than standard “false zero” reporting setups, where respondents may underreport a positive outcome level exclusively as a “false zero” (see Brown et al., 2018; Greene et al., 2018; Nguimkeu et al., 2019).
+The above definition also bears similarities with the binary selection setup in Klein et al.
+(2015).
+However, additional complications in the identification arise due to the partial observability of the outcome of interest, stemming from the unknown reporting decision (see also Poirier (1980)).
+That is, unlike a selection setup where the selection indicator is fully observed, $R(d)$ can be only partially inferred for each $d$ from the observed joint outcome of individual consumption and reporting decisions.
+Importantly, also note that in our setup reporting behavior is allowed to change in response to the treatment status $d$.
+For example, changes in the legalization of recreational marijuana usage for adults may alter the perceived harmfulness or stigma associated with its consumption, thereby also affecting individuals’ willingness to report truthfully.
+The introduction of $R(d)$ allows us to define new policy parameter of interest such as the for the treated on the reporting intention:
+$\tau^{r}(j\mid\mathbf{z}):=$ $\Pr\left(R(1)=j\mid D=1,\mathbf{Z}=\mathbf{z}\right)-\Pr\left(R(0)=j\mid D=1,\mathbf{Z}=\mathbf{z}\right),$ (8)
+for different levels $j$ and values $\mathbf{z}$ in the support of $\mathbf{Z}$ given $G=1$ and $T=1$, where $\mathbf{Z}$ denotes a covariate vector nonnested within $\mathbf{X}$.
+The parameter in (8) provides insights into the possible changes in the probabilities of reporting on individuals in response to the treatment.
+This may also shed light on how factors like stigma and shame influence the reporting behavior.
+In analogy to $Y^{\ast}(0)$, we conclude by defining the latent reporting intention variable $R^{\ast}(0)$:
+$R(0)=\left\{\begin{array}[]{lll}0&\text{ if }\quad\iota_{0}\leq R^{*}(0)\leq\iota_{1}\\ 1&\text{ if }\quad\iota_{1}<R^{*}(0)\leq\iota_{2}\\ 2&\text{ if }\quad\iota_{2}<R^{*}(0)\leq\iota_{3},\end{array}\right.$ (12)
+where $\iota_{0}<\iota_{1}<\iota_{2}<\iota_{3}$, setting again $\iota_{0}=-\infty$ and $\iota_{3}=\infty$, and (w.l.o.g.) $\iota_{1}=0$ and $\iota_{2}=1$.
+4.2.1 Counterfactual Distribution
+Building upon the observation rule outlined in (7) and the latent outcome formulation of $R(0)$ in (12), we consider semiparametric versions of our model for $Y_{gt}^{\ast}(0)$, where the distribution of $\varepsilon_{gt}$ remains unspecified and further structure is placed on the functional form of $\eta_{gt}(\cdot)$ and $\lambda_{gt}(\cdot)$, respectively.
+Specifically, for every $(g,t)\in\{0,1\}$, we assume that $\eta_{gt}(\mathbf{X}_{gt})=\overline{\mathbf{X}}_{gt}^{\prime}\overline{\boldsymbol{\eta}}_{gt}$ with $\overline{\mathbf{X}}_{gt}=(1,\mathbf{X}_{gt}^{\prime})^{\prime}$ and $\overline{\boldsymbol{\eta}}_{gt}=(\eta_{c,gt},\boldsymbol{\eta}_{gt}^{\prime})^{\prime}$, while $\lambda_{gt}(\mathbf{X}_{gt})=\lambda_{gt}>0$.
 Thus:
 $Y_{gt}^{\ast}(0)=\overline{\mathbf{X}}_{gt}^{\prime}\overline{\boldsymbol{\eta}}_{gt}+\lambda_{gt}\varepsilon_{gt}.$ (13)
-Equation (13) is a standard functional specification for ordered choice models (Klein and Sherman, 2002), and allows the median of $Y_{gt}^{\ast}(0)$ to vary with the covariates $\mathbf{X}_{gt}$, but not the scaling parameter.999 Note that interpretation of $\overline{\mathbf{X}}_{gt}^{\prime}\overline{\boldsymbol{\eta}}_{gt}$ as conditional median will be important for the identification of the intercept parameter $\eta_{c,gt}$ (e.g., see Manski, 1988).
-Letting $\overline{\mathcal{X}}_{gt}$ the support of $\overline{\mathbf{X}}_{gt}$, we make the following assumptions:
-Assumption 2.
-For each $g\in\{0,1\}$ and $t\in\{0,1\}$, $\varepsilon_{gt}$ is independent of $\mathbf{X}_{gt}$.
-Assumption 3.
-For each $(g,t)\neq(1,1)$, there exists at least one $k\in\{1,\ldots$, $K\}$ such that $\eta_{k,gt}\neq 0$ and
-$\operatorname{Pr}\left(X_{k,gt}\in\left(a_{1},a_{2}\right)\mid\overline{\mathbf{X}}_{-k,gt}\right)>0$
-with probability one for all open intervals $\left(a_{1},a_{2}\right)\subseteq\mathbb{R}$, where $\overline{\mathbf{X}}_{-k,gt}$ denotes the subvector of $\overline{\mathbf{X}}_{gt}$ excluding the $k$th slope.
-W.l.o.g., let $k=1$.
-Assumption 4.
-For each $g\in\{0,1\}$ and $t\in\{0,1\}$, $\overline{\mathcal{X}}_{gt}$ is not contained in a proper linear subspace of $\mathbb{R}^{k+1}$.
-Assumption 5.
-It holds that $\mathcal{X}_{11}\subseteq\cap_{(g,t)\neq(1,1)}\mathcal{X}_{gt}$.
-Assumption 2 states covariate independence.
-Although this is often challenging to justify in practice, it remains a standard condition in the semiparametric estimation and identification literature, playing a crucial role in the identification argument (cf.
-Athey and Imbens, 2006).
-Assumption 3 on the other hand requires the existence of a “special regressor” with large support, an assumption frequently used in models with discrete responses to ensure point identification of the parameters (see Manski, 1988; Lewbel, 2000b; Han and Vytlacil, 2017; Khan et al., 2021).
-[^10] In principle, other identifying strategies from the semiparametric literature, such as exploiting the differentiability with respect to the continuous variables (e.g., Klein and Spady, 1993), are also possible with more complicated arguments.
-Finally, Assumption 4 is a standard identification assumption ruling out perfect multicollinearity, while Assumption 5 postulates a common support of $\mathbf{X}_{11}$ and the other subgroups.
-We have the following identification result:
-Proposition 3.
-Suppose Assumptions 1-5 hold and consider the model in (13).
-Also, suppose (unconditional) distributional parallel trends hold as in Definition 1.
-Then, for each $g\in\{0,1\}$, $t\in\{0,1\}$ and $v\in\mathbb{R}$, $\overline{\boldsymbol{\eta}}_{gt}$, $\lambda_{gt}$ and $F_{\varepsilon_{gt}}(v)$ are identified.
-In particular, it holds for the counterfactual median and scaling parameters that:
-$\overline{\boldsymbol{\eta}}_{11}=\overline{\boldsymbol{\eta}}_{01}+\frac{\overline{\boldsymbol{\eta}}_{10}-\overline{\boldsymbol{\eta}}_{00}}{\lambda_{00}/\lambda_{01}}\quad\text{and}\quad\lambda_{11}=\frac{\lambda_{01}\lambda_{10}}{\lambda_{00}}.$
-In contrasting the identification result of the counterfactual distribution in Proposition 3 with the point identification in Theorem 4.4 of Athey and Imbens (2006) for the discrete CiC model with covariates, we note that while Proposition 3 imposes more structural constraints such as a linear index, it offers a direct pathway to an estimation procedure that naturally accommodates the discreteness of the outcome, as we will outline in Section 7.
-This presents a clear estimation advantage compared to the methods proposed by Athey and Imbens (2006), which are not specifically tailored to discrete outcomes or the inclusion of covariates.
-However, it is important to stress that, akin to Theorem 4.4 in Athey and Imbens (2006), the identification result in Proposition 3 relies on strong assumptions that, while commonly invoked in the literature on semiparametric discrete outcome models, may sometimes be hard to justify.
-5 Underreported Outcomes
-This section addresses the scenario where outcomes may be underreported.
-It considers cases where, for $d\in\{0,1\}$, $C_{gt}(d)\neq Y_{gt}(d)$ with positive probability, a situation often referred to as “partial observability”.
-In nonlinear models, such discrepancies between true and reported outcomes generally lead to a failure of (point) identification.
-This issue persists even in fully parametric models unless additional exclusion restrictions are assumed (cf.
-Poirier, 1980).
-Building upon the observation rule outlined in (1) and the latent outcome formulation of $R_{gt}(0)$ in (7), and drawing on the model for $Y_{gt}^{\ast}(0)$ presented in (13) from Section 4.2, we are now prepared to introduce a model for $R_{gt}^{\ast}(0)$.
-Similar in structure to the one for $Y_{gt}^{\ast}(0)$, this model incorporates additional covariates $\mathbf{Z}_{gt}$:
+Equation (13) is a standard functional specification for ordered choice models (Klein and Sherman, 2002), and allows the mean of $Y_{gt}^{\ast}(0)$ to vary with the covariates $\mathbf{X}_{gt}$, but not the scaling parameter.
+Similar to the structure of $Y_{gt}^{\ast}(0)$, we also model $R_{gt}^{\ast}(0)$ as a function of additional covariates $\mathbf{Z}_{gt}$:
 $R_{gt}^{\ast}(0)=\overline{\mathbf{Z}}_{gt}^{\prime}\overline{\boldsymbol{\pi}}_{gt}+\zeta_{gt}\nu_{gt},$ (14)
-where $\overline{\mathbf{Z}}_{gt}=(1,\mathbf{Z}_{gt}^{\prime})^{\prime}$ and $\overline{\boldsymbol{\pi}}_{gt}=(\pi_{c,gt},\boldsymbol{\pi}_{gt}^{\prime})^{\prime}$.
+where $\overline{\mathbf{Z}}_{gt}=(1,\mathbf{Z}_{gt}^{\prime})^{\prime}$, $\overline{\boldsymbol{\pi}}_{gt}=(\pi_{c,gt},\boldsymbol{\pi}_{gt}^{\prime})^{\prime}$ and $\zeta_{gt}>0$.
 The covariate vectors $\mathbf{X}_{gt}$ and $\mathbf{Z}_{gt}$ are allowed to share common elements.
-It is required, though, that each vector includes at least one distinct continuous variable not present in the other.
-W.l.o.g., we will assume that all elements, except for the first, in both $\mathbf{X}_{gt}$ and $\mathbf{Z}_{gt}$ are common.
-Consequently, the augmented vectors $\overline{\mathbf{X}}_{gt}$ and $\overline{\mathbf{Z}}_{gt}$ are of the same dimension $k+1$.
+Depending on the treatment parameter of interest, however, it will be required that one vector includes a distinct continuous variable not present in the other.
 Under Assumption 6 below, for each $j\in\{0,1,2\}$, the upper tail probability of consumption in the absence of treatment can be expressed as:
@@
 This suggests that the identification results presented in the aforementioned papers are not directly applicable to our context.
-Specifically, the approach of using a single excluded variable in the second argument, as employed in Theorem 2.1 by Chernozhukov et al.
-(2023) and Theorem 6.1 by Han and Vytlacil (2017), may not be sufficient for point identifying the parameters inside the marginal distribution in our model, nor for the marginal distribution itself.
-This insight is supported by the simulation results in Section 8.
-These simulations suggest that, even within the parametric framework and using fairly large sample sizes, some parameters can still be poorly estimated if only one exclusion restriction is in place.
+This insight is supported by the simulation results in Section LABEL:sec:simulation of the supplementary material.
+These simulations suggest that parameters from the model can, in some cases, still be poorly estimated even if one exclusion restriction is in place.
+In what follows, let $X_{1,gt}$ ($Z_{1,gt}$) denote the first element of the vector $\mathbf{X}_{gt}$ ($\mathbf{Z}_{gt}$), and $\mathbf{X}_{-1,gt}$ ($\mathbf{Z}_{-1,gt}$) the vector containing the remaining ones.
 We make the following assumption:
@@
 Assumption 7.
-(i) For each $g\in\{0,1\}$ and $t\in\{0,1\}$, the variables $X_{1,gt}$ and $Z_{1,gt}$ are excluded from the vectors $\mathbf{Z}_{gt}$ and $\mathbf{X}_{gt}$, respectively, while the other elements of them are common.
+(i) For each $g\in\{0,1\}$ and $t\in\{0,1\}$, (a) the variable $X_{1,gt}$ is excluded from the vector $\mathbf{Z}_{gt}$; (b) the variable $Z_{1,gt}$ is excluded from the vector $\mathbf{X}_{gt}$.
 (ii) For each $(g,t)\neq(1,1)$, $\eta_{1,gt}\neq 0$ and $\pi_{1,gt}\neq 0$ and there exists some Cartesian product of two open intervals $I_{1}\times I_{2}\subseteq\mathbb{R}\times\mathbb{R}$ such that
 $\operatorname{Pr}\left(\left(X_{1,gt},Z_{1,gt}\right)\in I_{1}\times I_{2}\mid\overline{\mathbf{W}}_{gt}\right)>0$
-with probability one, where $\mathbf{{W}}_{gt}$ denote the vector of regressors that are common to both $\mathbf{{X}}_{gt}$ and $\mathbf{{Z}}_{gt}$ and $\overline{\mathbf{W}}_{gt}=\left(1,\mathbf{W}_{gt}^{\prime}\right)^{\prime}$.
-(iii) For each $g\in\{0,1\}$ and $t\in\{0,1\}$, $\mathcal{\overline{X}}_{gt}$ and $\mathcal{\overline{Z}}_{gt}$ are not contained in a proper linear subspace of $\mathbb{R}^{k+1}$.
-Assumption 6(i) closely resembles Assumptions 1-2.
-While the covariate independence is restrictive, as previously discussed, it is important to note that we do not impose any constraints on the correlation between the unobservables $\varepsilon_{gt}$ and $\nu_{gt}$.
+with probability one, where $\overline{\mathbf{W}}_{gt}=\left(1,\mathbf{X}_{-1gt}^{\prime},\mathbf{W}_{-1gt}^{\prime}\right)^{\prime}$.
+(iii) For each $g\in\{0,1\}$ and $t\in\{0,1\}$, $\mathcal{\overline{X}}_{gt}$ and $\mathcal{\overline{Z}}_{gt}$ are not contained in a proper linear subspace of $\mathbb{R}^{d_{X_{gt}}+1}$ and $\mathbb{R}^{d_{Z_{gt}}+1}$, respectively, where $d_{X_{gt}}$ and $d_{Z_{gt}}$ denote the dimension of the vectors $\mathbf{X}_{gt}$ and $\mathbf{Z}_{gt}$.
+While the covariate independence in Assumption 6(i) is restrictive, it is important to note that we do not impose any constraints on the correlation between the unobservables $\varepsilon_{gt}$ and $\nu_{gt}$.
 This allowance caters to scenarios where certain types of “consumers” may be more or less likely to misreport than others.
-This aspect of our model is distinct from the literature on binary choice models with two-sided misreporting (or misclassification) that typically imposes (conditional) independence between the unobservables governing the outcome and reporting decisions (see Hausman et al., 1998; Lewbel, 2000a).
+This aspect of our model is distinct from the literature on binary choice models with two-sided misreporting (or misclassification) that typically imposes (conditional) independence between the unobservables governing the outcome and reporting decisions (see Hausman et al., 1998; Lewbel, 2000).
 On the other hand, the other parts of Assumption 6 impose further conditions on the joint and marginal distributions of the unobservables.
@@
 Specifically, the extrapolation is based on the fact that if two real analytic functions coincide on an open neighborhood, then they coincide everywhere, and in particular also in the limit (if it exists).
-Identification based on such extrapolation is in contrast to “identification at infinity” type arguments, which require the existence of excluded (instrumental) variables with a large support (e.g., Chamberlain, 1986; Lewbel, 2007; D’Haultfoeuille and Maurel, 2013).
+Identification based on such extrapolation is in contrast to “identification at infinity” type arguments, which require the existence of excluded (instrumental) variables with a large support (e.g., Chamberlain, 1986; Lewbel, 2007; Corradi and Gutknecht, 2023).
 Moreover, note that the real analytic property is satisfied by several well-known Copula families such as the Gaussian, Frank, Clayton, Plackett, Joe, or Gumbel Copula families.
-Indeed, while Assumptions 7(i) spells out the exclusion restrictions, Assumptions 7(ii) “only” imposes mild conditions on continuity of the excluded variables $X_{1,gt}$ and $Z_{1,gt}$, without requiring a “special regressor” as often necessitated in the semiparametric identification literature on discrete response models.
-On the other hand, the need for an excluded (continuous) variable in each decision equation is common to other identification work dealing with multiple latent outcome equations, see for instance Lewbel et al.
-(2021).
-Finally, Assumption 7(iii) extends Assumption 5 to the multivariate case.
+Indeed, while Assumptions 7 (i)-(a) as well as (i)-(b) postulate an exclusion restriction for the identification of the parameters of the consumption and the reporting decision, respectively, part (ii) is rather “mild” in its continuity requirement.
+[^7] Note that Assumption 7(i)-(b) together with Assumption 6(i) imply Assumption 3 from Subsection 4.1.
+Specifically, it requires one continuous element in each vector, one of which is the instrumental variable.
+Finally, note that under the assumption that both $\mathbf{X}_{gt}$ and $\mathbf{Z}_{gt}$ contain an excluded, continuous variable, we may actually not only identify counterfactual parameters for either the consumption or the reporting decision as in Proposition 3 below, but for both simultaneously.
 We obtain the following identification result:
 Proposition 3.
-Suppose Assumptions 6-7 hold and consider the models in (13)-(14).
-Also, suppose distributional parallel trends hold as in Definition 1 for $Y_{gt}^{*}(0)$ as well as $R_{gt}^{*}(0)$.
-Then, for each $g\in\{0,1\}$, $t\in\{0,1\}$ and $v\in\mathbb{R}$, $\overline{\boldsymbol{\eta}}_{gt}$, $\overline{\boldsymbol{\pi}}_{gt}$, $\lambda_{gt}$, $\zeta_{gt}$, $F_{-\varepsilon_{gt}}(v)$ and $F_{-\nu_{gt}}(v)$ are identified.
-In particular, it holds for the counterfactual quantities:
+Suppose Assumptions 6 holds and consider the models in (13)-(14).
+Also, suppose that the CiC model from Definition 1 holds for $Y_{gt}^{*}(0)$ as well as for $R_{gt}^{*}(0)$, and that $\mathcal{X}_{11}\subseteq_{(g,t)\neq(1,1)}\mathcal{X}_{gt}$.
+(i) Under Assumption 7 (i)-(a), for each $g\in\{0,1\}$, $t\in\{0,1\}$ and $v\in\mathbb{R}$, $\overline{\boldsymbol{\eta}}_{gt}$, $\lambda_{gt}$, and $F_{-\varepsilon_{gt}}(v)$ are identified.
+In particular, we have that:
 $\overline{\boldsymbol{\eta}}_{11}=\overline{\boldsymbol{\eta}}_{01}+\frac{\overline{\boldsymbol{\eta}}_{10}-\overline{\boldsymbol{\eta}}_{00}}{\lambda_{00}/\lambda_{01}}\quad\text{and}\quad\lambda_{11}=\frac{\lambda_{01}\lambda_{10}}{\lambda_{00}},$ (16)
-and
+(ii) Under Assumption 7 (i)-(b), for each $g\in\{0,1\}$, $t\in\{0,1\}$ and $v\in\mathbb{R}$, $\overline{\boldsymbol{\pi}}_{gt}$, $\zeta_{gt}$, and $F_{-\nu_{gt}}(v)$ are identified.
+In particular, we have that:
 $\overline{\boldsymbol{\pi}}_{11}=\overline{\boldsymbol{\pi}}_{01}+\frac{\overline{\boldsymbol{\pi}}_{10}-\overline{\boldsymbol{\pi}}_{00}}{\zeta_{00}/\zeta_{01}}\quad\text{and}\quad\zeta_{11}=\frac{\zeta_{01}\zeta_{10}}{\zeta_{00}}.$ (17)
-It should be emphasized that the distributional parallel trends solely set constraints on the change of the marginal distributions of counterfactual outcomes over time.
+It should be emphasized that the CiC conditions solely set constraints on the change of the marginal distributions of counterfactual outcomes over time.
 However, these assumptions do not place any constraints on the evolution of the dependence structure between the outcome and reporting decisions.
@@
 This is because the identification based on the real analytic assumption is not “constructive” and does not provide a direct pathway to construct estimators.
-As discussed earlier, however, due to the partially observable nature of the consumption and reporting decisions, it appears difficult to bypass this real analytic assumption without relying on excluded variables with large support.
-Remark 1: Note that, under the assumption of index sufficiency and instrument validity, the above assumption setup not only allows for specification tests of the functional form, but also for the presence of misreporting.
-Specifically, starting with the test of correct functional form and focusing on a specific choice of $(g,t)\neq(1,1)$, let $M\left(j;\mathbf{X}_{gt}^{\prime}\widetilde{\boldsymbol{\eta}}_{gt},\mathbf{Z}_{gt}^{\prime}\widetilde{\boldsymbol{\pi}}_{gt}\right):=\Pr\left(C_{gt}\geq j\mid\mathbf{X}_{gt}^{\prime}\widetilde{\boldsymbol{\eta}}_{gt},\mathbf{Z}_{gt}^{\prime}\widetilde{\boldsymbol{\pi}}_{gt}\right)$ denote the nonparametric tail probability of observed consumption at level $j\in\{0,1,2\}$, where $\widetilde{\boldsymbol{\eta}}_{gt}$ and $\widetilde{\boldsymbol{\pi}}_{gt}$ are the slope coefficient vectors normalized by $\|\boldsymbol{\eta}_{gt}\|$ and $\|\boldsymbol{\pi}_{gt}\|$, respectively.
-Note that this tail probability is nonparametrically identified under Assumption 7 and index sufficiency, see for instance Donkers and Schafgans (2008) or Ahn et al.
-(2018).
-Then, under Assumptions 6 and 7, we may construct a test for the null hypothesis:
-$H_{0}^{C}:\overline{C}_{gt}\left(F_{-{\varepsilon}_{gt}}\left(\frac{\overline{\mathbf{x}}^{\prime}\overline{\boldsymbol{\eta}}_{gt}-\kappa_{j}}{\lambda_{gt}}\right),F_{-{\nu}_{gt}}\left(\frac{\overline{\mathbf{z}}^{\prime}\overline{\boldsymbol{\pi}}_{gt}-\iota_{j}}{\zeta_{gt}}\right)\right)=M\left(j;\mathbf{x}^{\prime}\widetilde{\boldsymbol{\eta}}_{gt},\mathbf{z}^{\prime}\widetilde{\boldsymbol{\pi}}_{gt}\right)$
-for all $j\in\{0,1,2\}$ and $(\mathbf{x},\mathbf{z})\in\mathcal{X}_{gt}\times\mathcal{Z}_{gt}$, where $\mathcal{X}_{gt}$ and $\mathcal{Z}_{gt}$ denote compact subsets of the support of $\mathbf{X}_{gt}$ and $\mathbf{Z}_{gt}$, respectively, vs.
-$H_{A}^{C}:\overline{C}_{gt}\left(F_{-{\varepsilon}_{gt}}\left(\frac{\overline{\mathbf{x}}^{\prime}\overline{\boldsymbol{\eta}}_{gt}-\kappa_{j}}{\lambda_{gt}}\right),F_{-{\nu}_{gt}}\left(\frac{\overline{\mathbf{z}}^{\prime}\overline{\boldsymbol{\pi}}_{gt}-\iota_{j}}{\zeta_{gt}}\right)\right)\neq M\left(j;\mathbf{x}^{\prime}\widetilde{\boldsymbol{\eta}}_{gt},\mathbf{z}^{\prime}\widetilde{\boldsymbol{\pi}}_{gt}\right)$
-for at least some $j\in\{0,1,2\}$ and $(\mathbf{x},\mathbf{z})\in\mathcal{X}_{gt}\times\mathcal{Z}_{gt}$.
-Similarly, we may also use the fact that $M\left(j;\mathbf{X}_{gt}^{\prime}\widetilde{\boldsymbol{\eta}}_{gt},\mathbf{Z}_{gt}^{\prime}\widetilde{\boldsymbol{\pi}}_{gt}\right)$ is nonparametrically identified under our assumptions to test for the presence of misreporting in our semiparametric model setup.
-That is, given index sufficiency and Assumption 7, we may test:
-$H_{0}^{R}:M\left(j;\mathbf{x}^{\prime}\widetilde{\boldsymbol{\eta}}_{gt},\mathbf{z}^{\prime}\widetilde{\boldsymbol{\pi}}_{gt}\right)=\overline{M}\left(j;\mathbf{x}^{\prime}\widetilde{\boldsymbol{\eta}}_{gt}\right)$
-for all $j\in\{0,1,2\}$ and all $(\mathbf{x},\mathbf{z})\in\mathcal{X}_{gt}\times\mathcal{Z}_{gt}$ vs.
-$H_{0}^{R}:M\left(j;\mathbf{x}^{\prime}\widetilde{\boldsymbol{\eta}}_{gt},\mathbf{z}^{\prime}\widetilde{\boldsymbol{\pi}}_{gt}\right)\neq\overline{M}\left(j;\mathbf{x}^{\prime}\widetilde{\boldsymbol{\eta}}_{gt}\right)$
-for at least some $j\in\{0,1,2\}$ and $(\mathbf{x},\mathbf{z})\in\mathcal{X}_{gt}\times\mathcal{Z}_{gt}$, where $\overline{M}\left(j;\mathbf{X}_{gt}^{\prime}\widetilde{\boldsymbol{\eta}}_{gt}\right):=\Pr\left(C_{gt}\geq j\mid\mathbf{X}_{gt}^{\prime}\widetilde{\boldsymbol{\eta}}_{gt}\right)$.
-We comment on suitable test statistics for the above tests in Remark 3 in Section 7.2.
-6 Distributional Treatment Effects
-To identify the DTTs, $\tau_{11}^{c}(j)$, $j\in\{0,1,2\}$, as outlined in (2), we start by considering the scenario where $Y_{gt}(d)=C_{gt}(d)$ almost surely, for $d\in\{0,1\}$.
-Although modeling $Y_{11}(1)$ is per se not essential in DiD settings, it becomes indispensable when addressing scenarios involving misreporting.
-Therefore, to simplify the notation and facilitate the transition to the model with misreporting, we also postulate an analogous model structure for $Y_{11}(1)$ as for $Y_{11}(0)$ without misreporting.
+4.2.2 Distributional Treatment Effects
+To identify the different DTTs, we start by setting up a model for $Y_{11}(1)$ and $R_{11}(1)$, respectively.
+As modeling $Y_{11}(1)$ become indispensable with misreporting, we postulate an analogous model structure for $Y_{11}(1)$ as for $Y_{11}(0)$ without misreporting.
 Specifically, drawing on the model for $Y_{11}^{\ast}(0)$ in (13), we define the latent consumption in the presence of treatment as:
@@
 From this point forward, we will accordingly add the superscript $(0)$ to parameters in the absence of treatment.
-Given the similarity in the structure of $Y_{11}^{\ast}(1)$, we adopt assumptions similar to those for $Y_{11}^{\ast}(0)$ in Section 4.1 above:
-Assumption 8.
-$\varepsilon_{11}^{(1)}$ is strictly stationary within each group with zero median and unit variance, independent of $\mathbf{X}_{11}$, and has an unknown, continuous, and strictly monotonic CDF $F_{\varepsilon_{11}^{(1)}}$.
-Besides, $\eta_{1,11}^{(1)}\neq 0$.
-We readily obtain the following proposition for identification of the conditional DTTs for consumption, in the absence of misreporting:
-Proposition 5.
-Suppose Assumptions 1-5 and 8 hold, and consider the model in (13) and (18).
-Also, suppose distributional parallel trends hold as in Definition 1 for $Y_{gt}^{*}(0)$.
-Then, $\tau_{11}^{c}(j\mid{\mathbf{x}})$ is identified for all $j\in\{0,1,2\}$ and $\mathbf{x}\in\mathcal{X}_{11}$ as:
-$\tau_{11}^{c}(j\mid{\mathbf{x}})=$ $\left(F_{\varepsilon_{11}^{(1)}}\left(\frac{\kappa_{j+1}-\overline{\mathbf{x}}^{\prime}\overline{\boldsymbol{\eta}}^{(1)}_{11}}{\lambda^{(1)}_{11}}\right)-F_{\varepsilon_{11}^{(1)}}\left(\frac{\kappa_{j}-\overline{\mathbf{x}}^{\prime}\overline{\boldsymbol{\eta}}^{(1)}_{11}}{\lambda^{(1)}_{11}}\right)\right)$
-$-\left(F_{\varepsilon_{11}^{(0)}}\left(\frac{\kappa_{j+1}-\overline{\mathbf{x}}^{\prime}\overline{\boldsymbol{\eta}}^{(0)}_{11}}{\lambda_{11}^{(0)}}\right)-F_{\varepsilon_{11}^{(0)}}\left(\frac{\kappa_{j}-\overline{\mathbf{x}}^{\prime}\overline{\boldsymbol{\eta}}_{11}^{(0)}}{\lambda_{11}^{(0)}}\right)\right).$
-Next, we turn to the case where the outcome is subject to underreporting. Drawing on the model for $R_{11}^{\ast}(0)$ in (14), We define:
-$R_{11}^{\ast}(1)=\overline{\mathbf{Z}}_{11}^{\prime}\overline{\boldsymbol{\pi}}_{11}^{(1)}+\zeta_{11}^{(1)}\nu_{11}^{(1)}.$ (19)
-We adopt assumptions similar to those for $Y_{11}^{\ast}(0)$ in Section 4.2 above:
+Drawing on the model for $R_{11}^{\ast}(0)$ in (14), we define analogously:
+$R_{11}^{\ast}(1)=\overline{\mathbf{Z}}_{11}^{\prime}\overline{\boldsymbol{\pi}}_{11}^{(1)}+\zeta_{11}^{(1)}\nu_{11}^{(1)},$ (19)
+and adopt the following assumption:
 Assumption 8.
 $-\varepsilon_{11}^{(1)}$ and $-\nu_{11}^{(1)}$ are strictly stationary within each group with zero median and unit variance, jointly independent of $(\mathbf{X}_{11}^{\prime},\mathbf{Z}_{11}^{\prime})^{\prime}$, and have unknown, real analytic and strictly monotonic CDFs $F_{-\varepsilon_{11}^{(1)}}$ and $F_{-\nu_{11}^{(1)}}$, respectively.
-Besides, $\eta_{1,11}^{(1)}\neq 0$ and $\pi_{1,11}^{(1)}\neq 0$ and there exists some Cartesian product of two open intervals $I_{1}\times I_{2}\subseteq\mathbb{R}\times\mathbb{R}$ such that
-$\operatorname{Pr}\left(\left(X_{1,11},Z_{1,11}\right)\in I_{1}\times I_{2}\mid\overline{\mathbf{W}}_{11}\right)>0$
-with probability one, where $\overline{\mathbf{W}}_{11}$ is defined as in Assumption 7.
 Also, the copula function $\overline{C}_{11}^{(1)}(\cdot,\cdot)$ defined below is strictly monotonic and real analytic in each of its coordinates:
 $\Pr\left(C_{11}(1)\geq j\mid\mathbf{X}_{11},\mathbf{Z}_{11}\right)$ $=\overline{C}_{11}^{(1)}\left(F_{-{\varepsilon_{11}^{(1)}}}\left(\frac{\overline{\mathbf{X}}_{11}^{\prime}\overline{\boldsymbol{\eta}}_{11}^{(1)}-\kappa_{j}}{\lambda_{11}^{(1)}}\right),F_{-{\nu}_{11}^{(1)}}\left(\frac{\overline{\mathbf{Z}}_{11}^{\prime}\overline{\boldsymbol{\pi}}_{11}^{(1)}-\iota_{j}}{\zeta_{11}^{(1)}}\right)\right).$
+Finally, $\eta_{1,11}^{(1)}\neq 0$ and $\pi_{1,11}^{(1)}\neq 0$.
 We have the following result for identification of the conditional DTTs for consumption and reporting, respectively, in the presence of misreporting:
 Proposition 4.
-Suppose Assumptions 6-7 and 9 hold, and consider the model in (13)-(14) and (18)-(19).
-Also, suppose distributional parallel trends hold as in Definition 1 for $Y_{gt}^{*}(0)$ as well as $R_{gt}^{*}(0)$.
-Then, $\tau_{11}^{c}(j)$ and $\tau_{11}^{r}(j)$ are identified for all $j\in\{0,1,2\}$, $\mathbf{x}\in\mathcal{X}_{11}$ and $\mathbf{z}\in\mathcal{Z}_{11}$ as:
-$\tau_{11}^{c}(j\mid{\mathbf{x}})=$ $\left(F_{-\varepsilon_{11}^{(1)}}\left(\frac{\overline{\mathbf{x}}^{\prime}\overline{\boldsymbol{\eta}}^{(1)}_{11}-\kappa_{j+1}}{\lambda^{(1)}_{11}}\right)-F_{-\varepsilon_{11}^{(1)}}\left(\frac{\overline{\mathbf{x}}^{\prime}\overline{\boldsymbol{\eta}}^{(1)}_{11}-\kappa_{j}}{\lambda^{(1)}_{11}}\right)\right)$
+Suppose Assumptions 6, 7, and 8 hold, and consider the model in (13)-(14) and (18)-(19).
+Also, suppose the CiC model from Definition 1 holds for $Y_{gt}^{*}(0)$ as well as $R_{gt}^{*}(0)$.
+Then, for all $j\in\{0,1,2\}$, $\mathbf{x}\in\mathcal{X}_{11}$ and $\mathbf{z}\in\mathcal{Z}_{11}$:
+(i) Under Assumption 7 (i)-(a), $\tau^{c}(j)$ is identified as:
+$\tau^{c}(j\mid{\mathbf{x}})=$ $\left(F_{-\varepsilon_{11}^{(1)}}\left(\frac{\overline{\mathbf{x}}^{\prime}\overline{\boldsymbol{\eta}}^{(1)}_{11}-\kappa_{j+1}}{\lambda^{(1)}_{11}}\right)-F_{-\varepsilon_{11}^{(1)}}\left(\frac{\overline{\mathbf{x}}^{\prime}\overline{\boldsymbol{\eta}}^{(1)}_{11}-\kappa_{j}}{\lambda^{(1)}_{11}}\right)\right)$
 $-\left(F_{-\varepsilon_{11}^{(0)}}\left(\frac{\overline{\mathbf{x}}^{\prime}\overline{\boldsymbol{\eta}}^{(0)}_{11}-\kappa_{j+1}}{\lambda_{11}^{(0)}}\right)-F_{-\varepsilon_{11}^{(0)}}\left(\frac{\overline{\mathbf{x}}^{\prime}\overline{\boldsymbol{\eta}}^{(0)}_{11}-\kappa_{j+1}}{\lambda_{11}^{(0)}}\right)\right).$
-$\tau_{11}^{r}(j\mid{\mathbf{z}})=$ $\left(F_{-\nu_{11}^{(1)}}\left(\frac{\overline{\mathbf{z}}^{\prime}\overline{\boldsymbol{\pi}}^{(1)}_{11}-\iota_{j+1}}{\zeta^{(1)}_{11}}\right)-F_{-\nu_{11}^{(1)}}\left(\frac{\overline{\mathbf{z}}^{\prime}\overline{\boldsymbol{\pi}}^{(1)}_{11}-\iota_{j}}{\zeta^{(1)}_{11}}\right)\right)$
+(ii) Under Assumption 7 (i)-(b), $\tau^{r}(j)$ is identified as:
+$\tau^{r}(j\mid{\mathbf{z}})=$ $\left(F_{-\nu_{11}^{(1)}}\left(\frac{\overline{\mathbf{z}}^{\prime}\overline{\boldsymbol{\pi}}^{(1)}_{11}-\iota_{j+1}}{\zeta^{(1)}_{11}}\right)-F_{-\nu_{11}^{(1)}}\left(\frac{\overline{\mathbf{z}}^{\prime}\overline{\boldsymbol{\pi}}^{(1)}_{11}-\iota_{j}}{\zeta^{(1)}_{11}}\right)\right)$
 $-\left(F_{-\nu_{11}^{(0)}}\left(\frac{\overline{\mathbf{z}}^{\prime}\overline{\boldsymbol{\pi}}^{(0)}_{11}-\iota_{j+1}}{\zeta_{11}^{(0)}}\right)-F_{-\nu_{11}^{(0)}}\left(\frac{\overline{\mathbf{z}}^{\prime}\overline{\boldsymbol{\pi}}^{(0)}_{11}-\iota_{j}}{\zeta_{11}^{(0)}}\right)\right).$
@@
 This effect is characterized by the difference between potential self-reported outcome distributions under treatment and control, where the counterfactual copula function is involved.
-7 Estimation
-In this section, we discuss the estimation of the DTTs $\tau_{11}^{c}(j)$.
-For each observational unit $i$, let $G_{i}$ and $T_{i}$ denote binary indicators of group membership (with $G_{i}=1$ indicating being in the treated group) and time period (with $T_{i}=1$ indicating being after the treatment), respectively.
-Following Sant’Anna and Zhao (2020), we write the observed outcomes and the covariates as $(C_{it},\mathbf{X}_{it}^{\prime},\mathbf{Z}_{it}^{\prime})^{\prime}$, where the subscript $t$ denotes the time period.
-This allows to accommodate various sampling frameworks, including repeated cross-section sampling, where observations are drawn conditional on a specific group $G_{i}=g$ and time period $T_{i}=t$, and panel sampling that draws observation pairs conditional on $G_{i}=g$, see Assumption 10 below.
-Consequently, we define $C_{i}=C_{i1}T_{i}+C_{i0}(1-T_{i})$ and $Y_{i}=Y_{i1}T_{i}+Y_{i0}(1-T_{i})$, with regressors defined in a similar manner.
-It is important to recall that in the absence of “false zeros”, we observe $Y_{it}=C_{it}$ almost surely.
-7.1 Estimation Without Misreporting
-We start with the estimation of the counterfactual distribution in the absence of misreporting, as described in Section 4, focusing initially on a parametric model without covariates detailed in Section 4.1.
-Recalling that $\kappa_{0}=-\infty$, $\kappa_{1}=0$, $\kappa_{2}=1$, $\kappa_{3}=+\infty$, we consider a sample $\{C_{i},G_{i},T_{i}\}_{i=1}^{n}$ with the properties specified in Assumption 10 below.
-Also, suppose Assumptions 1 and 8 hold and consider the models in (12) and (18) (no covariates for the latter), where the distributions of $\varepsilon_{gt}^{(d)}$ are known for all $d=g\cdot t$, $g\in\{0,1\}$ and $t\in\{0,1\}$.
-To avoid notational complexity, we assume that the error term distributions are invariant across groups, time and treatment statuses, i.e., $F_{\varepsilon_{gt}^{(d)}}=F_{\varepsilon}$.
-We define $\boldsymbol{\theta}^{(d)}_{gt}:=\left(\eta^{(d)}_{gt},\lambda^{(d)}_{gt}\right)^{\prime}$ and aggregate these into $\boldsymbol{\theta}:=\left(\boldsymbol{\theta}^{(0)^{\prime}}_{00},\boldsymbol{\theta}^{(0)^{\prime}}_{01},\boldsymbol{\theta}^{(0)^{\prime}}_{10},\boldsymbol{\theta}^{(1)^{\prime}}_{11}\right)^{\prime}$.
-The ML estimator of $\boldsymbol{\theta}$ is formulated as follows:
-$\widehat{\boldsymbol{\theta}}:=$ $\arg\min_{\boldsymbol{\theta}\in\mathcal{B}}\sum_{i=1}^{n}\sum_{j=0}^{2}\mathbb{I}\left[Y_{i}=j\right]\left[\sum_{(g,t)\neq(1,1)}\mathbb{I}\left[G_{i}=g,T_{i}=t\right]\ln\left(F_{\varepsilon}\left(\frac{\kappa_{j+1}-\eta_{gt}^{(0)}}{\lambda_{gt}^{(0)}}\right)-F_{\varepsilon}\left(\frac{\kappa_{j}-\eta_{gt}^{(0)}}{\lambda_{gt}^{(0)}}\right)\right)\right.$
-$+\left.\mathbb{I}\left[G_{i}=1,T_{i}=1\right]\ln\left(F_{\varepsilon}\left(\frac{\kappa_{j+1}-\eta^{(1)}_{11}}{\lambda^{(1)}_{11}}\right)-F_{\varepsilon}\left(\frac{\kappa_{j}-\eta^{(1)}_{11}}{\lambda^{(1)}_{11}}\right)\right)\right].$ (20)
-where $\mathcal{B}$ denotes the parameter space, $\mathbb{I}\left[A\right]$ is the indicator function that equals 1 when $A$ is true and 0 otherwise.
-This step allows for the construction of DTT estimates, $\tau_{11}^{c}(j)$ for $j\in\{0,1,2\}$:
-$\widehat{\tau}^{c}_{11}(j):=\left[F_{\varepsilon}\left(\frac{\kappa_{j+1}-\widehat{\eta}^{(1)}_{11}}{\widehat{\lambda}^{(1)}_{11}}\right)-F_{\varepsilon}\left(\frac{\kappa_{j}-\widehat{\eta}^{(1)}_{11}}{\widehat{\lambda}^{(1)}_{11}}\right)\right]-\left[F_{\varepsilon}\left(\frac{\kappa_{j+1}-\widehat{\eta}^{(0)}_{11}}{\widehat{\lambda}^{(0)}_{11}}\right)-F_{\varepsilon}\left(\frac{\kappa_{j}-\widehat{\eta}^{(0)}_{11}}{\widehat{\lambda}^{(0)}_{11}}\right)\right]$ (21)
-where
-$\widehat{\eta}^{(0)}_{11}:=\widehat{\eta}^{(0)}_{01}+\frac{\widehat{\eta}^{(0)}_{10}-\widehat{\eta}^{(0)}_{00}}{\widehat{\lambda}^{(0)}_{00}/\widehat{\lambda}^{(0)}_{01}}\quad\text{and}\quad\widehat{\lambda}^{(0)}_{11}:=\frac{\widehat{\lambda}^{(0)}_{01}\widehat{\lambda}^{(0)}_{10}}{\widehat{\lambda}^{(0)}_{00}}.$
-For the inference on the treatment parameters $\tau_{11}^{c}(j)$, it is crucial to derive the asymptotic distribution of the ML estimator. To do so, we impose the following assumptions:
-Assumption 10.
-It holds that either (i) conditional on $G_{i}=g$ and $T_{i}=t$ with $g,t\in\{0,1\}$, $\{C_{i}\}_{i=1}^{n}$ is an i.i.d.
-random sample from the corresponding subpopulation or (ii) conditional on $G_{i}=g$ with $g\in\{0,1\}$, $\{C_{i0},C_{i1}\}_{i=1}^{n}$ is an i.i.d.
-random sample from the corresponding subpopulation, either $g=0$ or $g=1$.
-Assumption 11.
-The ML estimator of $\boldsymbol{\theta}$ defined in (7.1) satisfies a Central Limit Theorem:
-$\sqrt{n}\left(\widehat{\boldsymbol{\theta}}-\boldsymbol{\theta}\right)\stackrel{{\scriptstyle d}}{{\rightarrow}}N\left(\boldsymbol{0},\boldsymbol{\Sigma}\right),$
-where $\boldsymbol{\Sigma}$ is a positive definite matrix.
-The first part (i) of Assumption 10 addresses the case where repeated cross-section data is available, while part (ii) covers the case of panel data.
-While Assumption 10 only accommodates the “conditional” sampling scheme where group membership and time are fixed, the assumption could readily be extended to allow for other schemes where $G_{i}$ and $T_{i}$ are random (see Sant’Anna and Zhao, 2020).
-Assumption 11 on the other hand is a high-level condition, valid under the stated model assumptions as well as certain primitive conditions on the Hessian and the population shares of each group.
-In particular, it implies that $n_{gt}/n\rightarrow K$ (with $0<K<\infty$) as $n\rightarrow\infty$ for each $g\in\{0,1\}$, $t\in\{0,1\}$, where $n_{gt}$ denotes the number of observations in the group $g$ and time period $t$.
-Proposition 7.
-Suppose Assumptions 10-11 hold. Then, for each $j\in\{0,1,2\}$:
-$\sqrt{n}\left(\widehat{\tau}^{c}_{11}(j)-\tau_{11}^{c}(j)\right)\stackrel{{\scriptstyle d}}{{\rightarrow}}N\left(\boldsymbol{0},\frac{\partial\tau_{11}^{c}(j)}{\partial{{\boldsymbol{\theta}}}}^{\prime}\boldsymbol{\Sigma}\frac{\partial\tau_{11}^{c}(j)}{\partial{{\boldsymbol{\theta}}}}\right),$
-where the definition of $\frac{\partial\tau_{11}^{c}(j)}{\partial{{\boldsymbol{\theta}}}}$ is given in Section LABEL:sec:derivatives of the supplementary material.
-The components of the asymptotic variance can be readily estimated.
-Specifically, the asymptotic variance matrix $\boldsymbol{\Sigma}$ can be replaced by a standard heteroskedasticity robust estimator, denoted as $\widehat{\boldsymbol{\Sigma}}_{R}$, while the derivatives $\frac{\partial\tau_{11}^{c}(j)}{\partial{{\boldsymbol{\theta}}}}$ can be substituted with a consistent estimator $\widehat{\partial\tau_{11}^{c}(j)/\partial{{\boldsymbol{\theta}}}}$.
-Remark 2: Note that the identification argument for the semiparametric model from Subsection 4.2 can in principle be used to construct a plug-in estimator of the DTTs.
-Here, we provide an informal outline of the steps.
-Section LABEL:sec:proofEstimation of the supplementary material considers the estimation procedure for the DTTs more formally.
-Introducing notations for the rescaled and recentered quantities, let ${\boldsymbol{\eta}}_{gt}^{(d)^{*}}:=\frac{{\boldsymbol{\eta}}_{gt}^{(d)}}{\left\|{\boldsymbol{\eta}}_{gt}^{(d)}\right\|}$, $\kappa_{j,gt}^{(d)^{*}}:=\frac{\kappa_{j}}{\left\|{\boldsymbol{\eta}}_{gt}^{(d)}\right\|}$, $\widetilde{\varepsilon}_{gt}^{(d)}:=\eta_{c,gt}^{(d)}+\lambda_{gt}^{(d)}\varepsilon_{gt}^{(d)}$ and $\widetilde{\varepsilon}_{gt}^{(d)^{*}}:=\frac{\widetilde{\varepsilon}_{gt}^{(d)}}{\left\|{\boldsymbol{\eta}}_{gt}^{(d)}\right\|}$ for all $d=g\cdot t$, $g\in\{0,1\}$ and $t\in\{0,1\}$.
-For simplicity, we will assume that $\eta_{c,gt}^{(d)}=\eta_{c}$ and $\lambda_{gt}^{(d)}=\lambda$ for all $(g,t)\in\{0,1\}$ and $d\in\{0,1\}$, and comment on the case where these parameters differ at the end of this remark.
-Given the new notation, the model in (13) can thus be reformulated as:
-$Y_{gt}(d)=$ $\sum_{j=0}^{2}j\cdot\mathbb{I}\left[\kappa_{j,gt}^{(d)^{*}}<{\mathbf{X}}_{gt}^{\prime}{\boldsymbol{\eta}}^{(d)^{*}}_{gt}+\widetilde{\varepsilon}^{(d)^{*}}_{gt}\leq\kappa_{j+1,gt}^{(d)^{*}}\right].$ (22)
-The rescaled slope coefficient vector ${\boldsymbol{\eta}}^{(d)^{*}}_{gt}$ in (22) can be estimated using various rank estimators available in the literature (e.g., Sherman, 1993).
-Subsequently, the threshold parameter $\kappa_{2,gt}^{(d)^{*}}=1/\|{\boldsymbol{\eta}}^{(d)}_{gt}\|$ can be estimated using the $\sqrt{n}_{gt}$-consistent threshold estimator for semiparametric ordered choice models proposed by Klein and Sherman (2002).
-This yields an estimator for the scaling factor as
-$\widehat{\|\boldsymbol{\eta}^{(d)}_{gt}\|}:=1/\widehat{\kappa_{2,gt}^{(d)^{*}}}.$
-Consequently, the estimator for the unscaled parameter vector becomes:
-$\widehat{\boldsymbol{\eta}}^{(d)}_{gt}:=\widehat{\boldsymbol{\eta}}_{gt}^{(d)^{*}}\widehat{\|\boldsymbol{\eta}^{(d)}_{gt}\|}.$
-After estimating all parameters, we recover the distributions of $\widetilde{\varepsilon}_{gt}^{(d)}$ using, e.g., a standard nonparametric kernel estimator:
-$\widehat{F}_{\widetilde{\varepsilon}_{gt}^{(d)}}(v):=\widehat{\mathrm{E}}\left[\mathbb{I}\left[Y_{i}\leq 0\right]\mid\mathbf{X}_{i}^{\prime}\widehat{\boldsymbol{\eta}}_{gt}^{(d)}=-v,G_{i}=g,T_{i}=t\right].$
-With estimates for all groups and time periods at hand, and under strict stationarity for the counterfactual distribution from Assumption 1, we can construct the estimated conditional DTTs for $j\in\{0,1,2\}$ and $\mathbf{x}\in\mathcal{X}_{11}$ as outlined in Proposition 6:
-$\widehat{\tau}^{c}_{11}(j\mid\mathbf{x})$
-::=:= $\mathbb{I}\left[(\kappa_{j+1}-\mathbf{x}^{\prime}\widehat{\boldsymbol{\eta}}^{(1)}_{11}),(\kappa_{j}-\mathbf{x}^{\prime}\widehat{\boldsymbol{\eta}}^{(1)}_{11})\in\Psi_{11}^{(1)};\;(\kappa_{j+1}-\mathbf{x}^{\prime}\widehat{\boldsymbol{\eta}}^{(0)}_{11}),(\kappa_{j}-\mathbf{x}^{\prime}\widehat{\boldsymbol{\eta}}^{(0)}_{11})\in\Psi_{11}^{(0)}\right]$
-×$\times\left(\left[\widehat{F}_{{\widetilde{\varepsilon}}_{11}^{(1)}}\left(\kappa_{j+1}-\mathbf{x}^{\prime}\widehat{\boldsymbol{\eta}}^{(1)}_{11}\right)-\widehat{F}_{{\widetilde{\varepsilon}}_{11}^{(1)}}\left(\kappa_{j}-\mathbf{x}^{\prime}\widehat{\boldsymbol{\eta}}^{(1)}_{11}\right)\right]-\left[\widehat{F}_{{\widetilde{\varepsilon}}_{11}^{(0)}}\left(\kappa_{j+1}-\mathbf{x}^{\prime}\widehat{\boldsymbol{\eta}}^{(0)}_{11}\right)-\widehat{F}_{{\widetilde{\varepsilon}}_{11}^{(0)}}\left(\kappa_{j}-\mathbf{x}^{\prime}\widehat{\boldsymbol{\eta}}^{(0)}_{11}\right)\right]\right),$
-where $\widehat{\boldsymbol{\eta}}_{11}^{(0)}:=\widehat{\boldsymbol{\eta}}_{01}^{(0)}+\widehat{\boldsymbol{\eta}}_{10}^{(0)}-\widehat{\boldsymbol{\eta}}_{00}^{(0)}$ and $\widehat{F}_{{\widetilde{\varepsilon}}_{11}^{(0)}}(v):=\widehat{F}_{{\widetilde{\varepsilon}}_{10}^{(0)}}(v)$ for all $v\in\mathbb{R}$, while $\Psi_{11}^{(d)}$, $d\in\{0,1\}$, denote suitable trimming sets.
-Under regularity conditions on the different estimators, the above procedure yields consistent estimates of the DTTs of interest, see Section LABEL:sec:proofEstimation of the supplementary material for details.
-Finally, when the median (or mean) as well as the variance parameters actually differ, a slightly modified procedure can be adopted.
-In particular, we can construct estimators for the median (or mean) parameter $\eta_{c,gt}^{(d)}$ and the scaling parameter $\lambda_{gt}^{(d)}$ as follows:
+Remark 1: As the identification argument presented in this subsection is not constructive, estimation may be carried out via Maximum Likelihood (ML) relying on a specific choice for the copula function.
+In Section S.3.2 of the supplementary material, we present corresponding ML estimators for panel as well as repeated cross-sectional data, and discuss their asymptotic properties.
+Specifically, as all model parameters are estimated jointly, note that consistent estimation requires both Assumption 7 (i)-(a) and (i)-(b) to hold simultaneously.
+For a given choice of the marginal distribution $F_{-\varepsilon}$ and $F_{-\nu}$, respectively, the ML estimators of the conditional DTTs outlined in Proposition 4 are given by:
+$\widehat{\tau}^{c}(j\mid\mathbf{x})$ $:=\left[F_{-\varepsilon}\left(\frac{\overline{\mathbf{x}}^{\prime}\widehat{\overline{\boldsymbol{\eta}}}^{(1)}_{11}-\kappa_{j+1}}{\widehat{\lambda}^{(1)}_{11}}\right)-F_{-\varepsilon}\left(\frac{\overline{\mathbf{x}}^{\prime}\widehat{\overline{\boldsymbol{\eta}}}^{(1)}_{11}-\kappa_{j}}{\widehat{\lambda}^{(1)}_{11}}\right)\right]-\left[F_{-\varepsilon}\left(\frac{\overline{\mathbf{x}}^{\prime}\widehat{\overline{\boldsymbol{\eta}}}^{(0)}_{11}-\kappa_{j+1}}{\widehat{\lambda}^{(0)}_{11}}\right)-F_{-\varepsilon}\left(\frac{\overline{\mathbf{x}}^{\prime}\widehat{\overline{\boldsymbol{\eta}}}^{(0)}_{11}-\kappa_{j}}{\widehat{\lambda}^{(0)}_{11}}\right)\right]$
+$\widehat{\tau}^{r}(j\mid\mathbf{z})$ $:=\left[F_{-\nu}\left(\frac{\overline{\mathbf{z}}^{\prime}\widehat{\overline{\boldsymbol{\pi}}}^{(1)}_{11}-\iota_{j+1}}{\widehat{\zeta}^{(1)}_{11}}\right)-F_{-\nu}\left(\frac{\overline{\mathbf{z}}^{\prime}\widehat{\overline{\boldsymbol{\pi}}}^{(1)}_{11}-\iota_{j}}{\widehat{\zeta}^{(1)}_{11}}\right)\right]-\left[F_{-\nu}\left(\frac{\overline{\mathbf{z}}^{\prime}\widehat{\overline{\boldsymbol{\pi}}}^{(0)}_{11}-\iota_{j+1}}{\widehat{\zeta}^{(0)}_{11}}\right)-F_{-\nu}\left(\frac{\overline{\mathbf{z}}^{\prime}\widehat{\overline{\boldsymbol{\pi}}}^{(0)}_{11}-\iota_{j}}{\widehat{\zeta}^{(0)}_{11}}\right)\right]$
+for $j\in\{0,1,2\}$, $\mathbf{x}\in\mathcal{X}_{11}$ and $\mathbf{z}\in\mathcal{Z}_{11}$, where
+$\widehat{\overline{\boldsymbol{\eta}}}_{11}^{(0)}:=\widehat{\overline{\boldsymbol{\eta}}}_{01}^{(0)}+\frac{\widehat{\overline{\boldsymbol{\eta}}}_{10}^{(0)}-\widehat{\overline{\boldsymbol{\eta}}}_{00}^{(0)}}{\widehat{\lambda}_{00}^{(0)}/\widehat{\lambda}_{01}^{(0)}},\quad\widehat{\lambda}_{11}^{(0)}:=\frac{\widehat{\lambda}_{01}^{(0)}\widehat{\lambda}_{10}^{(0)}}{\widehat{\lambda}_{00}^{(0)}},\quad\widehat{\overline{\boldsymbol{\pi}}}_{11}^{(0)}:=\widehat{\overline{\boldsymbol{\pi}}}_{01}^{(0)}+\frac{\widehat{\overline{\boldsymbol{\pi}}}_{10}^{(0)}-\widehat{\overline{\boldsymbol{\pi}}}_{00}^{(0)}}{\widehat{\zeta}_{00}^{(0)}/\widehat{\zeta}_{01}^{(0)}}\quad\text{and}\quad\widehat{\zeta}_{11}^{(0)}:=\frac{\widehat{\zeta}_{01}^{(0)}\widehat{\zeta}_{10}^{(0)}}{\widehat{\zeta}_{00}^{(0)}}$
+are the corresponding ML estimates. Similarly, the construction of “marginalized” DTTs for the treated group after treatment (see, e.g., Chernozhukov et al., 2013) are:
 […]
-[…]
-for all $d=g\cdot t$, $g\in\{0,1\}$ and $t\in\{0,1\}$, where $\Psi^{(d)}_{n,gt}$ now denotes a trimming set on which the index density is strictly bounded away from zero and that slowly expands with the sample size (cf.
-Hansen, 2008).
-This leads to an alternative estimator for the unscaled error term distribution:
-$\widehat{F}_{{\varepsilon}_{gt}^{(d)}}(v):=\widehat{F}_{\widetilde{\varepsilon}_{gt}^{(d)}}\left(\widehat{\eta}_{c,gt}^{(d)}+\widehat{\lambda}_{gt}^{(d)}v\right),$
-for all $\left(\widehat{\eta}_{c,gt}^{(d)}+\widehat{\lambda}_{gt}^{(d)}v\right)\in\Psi^{(d)}_{n,gt}$, though deriving the asymptotic distribution in this case is more involved and beyond the scope of the paper.
-7.2 Estimation With Underreported Outcomes
-In this subsection, we turn to the estimation of the DTTs when outcomes are subject to underreporting, as elaborated in Section 5.
-Acknowledging the non-constructive nature of the identification argument presented therein, our subsequent exposition assumes that $\overline{C}_{gt}^{(d)}(\cdot,\cdot)$ is a parametric copula, specifically $\overline{C}_{gt}^{(d)}\left(\cdot,\cdot;\rho_{gt}^{(d)}\right)$, parameterized by a dependence parameter $\rho_{gt}^{(d)}$ for all $d=g\cdot t$, $g\in\{0,1\}$ and $t\in\{0,1\}$.
-Additionally, we postulate that the distributions of $-\varepsilon_{gt}^{(d)}$ and $-\nu_{gt}^{(d)}$ are known.
-For simplicity, we assume invariance across groups, time and treatment statuses in terms of copula functions and marginal distributions, i.e., $F_{\varepsilon_{gt}^{(d)}}=F_{\varepsilon}$, $F_{\nu_{gt}^{(d)}}=F_{\nu}$, and $\overline{C}_{gt}^{(d)}=\overline{C}$.
-This assumption can be relaxed straightforwardly if needed.
-We define $\widetilde{\boldsymbol{\theta}}^{(d)}_{gt}:=\left(\overline{\boldsymbol{\eta}}^{(d)\prime}_{gt},\lambda^{(d)}_{gt},\overline{\boldsymbol{\pi}}^{(d)\prime}_{gt},\zeta^{(d)}_{gt},\rho^{(d)}_{gt}\right)^{\prime}$ for all $d=g\cdot t$, $g\in\{0,1\}$ and $t\in\{0,1\},$ and aggregate these into $\widetilde{\boldsymbol{\theta}}:=\left(\widetilde{\boldsymbol{\theta}}^{(0)\prime}_{00},\widetilde{\boldsymbol{\theta}}^{(0)\prime}_{01},\widetilde{\boldsymbol{\theta}}^{(0)\prime}_{10},\widetilde{\boldsymbol{\theta}}^{(1)\prime}_{11}\right)^{\prime}$.
-With this, the estimator in (7.1) is modified to:
-$\widehat{\widetilde{\boldsymbol{\theta}}}:=$ $\arg\min_{\widetilde{\boldsymbol{\theta}}\in\mathcal{B}}\sum_{i=1}^{n}\sum_{j=0}^{2}\mathbb{I}\left[C_{i}=j\right]\left\{\sum_{(g,t)\neq(1,1)}\mathbb{I}\left[G_{i}=g,T_{i}=t\right]\times\right.$
-$\left.\ln\left[\overline{C}\left(F_{-\varepsilon}\left(\frac{\overline{\mathbf{X}}_{i}^{\prime}\overline{\boldsymbol{\eta}}_{gt}^{(0)}-\kappa_{j+1}}{\lambda_{gt}^{(0)}}\right),F_{-\nu}\left(\frac{\overline{\mathbf{Z}}_{i}^{\prime}\overline{\pi}_{gt}^{(0)}-\iota_{j+1}}{\zeta_{gt}^{(0)}}\right);\rho_{gt}^{(0)}\right)-\overline{C}\left(F_{-\varepsilon}\left(\frac{\overline{\mathbf{X}}_{i}^{\prime}\overline{\boldsymbol{\eta}}_{gt}^{(0)}-\kappa_{j}}{\lambda_{gt}^{(0)}}\right),F_{-\nu}\left(\frac{\overline{\mathbf{Z}}_{i}^{\prime}\overline{\boldsymbol{\pi}}_{gt}^{(0)}-\iota_{j}}{\zeta_{gt}^{(0)}}\right);\rho_{gt}^{(0)}\right)\right]\right.$
-$\left.+\mathbb{I}\left[G_{i}=1,T_{i}=1\right]\ln\left[\overline{C}\left(F_{-\varepsilon}\left(\frac{\overline{\mathbf{X}}_{i}^{\prime}\overline{\boldsymbol{\eta}}_{11}^{(1)}-\kappa_{j+1}}{\lambda_{11}^{(1)}}\right),F_{-\nu}\left(\frac{\overline{\mathbf{Z}}_{i}^{\prime}\overline{\pi}_{11}^{(1)}-\iota_{j+1}}{\zeta_{11}^{(1)}}\right);\rho_{11}^{(1)}\right)\right.\right.$ (23)
-$\left.\left.-\overline{C}\left(F_{-\varepsilon}\left(\frac{\overline{\mathbf{X}}_{i}^{\prime}\overline{\boldsymbol{\eta}}_{11}^{(1)}-\kappa_{j}}{\lambda_{11}^{(1)}}\right),F_{-\nu}\left(\frac{\overline{\mathbf{Z}}_{i}^{\prime}\overline{\boldsymbol{\pi}}_{11}^{(1)}-\iota_{j}}{\zeta_{11}^{(1)}}\right);\rho_{11}^{(1)}\right)\right]\right\}$
-This estimator is conceptually similar to the first step of the estimation procedure described in Nguimkeu et al.
-(2019), which deals with the estimation of linear models with a partially observed, endogenous regressor within a group and a time period.
-Using the estimates, we construct the estimated conditional DTTs as outlined in Proposition 6:
-$\widehat{\tau}^{c}_{11}(j\mid\mathbf{x})$ $:=\left[F_{-\varepsilon}\left(\frac{\overline{\mathbf{x}}^{\prime}\widehat{\overline{\boldsymbol{\eta}}}^{(1)}_{11}-\kappa_{j+1}}{\widehat{\lambda}^{(1)}_{11}}\right)-F_{-\varepsilon}\left(\frac{\overline{\mathbf{x}}^{\prime}\widehat{\overline{\boldsymbol{\eta}}}^{(1)}_{11}-\kappa_{j}}{\widehat{\lambda}^{(1)}_{11}}\right)\right]-\left[F_{-\varepsilon}\left(\frac{\overline{\mathbf{x}}^{\prime}\widehat{\overline{\boldsymbol{\eta}}}^{(0)}_{11}-\kappa_{j+1}}{\widehat{\lambda}^{(0)}_{11}}\right)-F_{-\varepsilon}\left(\frac{\overline{\mathbf{x}}^{\prime}\widehat{\overline{\boldsymbol{\eta}}}^{(0)}_{11}-\kappa_{j}}{\widehat{\lambda}^{(0)}_{11}}\right)\right]$
-$\widehat{\tau}^{r}_{11}(j\mid\mathbf{z})$ $:=\left[F_{-\nu}\left(\frac{\overline{\mathbf{z}}^{\prime}\widehat{\overline{\boldsymbol{\pi}}}^{(1)}_{11}-\iota_{j+1}}{\widehat{\zeta}^{(1)}_{11}}\right)-F_{-\nu}\left(\frac{\overline{\mathbf{z}}^{\prime}\widehat{\overline{\boldsymbol{\pi}}}^{(1)}_{11}-\iota_{j}}{\widehat{\zeta}^{(1)}_{11}}\right)\right]-\left[F_{-\nu}\left(\frac{\overline{\mathbf{z}}^{\prime}\widehat{\overline{\boldsymbol{\pi}}}^{(0)}_{11}-\iota_{j+1}}{\widehat{\zeta}^{(0)}_{11}}\right)-F_{-\nu}\left(\frac{\overline{\mathbf{z}}^{\prime}\widehat{\overline{\boldsymbol{\pi}}}^{(0)}_{11}-\iota_{j}}{\widehat{\zeta}^{(0)}_{11}}\right)\right]$
-for $j\in\{0,1,2\}$, $\mathbf{x}\in\mathcal{X}_{11}$ and $\mathbf{z}\in\mathcal{Z}_{11}$, where
-$\widehat{\overline{\boldsymbol{\eta}}}_{11}^{(0)}:=\widehat{\overline{\boldsymbol{\eta}}}_{01}^{(0)}+\frac{\widehat{\overline{\boldsymbol{\eta}}}_{10}^{(0)}-\widehat{\overline{\boldsymbol{\eta}}}_{00}^{(0)}}{\widehat{\lambda}_{00}^{(0)}/\widehat{\lambda}_{01}^{(0)}},\quad\widehat{\lambda}_{11}^{(0)}:=\frac{\widehat{\lambda}_{01}^{(0)}\widehat{\lambda}_{10}^{(0)}}{\widehat{\lambda}_{00}^{(0)}},\quad\widehat{\overline{\boldsymbol{\pi}}}_{11}^{(0)}:=\widehat{\overline{\boldsymbol{\pi}}}_{01}^{(0)}+\frac{\widehat{\overline{\boldsymbol{\pi}}}_{10}^{(0)}-\widehat{\overline{\boldsymbol{\pi}}}_{00}^{(0)}}{\widehat{\zeta}_{00}^{(0)}/\widehat{\zeta}_{01}^{(0)}}\quad\text{and}\quad\widehat{\zeta}_{11}^{(0)}:=\frac{\widehat{\zeta}_{01}^{(0)}\widehat{\zeta}_{10}^{(0)}}{\widehat{\zeta}_{00}^{(0)}}.$
-We also explore the construction of “marginalized” DTTs for the treated group after treatment (see, e.g., Chernozhukov et al., 2013):
-[…]
-Here, $\widehat{F}_{\mathbf{X}_{11}}(\cdot)$ represents an estimator of the covariate distribution. For instance:
-$\widehat{\tau}^{c}_{11}(j):=$ $\frac{1}{n_{11}}\sum_{i=1}^{n_{11}}\mathbb{I}\left[G_{i}=1,T_{i}=1\right]\mathbb{I}\left[\mathbf{X}_{i}\in\mathcal{X}_{11}\right]\left(F_{-\varepsilon}\left(\frac{\overline{\mathbf{X}}_{i}^{\prime}\widehat{\overline{\boldsymbol{\eta}}}^{(1)}_{11}-\kappa_{j+1}}{\widehat{\lambda}^{(1)}_{11}}\right)-F_{-\varepsilon}\left(\frac{\overline{\mathbf{X}}_{i}^{\prime}\widehat{\overline{\boldsymbol{\eta}}}^{(1)}_{11}-\kappa_{j}}{\widehat{\lambda}^{(1)}_{11}}\right)\right)$
-$-\frac{1}{n_{11}}\sum_{i=1}^{n_{11}}\mathbb{I}\left[G_{i}=1,T_{i}=1\right]\mathbb{I}\left[\mathbf{X}_{i}\in\mathcal{X}_{11}\right]\left(F_{-\varepsilon}\left(\frac{\overline{\mathbf{X}}_{i}^{\prime}\widehat{\overline{\boldsymbol{\eta}}}^{(0)}_{11}-\kappa_{j+1}}{\widehat{\lambda}^{(0)}_{11}}\right)-F_{-\varepsilon}\left(\frac{\overline{\mathbf{X}}_{i}^{\prime}\widehat{\overline{\boldsymbol{\eta}}}^{(0)}_{11}-\kappa_{j}}{\widehat{\lambda}^{(0)}_{11}}\right)\right).$
-The asymptotic distribution of these DTT variants can be derived similarly to the approach in Subsection 7.1, with standard errors constructed accordingly.
-Remark 3: Test statistics for the hypotheses $H_{0}^{C}$ vs.
-$H_{A}^{C}$ or $H_{0}^{R}$ vs.
-$H_{A}^{R}$, as discussed in Remark 1, may be readily constructed using standard Cramer-von Mises measures of distance.
-Specifically, for testing $H_{0}^{C}$ vs.
-$H_{A}^{C}$, the following statistic can be employed for some $(g.t)\neq(1,1)$:
-$\sum_{i=1}^{n}\sum_{j=1}^{2}\mathbb{I}\left[G_{i}=g,T_{i}=t\right]\mathbb{I}\left[(\mathbf{X}_{i},\mathbf{Z}_{i})\in\mathcal{X}_{gt}\times\mathcal{Z}_{gt}\right]$
-×$\times\left[\overline{C}\left(F_{-{\varepsilon}}\left(\frac{\overline{\mathbf{X}}_{i}^{\prime}\widehat{\overline{\boldsymbol{\eta}}}_{gt}^{(0)}-\kappa_{j}}{\widehat{\lambda}_{gt}^{(0)}}\right),F_{-{\nu}}\left(\frac{\overline{\mathbf{Z}}_{i}^{\prime}\widehat{\overline{\boldsymbol{\pi}}}_{gt}^{(0)}-\iota_{j}}{\widehat{\zeta}_{gt}^{(0)}}\right);\widehat{\rho}_{gt}^{(0)}\right)-\widehat{M}\left(j;\mathbf{X}_{i}^{\prime}\widehat{\widetilde{\boldsymbol{\eta}}}_{gt},\mathbf{Z}_{i}^{\prime}\widehat{\widetilde{\boldsymbol{\pi}}}_{gt}\right)\right]^{2},$
-where $\left(\widehat{\overline{\boldsymbol{\eta}}}_{gt}^{(0)^{\prime}},\widehat{\lambda}_{gt}^{(0)},\widehat{\overline{\boldsymbol{\pi}}}_{gt}^{(0)^{\prime}},\widehat{\zeta}^{(0)}_{gt},\widehat{\rho}_{gt}^{(0)}\right)$ are ML estimators discussed before, $\left(\widehat{\widetilde{\boldsymbol{\eta}}}_{gt}^{\prime},\widehat{\widetilde{\boldsymbol{\pi}}}_{gt}^{\prime}\right)$ denote multi-index estimators (Donkers and Schafgans, 2008; Ahn et al., 2018), and $\widehat{M}(j;\cdot,\cdot)$ is a suitable nonparametric estimator.
-A similar approach can be formulated for testing $H_{0}^{R}$ vs.
-$H_{A}^{R}$.
-The limiting distributions of both tests, while beyond the scope of the paper, are standard, and suitable resampling procedures can be used to construct asymptotically valid critical values.
-8 Simulation
-In this section, we present Monte Carlo simulation evidence to investigate the finite sample behavior of the ML estimator outlined in Subsection 7.2.
-Our focus is on assessing the finite sample bias and variance of the ML estimator, specifically in relation to the model parameters for marginal consumption and reporting distributions, as well as the dependence parameter in the copula function.
-The analysis spans various exclusion restriction scenarios, comparing cases with equal index and error variances to those where the index variance is less than the error variance.
-The design of our observed outcome equation for simulation aligns with that in (1), but streamlined by not incorporating the treatment effects framework and simplifying the notation for the potential outcomes and the $gt$ subscript.
-Specifically, we model the consumption and reporting decisions as:
-$C=\sum_{j=0}^{2}\mathbb{I}\left[\beta_{0}+\beta_{1}X+\beta_{2}W-\lambda\varepsilon>\kappa_{j}\right],$
-$R=\sum_{j=0}^{2}\mathbb{I}\left[\pi_{0}+\pi_{1}Z+\pi_{2}W-\zeta\nu>\iota_{j}\right].$
-Here, $W\sim\mathcal{N}(0,1)$ represents a common regressor, while $X$ and $Z$ are the “excluded” variables following a uniform distribution $\mathcal{U}\left(-\sqrt{3},\sqrt{3}\right)$ (detailed in Table 1).
-The joint distribution $(\varepsilon,\nu)^{\prime}$ is modeled using a Frank copula, denoted as $\overline{C}\left(F_{\varepsilon},F_{\nu};\rho\right)$, with the marginals being standard normal distributions.
-The Frank copula, representative of symmetric copulas, is defined as:
-$\overline{C}(v_{1},v_{2};\rho)=-\frac{1}{\rho}\ln\left(1+\frac{(\exp(-\rho v_{1})-1)(\exp(-\rho v_{2})-1)}{\exp(-\rho)-1}\right)$
-for $\rho\neq 0$. When $\rho=0$, it defaults to the independence copula.
-For the simulation purposes, the dependence parameter $\rho$ is set such that the corresponding Spearman’s $\rho$ is 0.50.5-0.5- 0.5.111111Results for cases where the Spearman’s $\rho$ equals $-0.25$, yielding qualitatively similar outcomes, are available in Subsection LABEL:ssec:frank of the supplementary material.
-This approach is intended to facilitate comparability between dependence parameters across copulas, each of which can have distinct interpretations (see Han and Lee, 2019).
-Additionally, in Section LABEL:sec:simulations of the supplementary material, we extend our analysis to include simulations with the Clayton Copula, known for its asymmetric properties.
-These simulations yield findings that are qualitatively similar to those obtained using the Frank copula, thus highlighting the robustness of our study.
-We consider sample sizes $n=\{2000,5000,10000\}$, with each size undergoing 2000 Monte Carlo replications.
-To assess finite sample bias and variance, we focus on the Mean Bias (MB) and the Root Mean Square Error (RMSE).
-Table 1 provides an overview of the various designs examined in our Monte Carlo study.
-We selected distributions and parameter values to ensure that the (unconditional) variances of the indices and the unobservables are equal to each other, with the exceptions of Cases 1-b and 2-b.
-Table 1: Simulation Design Cases
-Case No. of Exclusions Specification
-0 No Exclusion $C=\sum_{j=1}^{2}\mathbb{I}\left[-1+2\cdot W-2\cdot\varepsilon>\kappa_{j}\right]$
-$R=\sum_{j=1}^{2}\mathbb{I}\left[1.5-2\cdot W-2\cdot\nu>\iota_{j}\right]$
-1-a 1 Exclusion $C=\sum_{j=1}^{2}\mathbb{I}\left[-1+2\cdot W-2\cdot\varepsilon>\kappa_{j}\right]$
-$R=\sum_{j=1}^{2}\mathbb{I}\left[1.5-\sqrt{2}\cdot W-\sqrt{2}\cdot Z-2\cdot\nu>\iota_{j}\right]$
-1-b 1 Exclusion $C=\sum_{j=1}^{2}\mathbb{I}\left[-1+\sqrt{2}\cdot W-2\cdot\varepsilon>\kappa_{j}\right]$
-$R=\sum_{j=1}^{2}\mathbb{I}\left[1.5-W-Z-2\cdot\nu>\iota_{j}\right]$
-2-a 2 Exclusions $C=\sum_{j=1}^{2}\mathbb{I}\left[-1+\sqrt{2}\cdot W+\sqrt{2}\cdot X-2\cdot\varepsilon>\kappa_{j}\right]$
-$R=\sum_{j=1}^{2}\mathbb{I}\left[1.5-\sqrt{2}\cdot W-\sqrt{2}\cdot Z-2\cdot\nu>\iota_{j}\right]$
-2-b 2 Exclusions $C=\sum_{j=1}^{2}\mathbb{I}\left[-1+W+X-2\cdot\varepsilon>\kappa_{j}\right]$
-$R=\sum_{j=1}^{2}\mathbb{I}\left[1.5-W-Z-2\cdot\nu>\iota_{j}\right]$
-Case 0 represents a scenario without exclusion restrictions, where identification is expected to fail.
-In contrast, Cases 1-a and 1-b focus on designs with a single exclusion restriction, differing in their (unconditional) index variances.
-Case 1-a considers equal (unconditional) index and error variances, whereas Case 1-b explores a setting with a smaller index variance relative to that of the error.
-Extending this framework, Cases 2-a and 2-b, akin to Cases 1-a and 1-b, respectively, incorporate two distinct exclusion restrictions in each equation.
-Table 2: Frank Copula - Mean Bias
-Case 0
-Param. obs = 2000, $\rho$ = -0.5 obs = 5000, $\rho$ = -0.5 obs = 10000, $\rho$ = -0.5
-$\beta_{0}$ $45.387$$10.271$$0.289$
-$\beta_{1}$ $605.906$ $92.359$ $16.638$
-$\pi_{0}$ $62.434$$12.121$$1.343$
-$\pi_{1}$ $651.649$ $141.740$ $2.395$
-$\lambda$ $719.680$ $79.700$ $16.190$
-$\zeta$ $715.016$ $138.753$ $0.252$
-$\rho$ $0.393$ $0.300$ $0.218$
-Case 1-a
-Param. obs = 2000, $\rho$ = -0.5 obs = 5000, $\rho$ = -0.5 obs = 10000, $\rho$ = -0.5
-$\beta_{0}$$9.946$ $0.011$ $0.007$
-$\beta_{1}$ $124.558$ $0.110$ $0.046$
-$\pi_{0}$ $100.666$$0.229$$0.092$
-$\pi_{1}$$77.552$ $0.110$ $0.040$
-$\pi_{2}$$74.046$ $0.020$$0.002$
-$\lambda$ $115.175$ $0.076$ $0.031$
-$\zeta$ $91.241$$0.021$ $0.006$
-$\rho$ $0.253$ $0.144$ $0.066$
-Case 1-b
-Param. obs = 2000, $\rho$ = -0.5 obs = 5000, $\rho$ = -0.5 obs = 10000, $\rho$ = -0.5
-$\beta_{0}$$245.968$ $4.973$$0.011$
-$\beta_{1}$ $1,902.489$ $15.519$ $0.128$
-$\pi_{0}$ $252.689$$0.445$$0.341$
-$\pi_{1}$$154.198$ $0.178$ $0.138$
-$\pi_{2}$$149.381$ $0.034$ $0.040$
-$\lambda$ $2,350.347$ $20.812$ $0.160$
-$\zeta$ $270.014$$0.080$$0.078$
-$\rho$ $0.364$ $0.273$ $0.190$
-Case 2-a
-Param. obs = 2000, $\rho$ = -0.5 obs = 5000, $\rho$ = -0.5 obs = 10000, $\rho$ = -0.5
-$\beta_{0}$$0.035$$0.019$$0.005$
-$\beta_{1}$ $0.114$ $0.046$ $0.016$
-$\beta_{2}$ $0.118$ $0.046$ $0.017$
-$\pi_{0}$ $13.961$ $0.043$ $0.028$
-$\pi_{1}$$9.951$$0.040$$0.023$
-$\pi_{2}$$11.006$$0.038$$0.024$
-$\lambda$ $0.133$ $0.052$ $0.020$
-$\zeta$ $13.324$ $0.043$ $0.029$
-$\rho$ $0.026$ $0.007$ $0.001$
-Case 2-b
-Param. obs = 2000, $\rho$ = -0.5 obs = 5000, $\rho$ = -0.5 obs = 10000, $\rho$ = -0.5
-$\beta_{0}$$280.768$$3.627$$0.014$
-$\beta_{1}$ $603.093$ $9.119$ $0.055$
-$\beta_{2}$ $603.862$ $8.056$ $0.050$
-$\pi_{0}$ $238.591$ $0.152$ $0.063$
-$\pi_{1}$$139.640$$0.088$$0.038$
-$\pi_{2}$$147.751$$0.096$$0.034$
-$\lambda$ $1,122.068$ $16.560$ $0.091$
-$\zeta$ $285.927$ $0.167$ $0.071$
-$\rho$ $0.063$ $0.017$ $0.004$
-Notes: See Table 1 for the specific design of each case. All results are produced using $2000$ Monte Carlo replications.
-Table 3: Frank Copula - RMSE
-Case 0
-Param. obs = 2000, $\rho$ = -0.5 obs = 5000, $\rho$ = -0.5 obs = 10000, $\rho$ = -0.5
-$\beta_{0}$ $2,860.340$ $310.396$ $61.885$
-$\beta_{1}$ $8,978.860$ $2,086.819$ $816.737$
-$\pi_{0}$ $2,553.201$ $688.355$ $2.279$
-$\pi_{1}$ $7,481.836$ $2,942.766$ $4.219$
-$\lambda$ $7,083.483$ $1,754.262$ $714.109$
-$\zeta$ $5,732.446$ $2,576.872$ $2.182$
-$\rho$ $0.650$ $0.575$ $0.489$
-Case 1-a
-Param. obs = 2000, $\rho$ = -0.5 obs = 5000, $\rho$ = -0.5 obs = 10000, $\rho$ = -0.5
-$\beta_{0}$ $684.210$ $0.190$ $0.126$
-$\beta_{1}$ $3,169.837$ $0.471$ $0.285$
-$\pi_{0}$ $2,328.631$ $1.059$ $0.736$
-$\pi_{1}$ $1,702.444$ $0.685$ $0.474$
-$\pi_{2}$ $1,661.837$ $0.433$ $0.293$
-$\lambda$ $2,992.510$ $0.390$ $0.239$
-$\zeta$ $1,925.182$ $0.607$ $0.428$
-$\rho$ $0.488$ $0.362$ $0.233$
-Case 1-b
-Param. obs = 2000, $\rho$ = -0.5 obs = 5000, $\rho$ = -0.5 obs = 10000, $\rho$ = -0.5
-$\beta_{0}$ $5,103.763$ $211.961$ $0.187$
-$\beta_{1}$ $11,215.290$ $608.039$ $0.436$
-$\pi_{0}$ $3,290.975$ $1.583$ $1.146$
-$\pi_{1}$ $2,035.018$ $0.756$ $0.561$
-$\pi_{2}$ $1,936.292$ $0.576$ $0.341$
-$\lambda$ $13,803.720$ $833.790$ $0.522$
-$\zeta$ $3,457.443$ $1.040$ $0.719$
-rho $0.595$ $0.505$ $0.400$
-Case 2-a
-Param. obs = 2000, $\rho$ = -0.5 obs = 5000, $\rho$ = -0.5 obs = 10000, $\rho$ = -0.5
-$\beta_{0}$ $0.323$ $0.165$ $0.110$
-$\beta_{1}$ $0.621$ $0.265$ $0.175$
-$\beta_{2}$ $0.549$ $0.233$ $0.150$
-$\pi_{0}$ $616.635$ $0.476$ $0.335$
-$\pi_{1}$ $438.369$ $0.359$ $0.249$
-$\pi_{2}$ $484.978$ $0.308$ $0.215$
-$\lambda$ $0.712$ $0.291$ $0.191$
-$\zeta$ $587.512$ $0.415$ $0.288$
-$\rho$ $0.171$ $0.103$ $0.072$
-Case 2-b
-Param. obs = 2000, $\rho$ = -0.5 obs = 5000, $\rho$ = -0.5 obs = 10000, $\rho$ = -0.5
-$\beta_{0}$ $2,581.494$ $107.814$ $0.164$
-$\beta_{1}$ $4,770.145$ $256.687$ $0.261$
-$\beta_{2}$ $4,687.187$ $222.714$ $0.223$
-$\pi_{0}$ $3,211.289$ $0.958$ $0.587$
-$\pi_{1}$ $1,913.604$ $0.512$ $0.305$
-$\pi_{2}$ $2,048.146$ $0.454$ $0.261$
-$\lambda$ $8,605.971$ $459.471$ $0.410$
-$\zeta$ $3,794.507$ $0.889$ $0.509$
-$\rho$ $0.255$ $0.146$ $0.104$
-Notes: See Table 1 for the specific design of each case. All results are produced using $2000$ Monte Carlo replications.
-Analyzing the results presented in Table 2, we observe a pronounced MB across all scenarios for sample sizes of $n=2000$ and $n=5000$, which substantially diminishes at $n=10000$.
-This behavior aligns with the findings for the RMSE in Table 3.
-In Case 0, the absence of exclusion restrictions results in a persistent identification issue across all sample sizes, as evidenced by the excessive MB and RMSE.
-Notably, scenarios with a single (continuous) excluded variable in only one of the equations, specifically Cases 1-a and 1-b, also demonstrate substantial MB at $n=10000$, especially for intercepts, scaling parameters, and dependence parameters.
-Conversely, in Cases 2-a and 2-b, which feature two exclusion restrictions (with continuous excluded variables), MB and RMSE are substantially reduced compared to their counterparts in Cases 1-a and 1-b.
-Interestingly, the (unconditional) variance of the index emerges as a crucial factor, as demonstrated by the larger MB and RMSE in Cases 1-b and 2-b, where the index variance is substantially smaller than that of the error.
-[^12] The RMSE results for Cases 1-b and 2-b indicate that various parameters are very imprecisely estimated at smaller sample sizes, often attributed to outlier estimates.
-A detailed investigation reveals that this imprecision often originates from outlier estimates, further emphasizing the importance of a large index variance for precise estimation.
-In summary, our findings from this section suggest that employing at least two exclusion restrictions (one in each equation) effectively reduces the finite sample MB and RMSE when estimating model parameters of the joint distribution in a partially observable framework.
-Furthermore, a larger index variance contributes to precise estimation.
+Here, $\widehat{F}_{\mathbf{X}_{11}}(\cdot)$ represents an estimator of the covariate distribution.
+Remark 2: While Proposition 2 does not allow any conclusions about the relationship of $\overline{\boldsymbol{\eta}}_{gt}$ and $\lambda_{gt}$ across pre-treatment periods, note that the ML estimator mentioned in previous remark can be used to conduct “pre-treatment” tests of the CiC model structure in analogy to the popular “pre-trend” test.
+That is, using the consumption equation as illustration, we may test the implied CiC relationship in Equation (11) from Proposition 2 in pre-treatment periods $t=0,-1,-2,\ldots$.
+Specifically, under the assumption that Equation (11) holds also for any given pre-treatment period $t\in\{0,-1,-2,\ldots\}$, we may test:
+$H_{0}:\overline{\boldsymbol{\eta}}^{(0)}_{1t}-\overline{\boldsymbol{\eta}}^{(0)}_{0t}-\frac{\lambda^{(0)}_{0t}}{\lambda^{(0)}_{0(t-1)}}\left(\overline{\boldsymbol{\eta}}^{(0)}_{1(t-1)}-\overline{\boldsymbol{\eta}}^{(0)}_{0(t-1)}\right)=\mathbf{0}\quad\text{and}\quad\frac{\lambda^{(0)}_{1t}}{\lambda^{(0)}_{1(t-1)}}-\frac{\lambda^{(0)}_{0t}}{\lambda^{(0)}_{0(t-1)}}=0,$
+against
+$H_{A}:\overline{\boldsymbol{\eta}}^{(0)}_{1t}-\overline{\boldsymbol{\eta}}^{(0)}_{0t}-\frac{\lambda^{(0)}_{0t}}{\lambda^{(0)}_{0(t-1)}}\left(\overline{\boldsymbol{\eta}}^{(0)}_{1(t-1)}-\overline{\boldsymbol{\eta}}^{(0)}_{0(t-1)}\right)=\mathbf{0}\quad\text{and/ or}\quad\frac{\lambda^{(0)}_{1t}}{\lambda^{(0)}_{1(t-1)}}-\frac{\lambda^{(0)}_{0t}}{\lambda^{(0)}_{0(t-1)}}=0.$
+via a Likelihood Ratio test comparing an unrestricted log-likelihood as in Equation (15) with a restricted version that satisfies $\overline{\boldsymbol{\eta}}^{(0)}_{1t}=\overline{\boldsymbol{\eta}}^{(0)}_{0t}+\frac{\lambda^{(0)}_{0t}}{\lambda^{(0)}_{0(t-1)}}(\overline{\boldsymbol{\eta}}^{(0)}_{1(t-1)}-\overline{\boldsymbol{\eta}}^{(0)}_{0(t-1)})$ and $\lambda^{(0)}_{1t}=\frac{\lambda^{(0)}_{0t}\lambda^{(0)}_{1(t-1)}}{\lambda^{(0)}_{0(t-1)}}$, respectively.
 5 Empirical Application
@@
 The potential of cannabis serving as a gateway to more harmful drug use has also been documented (see Pudney, 2010, and references therein).
-On the contrary, various other countries have initiated a legalization process of the possession and consumption of certain categories of cannabis drugs for recreational and medicinal purposes, driven by aims to reduce drug-related crime and corresponding law enforcement costs, as well as to generate additional tax revenue.
-Notably, nearly one-quarter of the U.S.
-states have legalized recreational marijuana by now, regulating and taxing its sales (Brinkman and Mok-Lamme, 2019; Hansen et al., 2020).
 In this section, we investigate the effects of recreational marijuana legalization for adults in several U.S.
@@
 The outcome variable of interest (cons_30days), based on the survey responses to the question “On how many occasions have you used marijuana (weed, pot) or hashish (hash, hash oil) … in the last 30 days?”, is categorized into three levels: “0” (never used), “1” (used 1-2 times), and “2” (used more than 2 times).
+[^11] Table LABEL:Summary1 in Section LABEL:sec:additionalempirical of the supplement presents descriptive statistics of the outcome variable and other controls.
 Similar to Hollingworth et al.
@@
 On the other hand, “survey cooperation”, defined as the proportion of missing survey responses (based solely on questions common to all questionnaire forms and not directly related to drug use), is assumed to provide an indication of the respondents’ willingness to cooperate in the survey, unrelated to marijuana consumption itself, as the reporting decision occurs post-consumption only.
-Table 4 presents descriptive statistics of the outcome variable and other controls.
-As expected, marijuana consumption among 8th graders appears relatively low, with only 6% of the sample reporting usage over the past 30 days.
-Further, unreported results indicate that these characteristics remain largely stable across time within each group.
-Table 4: Summary Statistics
-Variable N Mean Std. Dev. Min Pctl. 25 Pctl. 75 Max
-cons_30days 46472
-… 1-2 occasions 1290 3%
-… 3 or more occasions 1375 3%
-… no use 43807 94%
-male 46472 0.48 0.5 0 0 1 1
-age_month 46472 169 6.2 122 165 173 231
-white 46472 0.47 0.5 0 0 1 1
-parent_edu_clg 46472 0.58 0.49 0 0 1 1
-state_unemp 46472 5.3 0.78 2.9 4.8 5.9 6.8
-state_median_income 46472 56552 7024 40037 51983 60925 75675
-state_price_m 46472 240 35 159 220 265 325
-trust 46472 0.96 0.072 0.28 0.96 1 1
-treated 46472 0.19 0.39 0 0 0 1
-after 46472 0.47 0.5 0 0 1 1
-Notes: Data for 8th graders from 2015-2018, excluding the states Oregon, District of Columbia, Colorado and Washington.
-treated = 1 if the respondent is from California, Maine, Massachusetts and Nevada, = 0 otherwise.
-after = 1 if the respondent is surveyed in 2017-2018, = 0 otherwise.
-The variables male, white, parent_edu_clg are binary indicators for self-reported gender status (male), ethnicity (white), and parental education level (college) equal to one if the category was true, and zero otherwise.
-For the other variables, please refer to the text for the definition.
-In Table 5, we present our initial findings from the unconditional analysis, as defined in (21), without misreporting and covariates.
+In Table 1, we present our initial findings for a model without underreporting and without additional control variables.
 Our first set of results (gdid) suggests that $\tau^{c}_{11}(j)$ is not statistically different from zero for $j\in\{0,1,2\}$ at any conventional levels.
@@
 The first panel of results (“Without Misreporting”) displays estimates of treatment effects $\tau_{11}(j)$, $j\in\{0,1,2\}$, for the model estimated without misreporting.
-The treatment effect estimates have been marginalized with respect to the distribution of $\mathbf{X}_{11}$ (see Equation (21)).
+The treatment effect estimates have been marginalized with respect to the distribution of $\mathbf{X}_{11}$.
 The second panel of results (“With Misreporting”) displays estimates of treatment effects $\tau_{11}(j)$ and $\tau_{11}^{r}(j)$, $j\in\{0,1,2\}$, for the model with misreporting.
@@
 6 Conclusion
-This paper develops a generalized Difference-in-Difference model for discrete, possibly underreported outcomes in repeated cross-sectional or panel data.
+This paper develops a Difference-in-Differences model for discrete, possibly underreported ordered outcomes in repeated cross-sectional or panel data.
 The models are based on the continuous CiC model of Athey and Imbens (2006).
-For the case without misreporting and covariates, we propose a parametric model that generalizes standard Probit or Logit DiD models which only accommodate mean shifts.
-When covariates are present (but outcomes are reported correctly), we relax the distributional assumption on the unobservables, and provide a point identification result for a semiparametric index model.
-The result is complementary to Athey and Imbens (2006) and offers a direct pathway to estimation of the structural model and corresponding treatment effects.
-In the second step, we extend the CiC model for discrete outcomes to partially observable outcomes.
+We extend this CiC model for discrete outcomes to partially observable outcomes.
 Special features of this extended model are that it allows the underreporting behavior to be correlated with the outcome of interest and that underreporting may depend on the treatment status itself.
 We provide point identification conditions for a semiparametric version of the model based on a real analytic extrapolation assumption, and suggest a flexible estimator for a parametric version of the model.
-A set of Monte Carlo simulations investigates the finite sample performance of the corresponding ML estimator under different reporting and exclusion restriction scenarios.
-The results suggest that a single (or no) exclusion restriction may yield poor finite sample performance in the partially observable outcome framework.
 We apply our methodology to investigate the distributional effects of the legalization of recreational marijuana usage for adults in the U.S.
@@
 Importantly, correcting for misreporting increases effect sizes by more than 50%.
-Appendix
-Appendix A Alternative Characterization: Distributional Parallel Trends
-While Proposition 1 outlines the relationships among the functional parameters across different groups and over time, the subsequent result provides a further characterization of how models that satisfy distributional parallel trends are generated:
-Proposition A.1.
-Suppose Assumption 1 holds.
-Then, distributional parallel trends from Definition 9 holds if and only if there exist real-valued functions $\alpha_{t}(\mathbf{x}),\mu_{g}(\mathbf{x}),\beta_{t}(\mathbf{x})$ and $\sigma_{g}(\mathbf{x})$ such that $\beta_{t}(\mathbf{x})\sigma_{g}(\mathbf{x})>0$, $g\in\{0,1\}$ and $t\in\{0,1\}$, such that
-$\eta_{gt}(\mathbf{x})=\alpha_{t}(\mathbf{x})+\beta_{t}(\mathbf{x})\mu_{g}(\mathbf{x})\quad\text{and}\quad\lambda_{gt}(\mathbf{x})=\beta_{t}(\mathbf{x})\sigma_{g}(\mathbf{x})$ (25)
-for all covariate values $\mathbf{x}\in\cap_{g,t}\mathcal{X}_{gt}$,
-Proposition A.1 demonstrates that the distributional parallel trends hold if and only if the functional median (mean) and the scaling parameters admit a specific form of separability in group and time.
-Specifically, each functional median (mean) parameter $\eta_{gt}(\mathbf{x})$ can be expressed as a bilinear form comprising a time-trend component $(\alpha_{t}(\mathbf{x}),\beta_{t}(\mathbf{x}))^{\prime}$ and a group-specific component $(1,\mu_{g}(\mathbf{x}))^{\prime}$, and each scaling parameter $\lambda_{gt}(\mathbf{x})$ follows a similar structure involving $\beta_{t}(\mathbf{x})$ and $\sigma_{g}(\mathbf{x})$.
-We highlight that this kind of separability imposes no substantial limitations on how the time trend, the group membership, or the observed heterogeneity affect the outcome variable.
-Furthermore, given the representations in (25), it becomes evident that our model is nonnested within the discrete CiC models with covariates considered by Athey and Imbens (2006), where, when fixing the covariates at specific values, the group membership only allows to affect the outcome variable through the distribution of the unobservable (that is assumed to be within-group time invariant and independent of covariates).
-In contrast, the group membership can still alter the way the covariates affect the outcome variable in our case.
-Provided identification of the functional median (mean) and scaling parameters, Corollary A.2 identifies the underlying structure in the representations (25) up to normalization:
-Corollary A.2.
-Suppose that the functional parameters $\eta_{gt}(\mathbf{x})$ and $\lambda_{gt}(\mathbf{x})$ are identified for all covariate values $\mathbf{x}\in\cap_{g,t}\mathcal{X}_{gt}$, $g\in\{0,1\}$ and $t\in\{0,1\}$.
-If we adopt $\beta_{0}(\mathbf{x})=1$ as an normalization assumption, then $\beta_{1}(\mathbf{x})$ and $\sigma_{g}(\mathbf{x})$ are identified as
-$\beta_{1}(\mathbf{x})=\frac{\lambda_{01}(\mathbf{x})}{\lambda_{00}(\mathbf{x})}\quad\text{and}\quad\sigma_{g}(\mathbf{x})=\lambda_{g0}(\mathbf{x}).$
-for $g\in\{0,1\}$. If we further adopt $\alpha_{0}(\mathbf{x})=0$ as an normalization assumption, then $\alpha_{1}(\mathbf{x})$ and $\mu_{g}(\mathbf{x})$ are identified as
-$\alpha_{1}(\mathbf{x})=\eta_{01}(\mathbf{x})-\eta_{00}(\mathbf{x})\frac{\lambda_{01}(\mathbf{x})}{\lambda_{00}(\mathbf{x})}\quad\text{and}\quad\mu_{g}(\mathbf{x})=\eta_{g0}(\mathbf{x})$
-for $g\in\{0,1\}$.
-It should be noted that the choice of normalizations is arbitrary.
-It merely alters the interpretation of the functional parameters but places no additional restrictions.
-Here, we have chosen to normalize the common time trend in the pre-treatment period to facilitate the interpretation of heterogeneity among the groups.
-Alternatively, a different choice of normalizations could be setting $\mu_{0}(\mathbf{x})=0$ and $\sigma_{0}(\mathbf{x})=1$.
-In light of the structure in (25), we construct a model for the latent outcome $Y^{\ast}_{gt}(0)$, paralleling the features in the assumptions for the continuous CiC model of Athey and Imbens (2006) without covariates, as an example that satisfies the distributional parallel trends in our model with covariates.
-We define
-$Y^{*}_{gt}(0)=h_{t}(\mathbf{X}_{gt},\varepsilon_{gt}),$ (26)
-where $h_{t}(\cdot,\cdot)$ denotes a real-valued function that possibly changes over time. We make the following assumptions:
-Assumption A.1.
-For each $t\in\{0,1\},\mathbf{x}\in\mathbb{R}^{k}$ and $u\in\mathbb{R}$:
-$h_{t}(\mathbf{x},u)=\alpha_{t}(\mathbf{x})+\beta_{t}(\mathbf{x})u,$
-where $\alpha_{t}:\mathbb{R}^{d_{X_{gt}}}\rightarrow\mathbb{R}$ and $\beta_{t}:\mathbb{R}^{d_{X_{gt}}}\rightarrow\mathbb{R}_{+}$ are real-valued functions.
-Assumption A.2.
-$\varepsilon_{gt}$ is strictly stationary within each group given $\mathbf{X}_{gt}$.
-Given Assumptions A.1-A.2 and denoting the conditional expectation and variance of $\varepsilon_{gt}$ given $\mathbf{X}_{gt}$ as $\mu_{g}(\mathbf{X}_{gt})$ and $\sigma^{2}_{g}(\mathbf{X}_{gt})$, respectively, we obtain
-$\mathrm{E}\left[Y_{gt}^{*}(0)\mid\mathbf{X}_{gt}=\mathbf{x}\right]$ $=\alpha_{t}(\mathbf{x})+\beta_{t}(\mathbf{x})\mu_{g}(\mathbf{x})=:\eta_{gt}(\mathbf{x}),$
-$\mathrm{Var}\left[Y_{gt}^{*}(0)\mid\mathbf{X}_{gt}=\mathbf{x}\right]$ $=\left(\beta_{t}(\mathbf{x})\sigma_{g}(\mathbf{x})\right)^{2}=:\lambda^{2}_{gt}(\mathbf{x}),$
-for all $\mathbf{x}\in\mathcal{X}_{gt}$, which thus satisfies the distributional parallel trends in our model with covariates.
-Appendix B Proofs
-Proof of Proposition 1.
-For the “if” part, under Assumption 1
-$F_{Y^{*}_{11}(0)\mathbf{X}_{11}}(v\mid\mathbf{x})$ $=F_{\varepsilon_{11}\mid\mathbf{X}_{11}}\left(\frac{v-\eta_{11}(\mathbf{x})}{\lambda_{11}(\mathbf{x})}\mid\mathbf{x}\right)$
-$=F_{\varepsilon_{11}\mid\mathbf{X}_{11}}\left(\frac{v-\eta_{01}(\mathbf{x})-\frac{\eta_{10}(\mathbf{x})-\eta_{00}(\mathbf{x})}{\lambda_{00}(\mathbf{x})/\lambda_{01}(\mathbf{x})}}{\frac{\lambda_{01}(\mathbf{x})\lambda_{10}(\mathbf{x})}{\lambda_{00}(\mathbf{x})}}\mid\mathbf{x}\right)$
-$=F_{\varepsilon_{10}\mid\mathbf{X}_{10}}\left(\frac{\eta_{00}(\mathbf{x})+\lambda_{00}(\mathbf{x})F_{\varepsilon_{00}\mid\mathbf{X}_{00}}^{-1}\left(F_{\varepsilon_{01}\mid\mathbf{X}_{01}}\left(\frac{v-\eta_{01}(\mathbf{x})}{\lambda_{01}(\mathbf{x})}\mid\mathbf{x}\right)\mid\mathbf{x}\right)-\eta_{10}(\mathbf{x})}{\lambda_{10}(\mathbf{x})}\mid\mathbf{x}\right)$
-$=F_{Y^{*}_{10}(0)\mid\mathbf{X}_{10}}\left(F_{Y^{*}_{00}(0)\mid\mathbf{X}_{00}}^{-1}\left(F_{Y^{*}_{01}(0)\mid\mathbf{X}_{01}}(v\mid\mathbf{x})\mid\mathbf{x}\right)\mid\mathbf{x}\right),$
-for all $v\in\mathbb{R}$ and for all covariate values $\mathbf{x}\in\cap_{g,t}\mathcal{X}_{gt}$, where we used strict stationarity $\varepsilon_{g1}\mid X_{g1}\sim\varepsilon_{g0}\mid X_{g0}$ in the third equality, and $F_{Y^{*}_{gt}(0)\mid\mathbf{X}_{gt}}^{-1}(q\mid\mathbf{x})=\eta_{gt}(\mathbf{x})+\lambda_{gt}(\mathbf{x})F_{\varepsilon_{gt}}^{-1}(q\mid\mathbf{x})$ in the last equality.
-For the “only if” part, let $\mathbf{x}\in\cap_{g,t}\mathcal{X}_{gt}$. Given Assumption 1 and the distributional parallel trends, there exists an open interval $I$ such that
-$F_{Y^{*}_{11}(0)\mid\mathbf{X}_{11}}(v\mid\mathbf{x})$ $=F_{\varepsilon_{11}\mid\mathbf{X}_{11}}\left(\frac{v-\eta_{11}(\mathbf{x})}{\lambda_{11}(\mathbf{x})}\mid\mathbf{x}\right),$
-which is equal to
-$\ F_{Y^{*}_{10}(0)\mid\mathbf{X}_{10}}\left(F_{Y^{*}_{00}(0)\mid\mathbf{X}_{00}}^{-1}\left(F_{Y^{*}_{01}(0)\mid\mathbf{X}_{01}}(v\mid\mathbf{x})\mid\mathbf{x}\right)\mid\mathbf{x}\right)$
-$=F_{\varepsilon_{10}\mid\mathbf{X}_{10}}\left(\frac{\eta_{00}(\mathbf{x})+\lambda_{00}(\mathbf{x})F_{\varepsilon_{00}\mid\mathbf{X}_{00}}^{-1}\left(F_{\varepsilon_{01}\mid\mathbf{X}_{01}}\left(\frac{v-\eta_{01}(\mathbf{x})}{\lambda_{01}(\mathbf{x})}\mid\mathbf{x}\right)\mid\mathbf{x}\right)-\eta_{10}(\mathbf{x})}{\lambda_{10}(\mathbf{x})}\mid\mathbf{x}\right)$
-for all $v\in I$. The strict stationarity $\varepsilon_{g1}\mid\mathbf{X}_{g1}\sim\varepsilon_{g0}\mid\mathbf{X}_{g0}$ further implies a polynomial in $v$ on the open interval $I$:
-$-\frac{\eta_{11}(\mathbf{x})}{\lambda_{11}(\mathbf{x})}-\frac{\eta_{00}(\mathbf{x})-\lambda_{00}(\mathbf{x})\frac{\eta_{01}(\mathbf{x})}{\lambda_{01}(\mathbf{x})}-\eta_{10}(\mathbf{x})}{\lambda_{10}(\mathbf{x})}+\left(\frac{1}{\lambda_{11}(\mathbf{x})}-\frac{\lambda_{00}(\mathbf{x})}{\lambda_{01}(\mathbf{x})\lambda_{10}(\mathbf{x})}\right)v=0.$
-Note that the polynomial is well defined as $\lambda_{gt}(\mathbf{x})>0$ for all $(g,t)\in\{0,1\}$.
-However, a polynomial on an open interval gives rise to the zero function if and only if it is the zero polynomial, that is,
-$-\frac{\eta_{11}(\mathbf{x})}{\lambda_{11}(\mathbf{x})}-\frac{\eta_{00}(\mathbf{x})-\lambda_{00}(\mathbf{x})\frac{\eta_{01}(\mathbf{x})}{\lambda_{01}(\mathbf{x})}-\eta_{10}(\mathbf{x})}{\lambda_{10}(\mathbf{x})}=0\quad\text{and}\quad\frac{1}{\lambda_{11}(\mathbf{x})}-\frac{\lambda_{00}(\mathbf{x})}{\lambda_{01}(\mathbf{x})\lambda_{10}(\mathbf{x})}=0,$
-which implies
-$\eta_{11}(\mathbf{x})=\eta_{01}(\mathbf{x})+\frac{\eta_{10}(\mathbf{x})-\eta_{00}(\mathbf{x})}{\lambda_{00}(\mathbf{x})/\lambda_{01}(\mathbf{x})}\quad\text{and}\quad\lambda_{11}(\mathbf{x})=\frac{\lambda_{01}(\mathbf{x})\lambda_{10}(\mathbf{x})}{\lambda_{00}(\mathbf{x})}.$
-Proof of Proposition 2.
-Let $(g,t)\neq(1,1)$. Since $\kappa_{1}$ and $\kappa_{2}$ are normalized to 0 and 1, we can write the observed fitted probabilities:
-$\mathrm{Pr}\left(Y_{gt}=0\right)=F_{\varepsilon_{gt}}\left(-\frac{\eta_{gt}}{\lambda_{gt}}\right)\quad\text{and}\quad\mathrm{Pr}\left(Y_{gt}\leq 1\right)=F_{\varepsilon_{gt}}\left(\frac{1-\eta_{gt}}{\lambda_{gt}}\right).$
-Given $F_{\varepsilon_{gt}}$ is invertible and the probabilities are nonzero by Assumptions 1, we build the system of equations:
-$a+F_{\varepsilon_{gt}}^{-1}\left(\mathrm{Pr}\left(Y_{gt}=0\right)\right)b=0$
-$a+F_{\varepsilon_{gt}}^{-1}\left(\mathrm{Pr}\left(Y_{gt}\leq 1\right)\right)b=1.$
-If $F_{\varepsilon_{gt}}^{-1}\left(\mathrm{Pr}\left(Y_{gt}=0\right)\right)$ and $F_{\varepsilon_{gt}}^{-1}\left(\mathrm{Pr}\left(Y_{gt}\leq 1\right)\right)$ are known, then $(\eta_{gt},\lambda_{gt})$ is uniquely recovered by
-$\left(\begin{array}[]{ll}\eta_{gt}\\ \lambda_{gt}\end{array}\right)$ $=\left(\begin{array}[]{ll}1&F_{\varepsilon_{gt}}^{-1}\left(\mathrm{Pr}\left(Y_{gt}=0\right)\right)\\ 1&F_{\varepsilon_{gt}}^{-1}\left(\mathrm{Pr}\left(Y_{gt}\leq 1\right)\right)\end{array}\right)^{-1}\left(\begin{array}[]{ll}0\\ 1\end{array}\right)$
-$=\left(\begin{array}[]{ll}\frac{-F_{\varepsilon_{gt}}^{-1}\left(\mathrm{Pr}\left(Y_{gt}=0\right)\right)}{F_{\varepsilon_{gt}}^{-1}\left(\mathrm{Pr}\left(Y_{gt}\leq 1\right)\right)-F_{\varepsilon_{gt}}^{-1}\left(\mathrm{Pr}\left(Y_{gt}=0\right)\right)}\\ \frac{1}{F_{\varepsilon_{gt}}^{-1}\left(\mathrm{Pr}\left(Y_{gt}\leq 1\right)\right)-F_{\varepsilon_{gt}}^{-1}\left(\mathrm{Pr}\left(Y_{gt}=0\right)\right)}\end{array}\right),$
-where the expressions on the RHS are well defined by Assumption 1. Consequently, given the distributional parallel trends, $\eta_{11}$ and $\lambda_{11}$ are identified by Proposition 1 as
-$\eta_{11}=\eta_{01}+\frac{\eta_{10}-\eta_{00}}{\lambda_{00}/\lambda_{01}}\quad\text{ and }\quad\lambda_{11}=\frac{\lambda_{01}\lambda_{10}}{\lambda_{00}}.$
-However, if $F_{\varepsilon_{gt}}^{-1}\left(\mathrm{Pr}\left(Y_{gt}=0\right)\right)$ and $F_{\varepsilon_{gt}}^{-1}\left(\mathrm{Pr}\left(Y_{gt}\leq 1\right)\right)$ are unknowns for $(g,t)\neq(1,1)$, then the system becomes a nonlinear system with four unknowns.
-In such a case, there are infinitely many solutions and so infinitely many observationally equivalent $(\eta_{gt},\lambda_{gt})$ contradicting point identification.
-∎
-Proof of Proposition 3.
-Denoting $\overline{\boldsymbol{\eta}}_{gt}^{*}:=\frac{\overline{\boldsymbol{\eta}}_{gt}}{\left\|\overline{\boldsymbol{\eta}}_{gt}\right\|}$, $\kappa_{j,gt}^{*}=\frac{\kappa_{j}}{\left\|\overline{\boldsymbol{\eta}}_{gt}\right\|}$, $\widetilde{\varepsilon}_{gt}:=\lambda_{gt}\varepsilon_{gt}$ and $\widetilde{\varepsilon}_{gt}^{*}:=\frac{\widetilde{\varepsilon}_{gt}}{\left\|\overline{\boldsymbol{\eta}}_{gt}\right\|}$, the model can be reformulated as:
-$Y_{gt}(0)=$ $\sum_{j=0}^{2}j\cdot\mathbb{I}\left[\kappa_{j,gt}^{*}\leq\overline{\mathbf{X}}_{gt}^{\prime}\overline{\boldsymbol{\eta}}^{*}_{gt}+\widetilde{\varepsilon}^{*}_{gt}\leq\kappa_{j+1,gt}^{*}\right].$
-Let $(g,t)\neq(1,1)$.
-Since Assumptions 2-4 imply conditions SI, X3, and X5 in Manski (1988), we have by Corollary 5 therein that the parameter vector $\overline{\boldsymbol{\eta}}_{gt}^{*}$ is identified.
-As a consequence, the distribution function $F_{\widetilde{\varepsilon}_{gt}^{*}}$ of the rescaled error term $\widetilde{\varepsilon}_{gt}^{*}$, defined by $F_{\widetilde{\varepsilon}_{gt}^{*}}(v)=F_{\varepsilon_{gt}}\left(\frac{\left\|\overline{\boldsymbol{\eta}}_{gt}\right\|}{\lambda_{gt}}v\right)$, is identified on $\mathbb{R}$ by
-$F_{\widetilde{\varepsilon}_{gt}^{*}}\left(-\overline{\mathbf{X}}_{gt}^{\prime}\overline{\boldsymbol{\eta}}_{gt}^{*}\right)=$ $\mathrm{E}\left[\Pr\left(Y_{gt}\leq 0\mid\overline{\mathbf{X}}_{gt}\right)\mid\overline{\mathbf{X}}_{gt}^{\prime}\overline{\boldsymbol{\eta}}_{gt}^{*}\right].$
-Given that the inverse function $F_{\widetilde{\varepsilon}_{gt}^{*}}^{-1}$ exists under Assumption 1, we identify
-$\left\|\overline{\boldsymbol{\eta}}_{gt}\right\|=\frac{1}{\kappa_{2,gt}^{*}}=\frac{1}{\mathrm{E}\left[F_{\widetilde{\varepsilon}_{gt}^{*}}^{-1}\left(\mathrm{E}\left[\Pr\left(Y_{gt}\leq 1\mid\mathbf{X}_{gt}\right)\mid\overline{\mathbf{X}}_{gt}^{\prime}\overline{\boldsymbol{\eta}}_{gt}^{*}\right]\right)+\overline{\mathbf{X}}_{gt}^{\prime}\overline{\boldsymbol{\eta}}_{gt}^{*}\right]}.$
-As an immediate consequence, $\overline{\boldsymbol{\eta}}_{gt}$ is identified.
-The distribution function $F_{\widetilde{\varepsilon}_{gt}}$ of the rescaled error term $\widetilde{\varepsilon}_{gt}$, defined by $F_{\widetilde{\varepsilon}_{gt}}\left(v\right):=F_{\varepsilon_{gt}}\left(\frac{v}{\lambda_{gt}}\right)$, is identified by a similar argument from above.
-In fact, given identification of the rescaled error term distribution, $\lambda_{gt}$ is also straightforwardly identified as the standard deviation of that distribution, as well as $F_{{\varepsilon}_{gt}}$.
-Now, focusing on $(g,t)=(1,1)$, identification of $\overline{\mathbf{x}}^{\prime}\overline{\boldsymbol{\eta}}_{11}$ and $\lambda_{11}$ directly follow from Proposition 1:
-$\overline{\mathbf{x}}^{\prime}\overline{\boldsymbol{\eta}}_{11}=\overline{\mathbf{x}}^{\prime}\left(\overline{\boldsymbol{\eta}}_{01}+\frac{\overline{\boldsymbol{\eta}}_{10}-\overline{\boldsymbol{\eta}}_{00}}{\lambda_{00}/\lambda_{01}}\right)\quad\text{and}\quad\lambda_{11}=\frac{\lambda_{01}\lambda_{10}}{\lambda_{00}},$
-for all covariate values $\mathbf{x}\in\mathcal{X}_{11}$.
-As $\overline{\mathcal{X}}_{11}$ does not lie in a proper linear subspace of $\mathbb{R}^{k+1}$ by Assumption 5, $\overline{\boldsymbol{\eta}}_{11}$ is identified as
-$\overline{\boldsymbol{\eta}}_{11}=\overline{\boldsymbol{\eta}}_{01}+\frac{\overline{\boldsymbol{\eta}}_{10}-\overline{\boldsymbol{\eta}}_{00}}{\lambda_{00}/\lambda_{01}}.$
-The identification of $F_{{\varepsilon}_{11}}$ directly follows from the strict stationarity in Assumption 1. ∎
-Proof of Proposition 4.
-Denote $\widetilde{\varepsilon}_{gt}:=\lambda_{gt}\varepsilon_{gt}$ and $\widetilde{\nu}_{gt}:=\zeta_{gt}\nu_{gt}$, as well as the composite distribution functions $F_{-\widetilde{\varepsilon}_{gt}}$ and $F_{-\widetilde{\nu}_{gt}}$ by
-$F_{-\widetilde{\varepsilon}_{gt}}\left(v\right):=F_{-\varepsilon_{gt}}\left(\frac{v}{\lambda_{gt}}\right)\quad\text{and}\quad F_{-\widetilde{\nu}_{gt}}\left(v\right):=F_{-\nu_{gt}}\left(\frac{v}{\zeta_{gt}}\right).$
-for any $v\in\mathbb{R}$. Under Assumption 6, it holds that
-$\Pr\left(C_{gt}(0)\geq j\mid\mathbf{X}_{gt},\mathbf{Z}_{gt}\right)$ $=\Pr\left(Y_{gt}(0)\geq j,R_{gt}(0)\geq j\mid\mathbf{X}_{gt},\mathbf{Z}_{gt}\right)$
-$=\Pr\left(\overline{\mathbf{X}}_{gt}^{\prime}\overline{\boldsymbol{\eta}}_{gt}+\lambda_{gt}\varepsilon_{gt}\geq\kappa_{j},\overline{\mathbf{Z}}_{gt}^{\prime}\overline{\boldsymbol{\pi}}_{gt}+\zeta_{gt}\nu_{gt}\geq\iota_{j}\right)$
-$=\overline{C}_{gt}\left(F_{-\widetilde{\varepsilon}_{gt}}\left(\overline{\mathbf{X}}_{gt}^{\prime}\overline{\boldsymbol{\eta}}_{gt}-\kappa_{j}\right),F_{-\widetilde{\nu}_{gt}}\left(\overline{\mathbf{Z}}_{gt}^{\prime}\overline{\boldsymbol{\pi}}_{gt}-\iota_{j}\right)\right).$
-Let $(g,t)\neq(1,1)$. By Assumptions 6 and 7, the signs of $\eta_{1,gt}$ and $\pi_{1,gt}$ are nonzero and identified as:
-$\operatorname{sgn}\left(\mathrm{E}\left[\frac{\partial}{\partial Z_{1,gt}}\Pr\left(C_{gt}\geq j|\mathbf{W}_{gt},X_{1,gt},Z_{1,gt}\right)\right]\right)$
-and
-$\operatorname{sgn}\left(\mathrm{E}\left[\frac{\partial}{\partial X_{1,gt}}\Pr\left(C_{gt}\geq j|\mathbf{W}_{gt},X_{1,gt},Z_{1,gt}\right)\right]\right),$
-respectively, for some $1\leq j\leq 2$, where $\operatorname{sgn}(x)=1\{x>0\}-1\{x<0\}$ is the sign function.
-W.l.o.g.
-let $\pi_{1,gt}>0$ for the rest of the proof, since otherwise we could consider $-\pi_{1,gt}$.
-Also, as the below exposition applies to each $1\leq j\leq 2$, we focus on a given $j$ level for simplicity.
-Under Assumption 6, given any $\mathbf{w}\in\mathbb{R}^{k-1}$, the function
-$(x,z)\mapsto{c}_{\mathbf{w},gt}(x,z):=\overline{C}_{gt}\left(F_{-\widetilde{\varepsilon}_{gt}}\left(\eta_{1,gt}x+\overline{\mathbf{w}}^{\prime}\overline{\boldsymbol{\eta}}_{-1,gt}-\kappa_{j}\right),F_{-\widetilde{\nu}_{gt}}\left(\pi_{1,gt}z+\overline{\mathbf{w}}^{\prime}\overline{\boldsymbol{\pi}}_{-1,gt}-\iota_{j}\right)\right).$
-is real analytic in each of its coordinates on $\mathbb{R}$, since the elementary functions are real analytic and the composition of real analytic functions are real analytic.
-Let $\eta^{*}_{1,gt},\lambda^{*}_{gt},\pi^{*}_{1,gt},\zeta^{*}_{gt}$ be arbitrary parameters in $\mathbb{R}$ with $\eta^{*}_{1,gt},\pi^{*}_{1,gt}>0$ and $\lambda^{*}_{gt},\zeta^{*}_{gt}>0$, $\overline{C}^{*}_{gt}$ an arbitrary copula function, and $F^{*}_{1,gt}$, $F^{*}_{2,gt}$ arbitrary distribution function such that $F^{*}_{l,gt}(x)=0.5$ if and only if $x=0$ for $l=1,2$, and
-$\Pr\left(C_{gt}\geq j\mid\mathbf{X}_{gt},\mathbf{z}_{gt}\right)$ $=\overline{C}_{gt}^{*}\left(F^{*}_{1,gt}\left(\overline{\mathbf{X}}_{gt}^{\prime}\overline{\boldsymbol{\eta}}^{*}_{gt}-\kappa_{j}\right),F^{*}_{2,gt}\left(\overline{\mathbf{Z}}_{gt}^{\prime}\overline{\boldsymbol{\pi}}^{*}_{gt}-\iota_{j}\right)\right).$
-Furthermore, given any $\mathbf{w}\in\mathbb{R}^{k-1}$, the composite function
-$(x,z)\mapsto{c}^{*}_{\mathbf{w},gt}(x,z):={\overline{C}}^{*}_{gt}\left(F^{*}_{1,gt}\left({{\eta}^{*}_{1,gt}x+\overline{\mathbf{w}}^{\prime}\overline{{\boldsymbol{\eta}}}^{*}_{-1,gt}-\kappa_{j}}\right),F^{*}_{2,gt}\left({\pi}^{*}_{1,gt}z+\overline{\mathbf{w}}^{\prime}\overline{{\boldsymbol{\pi}}}^{*}_{-1,gt}-\iota_{j}\right)\right)$
-is real analytic in each of its coordinates on $\mathbb{R}$. Now, fix $\mathbf{w}\in\mathcal{W}_{gt}$. Then, the composite function
-$(x,z)\mapsto{c}_{\mathbf{w},gt}(x,z)-{c}^{*}_{\mathbf{w},gt}(x,z)$
-is real analytic in each of its coordinates, and, by Assumption 7, is zero on the Cartesian product of two open intervals $I_{1}\times I_{2}$.
-This in turn implies that for any $x_{0}\in I_{1}$ and $z_{0}\in I_{2}$, the functions
-$z\mapsto{c}_{\mathbf{w},gt}(x_{0},z)-{c}^{*}_{\mathbf{w},gt}(x_{0},z),$
-$x\mapsto{c}_{\mathbf{w},gt}(x,z_{0})-{c}^{*}_{\mathbf{w},gt}(x,z_{0}),$
-are real analytic and zero on the open intervals $I_{2}$ and $I_{1}$, respectively.
-Focusing on the former function for identification of $\boldsymbol{\eta}_{gt}$, the identity theorem for real analytic functions guarantees that ${c}_{\mathbf{w},gt}(x_{0},\cdot)$ and ${c}^{*}_{\mathbf{w},gt}(x_{0},\cdot)$ coincide everywhere on $\mathbb{R}$.
-In fact, the limits $\lim_{z\rightarrow\infty}{c}_{\mathbf{w},gt}(x_{0},z)$ and $\lim_{z\rightarrow\infty}{c}^{*}_{\mathbf{w},gt}(x_{0},z)$ exist, are unique and coincide:
-$\lim_{z\rightarrow\infty}{c}_{\mathbf{w},gt}(x_{0},z)=$ $\lim_{z\rightarrow\infty}\overline{C}_{gt}\left(F_{-\widetilde{\varepsilon}_{gt}}\left(\eta_{1,gt}x_{0}+\overline{\mathbf{w}}^{\prime}\overline{\boldsymbol{\eta}}_{-1,gt}-\kappa_{j}\right),F_{-\widetilde{\nu}_{gt}}\left(\pi_{1,gt}z+\overline{\mathbf{w}}^{\prime}\overline{\boldsymbol{\pi}}_{-1,gt}-\iota_{j}\right)\right)$
-$=$ $F_{-\widetilde{\varepsilon}_{gt}}\left(\eta_{1,gt}x_{0}+\overline{\mathbf{w}}^{\prime}\overline{\boldsymbol{\eta}}_{-1,gt}-\kappa_{j}\right)$
-$=\lim_{z\rightarrow\infty}{c}^{*}_{\mathbf{w},gt}(x_{0},z)=$ $\lim_{z\rightarrow\infty}\overline{C}^{*}_{gt}\left(F^{*}_{1,gt}\left(\eta^{*}_{1,gt}x_{0}+\overline{\mathbf{w}}^{\prime}\overline{\boldsymbol{\eta}}^{*}_{-1,gt}-\kappa_{j}\right),F^{*}_{2,gt}\left(\pi^{*}_{1,gt}z+\overline{\mathbf{w}}^{\prime}\overline{\boldsymbol{\pi}}^{*}_{-1,gt}-\iota_{j}\right)\right)$
-$=$ $F^{*}_{1,gt}\left(\eta^{*}_{1,gt}x_{0}+\overline{\mathbf{w}}^{\prime}\overline{\boldsymbol{\eta}}^{*}_{-1,gt}-\kappa_{j}\right).$ (27)
-Because such identity holds for every $x_{0}\in I_{1}$, given Assumption 7 and by applying again the identity theorem for real analytic functions, the real analytic composite function
-$x\mapsto F_{-\widetilde{\varepsilon}_{gt}}\left(\eta_{1,gt}x+\overline{\mathbf{w}}^{\prime}\overline{\boldsymbol{\eta}}_{-1,gt}-\kappa_{j}\right)-F^{*}_{1,gt}\left(\eta^{*}_{1,gt}x+\overline{\mathbf{w}}^{\prime}\overline{\boldsymbol{\eta}}^{*}_{-1,gt}-\kappa_{j}\right)$
-is then zero on $\mathbb{R}$.
-In fact, as $\eta_{1,gt}\neq 0$, the individual functions also coincide at the point $x^{0}=\frac{\kappa_{j}-\overline{\mathbf{w}}^{\prime}\overline{\boldsymbol{\eta}}_{-1,gt}}{\eta_{1,gt}}$, such that $F_{-\widetilde{\varepsilon}_{gt}}$ is evaluated at its median:
-$0.5=F_{-\widetilde{\varepsilon}_{gt}}\left(0\right)=F^{*}_{1,gt}\left(\eta^{*}_{1,gt}\left(\frac{\kappa_{j}-\overline{\mathbf{w}}^{\prime}\overline{\boldsymbol{\eta}}_{-1,gt}}{\eta_{1,gt}}\right)+\overline{\mathbf{w}}^{\prime}\overline{\boldsymbol{\eta}}^{*}_{-1,gt}-\kappa_{j}\right),$
-However, it must be the case that $F^{*}_{1,gt}$ is evaluated at its median, too:
-$\eta^{*}_{1,gt}\left(\frac{\kappa_{j}-\overline{\mathbf{w}}^{\prime}\overline{\boldsymbol{\eta}}_{-1,gt}}{\eta_{1,gt}}\right)+\overline{\mathbf{w}}^{\prime}\overline{\boldsymbol{\eta}}^{*}_{-1,gt}-\kappa_{j}=0.$
-Rearrange the terms:
-$\left(\kappa_{j}-\overline{\mathbf{w}}^{\prime}\overline{{\boldsymbol{\eta}}}_{-1,gt}\right)\eta^{*}_{1,gt}+\eta_{1,gt}\overline{\mathbf{w}}^{\prime}\overline{\boldsymbol{\eta}}^{*}_{-1,gt}=\kappa_{j}\eta_{1,gt}$
-As the above exposition applies to each $1\leq j\leq 2$, we can build a system of linear equations:
-$-\overline{\mathbf{w}}^{\prime}\overline{{\boldsymbol{\eta}}}_{-1,gt}a+\eta_{1,gt}b$ $=0,$
-$\left(1-\overline{\mathbf{w}}^{\prime}\overline{{\boldsymbol{\eta}}}_{-1,gt}\right)a+\eta_{1,gt}b$ $=\eta_{1,gt}.$
-This system necessarily has full rank:
-$\operatorname{rank}\left(\left(\begin{array}[]{ll}-\overline{\mathbf{w}}^{\prime}\overline{{\boldsymbol{\eta}}}_{-1,gt}&\eta_{1,gt}\\ 1-\overline{\mathbf{w}}^{\prime}\overline{{\boldsymbol{\eta}}}_{-1,gt}&\eta_{1,gt}\end{array}\right)\right)=2.$
-Thus, $\left(\eta^{*}_{1,gt},\overline{\mathbf{w}}^{\prime}\overline{\boldsymbol{\eta}}^{*}_{-1,gt}\right)$ is uniquely determined as the solution in $(a,b)$ to the above system.
-Solving the system, we obtain
-$\left(\begin{array}[]{ll}\eta^{*}_{1,gt}\\ \overline{\mathbf{w}}^{\prime}\overline{\boldsymbol{\eta}}^{*}_{-1,gt}\end{array}\right)$ $=\left(\begin{array}[]{ll}-\overline{\mathbf{w}}^{\prime}\overline{{\boldsymbol{\eta}}}_{-1,gt}&\eta_{1,gt}\\ 1-\overline{\mathbf{w}}^{\prime}\overline{{\boldsymbol{\eta}}}_{-1,gt}&\eta_{1,gt}\end{array}\right)^{-1}\left(\begin{array}[]{ll}0\\ {\eta}_{1,gt}\end{array}\right)$
-$=-\frac{1}{\eta_{1,gt}}\left(\begin{array}[]{ll}\eta_{1,gt}&-\eta_{1,gt}\\ \overline{\mathbf{w}}^{\prime}\overline{{\boldsymbol{\eta}}}_{-1,gt}-1&-\overline{\mathbf{w}}^{\prime}\overline{{\boldsymbol{\eta}}}_{-1,gt}\end{array}\right)\left(\begin{array}[]{ll}0\\ {\eta}_{1,gt}\end{array}\right)$
-$=\left(\begin{array}[]{ll}\eta_{1,gt}\\ \overline{\mathbf{w}}^{\prime}\overline{{\boldsymbol{\eta}}}_{-1,gt}\end{array}\right).$
-As a result, $\eta_{1,gt}$ and $\overline{\mathbf{w}}^{\prime}\overline{{\boldsymbol{\eta}}}_{-1,gt}$ are identified.
-Furthermore, as the above exposition applies to every $w\in\mathcal{W}_{gt}$, and $\overline{\mathcal{W}}_{gt}$ does not lie in a proper linear subspace of $\mathbb{R}^{k}$ by Assumption 7, $\overline{\boldsymbol{\eta}}_{-1,gt}$ is identified.
-This implies in turn that
-$F_{-{\widetilde{\varepsilon}}_{gt}}\left(v\right)=F^{*}_{1,gt}\left(v\right),$
-for all $v\in\mathbb{R}$, establishing identification of $F_{-{\widetilde{\varepsilon}}_{gt}}$.
-Consequently, $\lambda_{gt}$ is identified as the standard deviation of this distribution, as well as $F_{{\varepsilon}_{gt}}$.
-Identification of $\overline{\boldsymbol{\pi}}_{gt}$, $\zeta_{gt}$ and $F_{-{\nu}_{gt}}$ is analogous.
-Now, focusing on $(g,t)=(1,1)$, given the distributional parallel trends, identification of $\overline{\mathbf{x}}^{\prime}\overline{\boldsymbol{\eta}}_{11}$ and $\lambda_{11}$ directly follow from Proposition 1:
-$\overline{\mathbf{x}}^{\prime}\overline{\boldsymbol{\eta}}_{11}=\overline{\mathbf{x}}^{\prime}\left(\overline{\boldsymbol{\eta}}_{01}+\frac{\overline{\boldsymbol{\eta}}_{10}-\overline{\boldsymbol{\eta}}_{00}}{\lambda_{00}/\lambda_{01}}\right)\quad\text{and}\quad\lambda_{11}=\frac{\lambda_{01}\lambda_{10}}{\lambda_{00}},$
-for all covariate values $\mathbf{x}\in\mathcal{X}_{11}$.
-As $\overline{\mathcal{X}}_{11}$ does not lie in a proper linear subspace of $\mathbb{R}^{k+1}$ by Assumption 5, $\overline{\boldsymbol{\eta}}_{11}$ is identified as
-$\overline{\boldsymbol{\eta}}_{11}=\overline{\boldsymbol{\eta}}_{01}+\frac{\overline{\boldsymbol{\eta}}_{10}-\overline{\boldsymbol{\eta}}_{00}}{\lambda_{00}/\lambda_{01}}.$
-Identification of $\overline{\boldsymbol{\pi}}_{11}$ and $\zeta_{gt}$ is analogous:
-$\overline{\boldsymbol{\pi}}_{11}=\overline{\boldsymbol{\pi}}_{01}+\frac{\overline{\boldsymbol{\pi}}_{10}-\overline{\boldsymbol{\pi}}_{00}}{\zeta_{00}/\zeta_{01}}\quad\text{ and }\quad\zeta_{11}=\frac{\zeta_{01}\zeta_{10}}{\zeta_{00}}.$
-Proof of Proposition 5.
-The result follows by applying Proposition 3 for the identification of the distribution of $Y_{11}^{\ast}(0)$, and Assumption 8. ∎
-Proof of Proposition 6.
-The result follows by applying Proposition 4 for the identification of the marginal distribution of $Y_{11}^{\ast}(0)$, and by using Assumption 9 together with arguments from the proof of Proposition 4 to identify the marginal distribution of $Y_{11}^{\ast}(1)$.
-∎
-Proof of Proposition 7.
-The proof follows from Assumptions 1, 10, and 11 as well as an application of the Delta Method. ∎